### v1

# Get the Agents Drunk: Memory Perturbations in Autonomous Agent-based Recommender Systems

Abstract.
Large language model-based agents are increasingly used in recommender systems (Agent4RSs) to achieve personalized behavior modeling. Specifically, Agent4RSs introduces memory mechanisms that enable the agents to autonomously learn and self-evolve from real-world interactions. However, to the best of our knowledge, how robust Agent4RSs are remains unexplored. As such, in this paper, we propose the first work to attack Agent4RSs by perturbing agents’ memories, not only to uncover their limitations but also to enhance their security and robustness, ensuring the development of safer and more reliable AI agents.
Given the security and privacy concerns, it is more practical to launch attacks under a black-box setting, where the accurate knowledge of the victim models cannot be easily obtained. Moreover, the practical attacks are often stealthy to maximize the impact. To this end, we propose a novel practical attack framework named DrunkAgent. DrunkAgent consists of a generation module, a strategy module, and a surrogate module. The generation module aims to produce effective and coherent adversarial textual triggers, which can be used to achieve attack objectives such as promoting the target items. The strategy module is designed to ‘get the target agents drunk’ so that their memories cannot be effectively updated during the interaction process. As such, the triggers can play the best role. Both of the modules are optimized on the surrogate module to improve the transferability and imperceptibility of the attacks. By identifying and analyzing the vulnerabilities, our work provides critical insights that pave the way for building safer and more resilient Agent4RSs. Extensive experiments across various real-world datasets demonstrate the effectiveness of DrunkAgent.
Adversarial Attacks, Recommender Systems, Large Language Model, Generative Agents, Collaborative Filtering
copyright: noneccs: Security and privacy Web application security
1. Introduction
Recommender systems (RSs) have evolved into a vital component of the modern digital ecosystem. From e-commerce sites (e.g., Amazon, Taobao) and streaming platforms (e.g., YouTube, TikTok) to social media networks (e.g., Facebook, Instagram), RSs power personalized experience, drive user engagement of these online platforms, and further influence market dynamics through item promotions (Yang et al., 2024, 2023). The foundation of the RSs lies in modeling users’ preferences based on their historical interactions with items, such as purchases and ratings, known as collaborative filtering (CF). Despite the success of CF techniques (Koren et al., 2009; He et al., 2017, 2020), they often struggle with data sparsity and cold-start problems (Sachdeva and McAuley, 2020). To alleviate the problems, the auxiliary information with rich semantics (e.g., item descriptions and user reviews) has been integrated into RSs (Zhang et al., 2019). Given the increasing volume and complexity of recommendation data, the RSs encounter difficulties in scaling effectively and capturing textual knowledge about users and items sufficiently, due to they generally rely on traditional deep learning (DL) models like CNNs and RNNs (Zhao et al., 2023).
Figure 1. An exemplar of attacks on agent-based recommender systems, where the target item’s descriptions have been manipulated to perturb the memory of the target agent for promoting the target item to more normal users. By investigating this, our work reveals emerging vulnerabilities in memory-augmented agents and provides insights for building more secure recommender systems.
Recently, large language models (LLMs) have shown potential to mitigate the above problems and revolutionize RSs (Zhao et al., 2023), as they are primarily built upon advanced transformer-based architectures and pretrained on diverse and extensive datasets. Most of works carefully design textual prompts for LLMs based on user-item interactions (Geng et al., 2022; Bao et al., 2023; Zhang et al., 2023; Wu et al., 2024). As such, the abundant knowledge inside the LLM’s training corpora can be fully exploited. However, the LLM-based RSs (LLM4RSs) struggle to capture underlying behavioral patterns, due to the gap between universal language modeling and personalized behavior modeling (Zhang et al., 2024b). Integrating agents built upon LLMs equipped with additional modules offers a promising solution (Durante et al., 2024; Huang et al., 2024), allowing to autonomously explore and learn from external sources. Among all the added modules, memory is a key component to support agent-environment [^1] In a narrow sense, environment is the object that the agent needs to interact with to accomplish the task. More broadly, environment can be any contextual factors (e.g., item titles, user-item interactive behaviors) that influence the agent’s decisions (Zhang et al., 2024a). interactions, which plays a crucial role in helping agents accumulate historical experience, retrieve informative knowledge and self-evolve in real-world environments (Zhang et al., 2024a; Wang et al., 2024b). Consequently, the agent-based RSs with powerful LLMs (Huang et al., 2024; Zhang et al., 2024b) can also leverage memory modules that store the information explored from the external environments to further improve the recommendation quality. Fig.  1 gives an example of an item agent interacting with its situated environment (e.g., users and other items), where the learning knowledge will be placed in its memory. Despite the aforementioned success, the safety vulnerability of the generative agent-based RSs (Agent4RSs) under adversarial attacks has yet to be explored. Such attacks can gain insights into the robustness of the systems and help develop effective defensive mechanisms. As recommender agents are integrated into more complex real-world environments, understanding how these attacks might compromise their behaviors is critical for building trustworthy and robust recommender agents (Chua et al., 2024).
Due to security and privacy considerations, the access of accurate knowledge of the victim model, including architectures, gradients, parameters, etc., is generally restricted in practice. As such, practical attacks are mainly under a black-box setting (Ning et al., 2024; Yang et al., 2024; Lin et al., 2022; Zhang et al., 2024d; Zeng et al., 2023). Practical attacks are often required to have transferability (i.e., effective against different black-box RSs). Moreover, to maximize the impact of the attack and the number of affected users, practical attacks are often rewarded by imperceptibility (i.e., difficult to be discovered) (Liu and Larson, 2021).
Conventional black-box shilling attacks (Wang et al., 2024a), which inject well-crafted fake user profiles (e.g., a set of users’ numerical ratings on items) into the victim model for certain attack objectives (e.g., promote their target items for profits), have been shown to effectively poison DL-based RSs. However, most attacks have difficulty in processing textual inputs (e.g., item titles and descriptions) and lack context awareness, making it challenging to effectively target LLM-empowered RSs that rely primarily on text inputs to generate natural language responses (Ning et al., 2024). Item providers are generally responsible for creating and maintaining their own product representations on the platforms with recommendation services in practice (Yang et al., 2024; Cohen et al., 2021; Liu and Larson, 2021) (e.g., YouTube allows creators to upload item titles and Amazon enables sellers to upload item descriptions), and current RSs depend heavily on these representations (Zhang et al., 2024d; Zhao et al., 2023). As such, security vulnerabilities in these systems naturally emerge. The vulnerabilities can easily be exploited by profit-driven merchants manipulating product representations, as shown in Fig.  1. Compared to the shilling attacks that involve hiring online writers to introduce fake reviews to indirectly promote the items, this type of attack is more direct, efficient, and lower in economic cost (Zhang et al., 2024d).
Recent black-box textual attacks on LLM4RSs are far from practical. Classical text attacks that are successful in text mis-classifications, such as TextFooler (Jin et al., 2020) and PuncAttack (Formento et al., 2023), have been directly applied for LLM4RSs (Zhang et al., 2024d). However, the effectiveness of the attacks is limited, due to there is a gap between the attack goals of the text attacks and RS attacks. Trivial and ChatGPT attacks (Zhang et al., 2024d), on the other hand, are proposed to be tailored for LLM4RSs, which increases the items’ attractiveness by appending positive words to the titles. Nevertheless, such attacks are not optimized on any black-box RSs, making them poorly transferable to the Agent4RSs. Moreover, since the attacks mainly target modifications to item titles, where titles’ brevity and frequent exposure to users tend to make even small perturbations noticeable.
To tackle the above problems, in this paper, we propose a new practical attack framework named DrunkAgent. Instead of manipulating the item titles, to enhance the imperceptibility of attacks, we propose to perturb the description of the target item, where length and context can easily hide changes. DrunkAgent is composed of three modules: generation module, strategy module and surrogate module. The generation module is designed to produce effective adversarial textual description triggers, which can be used to promote the target items to as many normal users as possible on the agent-powered RSs. The triggers are coherent, semantically meaningful and human-readable text, making the attack stealthy. The strategy module aims to ‘get the target agents drunk’, that is, to confuse the agents’ memory mechanisms so that they cannot be effectively updated during agent-environment interactions. In this way, the triggers can play the best role. To avoid frequently querying the victim models to induce suspicion and mimic the black-box settings in the real world, the generation module and the strategy module are optimized on the surrogate module to improve the transferability and imperceptibility of the practical attacks.
Beyond demonstrating the attack capabilities of DrunkAgent, our broader goal is to reveal and understand the overlooked vulnerabilities in Agent4RSs. Specifically, by analyzing how agents’ memory mechanisms can be influenced through carefully crafted and stealthy textual triggers, our work provides a new lens into the emerging attack surfaces introduced by memory-augmented, LLM-based agents. We believe that this exploration can offer valuable insights to the community for designing more resilient agent architectures, guiding future efforts on defense mechanisms such as memory validation or input sanitization.
Our main contributions are summarized as follows:
  • We identify a safety vulnerability in autonomous agent-based RSs with powerful LLMs. To the best of our knowledge, we present the first formal and systematic study of adversarial textual attacks on the agent-powered RSs, aiming to inform future robustness research.
  • We propose to perturb the description of the target item for effective memory perturbations, which is an unnoticeable and crucial property of current agentic recommendations.
  • We propose a novel practical attack framework on the agent-powered RSs named DrunkAgent, which can create adversarial descriptions of target items with transferability and imperceptibility.
  • Our DrunkAgent significantly outperforms state-of-the-art black-box attacks against various LLM-based agentic RSs on real-world datasets in terms of transferability and imperceptibility. Moreover, state-of-the-art defense strategies are ineffective to DrunkAgent, demonstrating the need for improved defense mechanisms to guide more secure Agent4RSs.
2. Problem Formulation
2.1. Victim LLM-powered Agentic RSs
We use $\mathcal{Y}=\{y_{u,v}:u\in\mathcal{U},v\in\mathcal{V}\}$ to denote the records of the user-item interaction matrix in the recommendation space, where $\mathcal{U}$ and $\mathcal{V}$ represent the set of real users and the item universe, respectively. $\mathcal{V}_{u}=\{v\in\mathcal{V}:y_{u,v}\neq 0\}$ indicates the set of items that have been interacted by $u$ (i.e., the user’s profile). LLM-powered RSs generally conduct recommendations by integrating user historical behavior sequences and/or item and/or user features with personalized prompt templates (Geng et al., 2022; Zhao et al., 2023; Bao et al., 2023; Zhang et al., 2023). Let $\mathcal{X}_{u,v}$ represent the prompts of LLM-powered RSs (denoted as $\text{LLM}_{\Theta}$ with parameters $\Theta$), the preference function can be formulated as below.
(1) $\mathcal{R}_{u}=f_{\text{LLM}_{\Theta}}(\mathcal{X}_{u,v}),$
where $\mathcal{X}_{u,v}=\mathcal{P}\oplus m_{u}\oplus\mathcal{M}_{v}^{u}\oplus\mathcal{C}_{v}^{u}$ and $\mathcal{R}_{u}$ is the recommendation results for $u$. $\oplus$ denotes the textual string concatenation. $\mathcal{P}$ stands for prompt templates of recommendations. $m_{u}$ indicates user features and $\mathcal{M}_{v}^{u}=\{m_{v}:v\in\mathcal{V}_{u}\}$ represents the feature set of $u$’s interacted items. $\mathcal{C}_{v}^{u}$ denotes the feature set of candidate items that are not interacted by $u$, where the amount often depends on the type of task (Xu et al., 2021). $|\mathcal{C}_{v}^{u}|>1$ if the task is from implicit feedback such as top-$K$ recommendations, and $|\mathcal{C}_{v}^{u}|=1$ if it is formulated as an explicit feedback task such as rating predictions. In addition to basic textual metadata (e.g., item titles), the features in the vanilla LLM-based RSs include user/item ID information, while the features in the generative agent-based RSs involve the memories of user/item agents. Note that all the elements of $\mathcal{X}_{u,v}$ are optional, which relies on its design.
2.2. Threat Model
2.2.1. Attacker’s Knowledge:
As the works (Ning et al., 2024; Yang et al., 2024; Lin et al., 2022; Zhang et al., 2024d; Zeng et al., 2023) point out, the inherent details (e.g., the architectures, gradients and parameters) of the victim RSs often cannot be obtained in practice. As such, our attack is performed under the black-box settings. In addition, the attacker is only able to access publicly available recommendation data (e.g., item titles and user reviews that can be collected directly by browsing or writing crawlers), due to the openness nature of RSs (Si and Li, 2020; Yang et al., 2023).
2.2.2. Attacker’s Capability:
Unlike another state-of-the-art LLM-powered RS attack (Ning et al., 2024) with an unrealistic assumption of $\mathcal{P}$ and $\mathcal{V}_{u}$ can be easily perturbed (that will be detailed in Section 5), we consider that the attackers have limited perturbation capability in practice. Specifically, they can only interfere with the representation of the target item, which is feasible and can be easily accomplished via a dedicated API, as the existing practical attacks do (Zhang et al., 2024d; Yang et al., 2024; Liu and Larson, 2021).
2.2.3. Attacker’s Objectives:
As mentioned in Section 1, two objectives: transferability and imperceptibility are considered for the practical attack. Given that push attacks are generally used by unscrupulous parties on RSs for profits in practice (Si and Li, 2020) with the aim of promoting target items to as many normal users as possible (Zhang et al., 2021), we thus focus on such attacks in this paper, as normally do in recent works (Wang et al., 2024a; Zhang et al., 2024d). Our first goal is to maximize the hit ratio of the target item $t$ on a surrogate RS (denoted as $\text{Agent}_{\Phi}$ with parameters $\Phi$) that is different from the victim black-box RS, ensuring transferability. Our second goal is to make the descriptions of $t$ as stealthy as possible, achieving imperceptibility. As such, the surrogate RS is used to optimize attacks to avoid frequently querying the victim models to induce suspicion.
Figure 2. Practical Attack Framework DrunkAgent Overview
3. DrunkAgent
In this section, we present a practical attack framework, DrunkAgent, for gaining insights into the robustness of autonomous agent-based RSs with powerful LLMs and related defense models in practice. DrunkAgent is composed of three attack modules. Fig.  2 gives an overview of DrunkAgent, where each component will be described in the following.
3.1. Surrogate Module
We leverage a surrogate RS $\Phi$ to measure how effective the attack is under the black-box settings, as inspired by the existing works (Lin et al., 2022; Tang et al., 2020; Yang et al., 2023, 2024). It is used to optimize the generation module and the strategy module, making the attack transferable and imperceptible.
Since our attack targets SOTA LLM-powered RSs, we adopt powerful LLMs as recommendation agents. Given the transparency, flexibility and scalability of the open-source LLM models, they can be considered for building this module. Moreover, due to our attack focus on adversarial text description generation tasks, we consider an autoregressive language model trained with a causal language modeling objective (Radford et al., 2019; Brown et al., 2020). There are some notable models, such as LLaMA (Touvron et al., 2023a), LLaMA 2 (Touvron et al., 2023b) and LLaMA 3 (Dubey et al., 2024), BLOOM 176B (Le Scao et al., 2023) and Falcon 180B (Almazrouei et al., 2023). Currently, Meta’s LLaMA family has become the most influential open LLM and has been widely adopted in the recent recommendation community (Zhao et al., 2023; Bao et al., 2023). To ensure the effectiveness of attacks, we apply the latest generative model: LLaMA 3, so that valuable feedback can be provided even with limited available resources. Specifically, LLaMA 3 uses an optimized transformer architecture (Vaswani et al., 2017). Its instruction-tuned version uses supervised fine-tuning and reinforcement learning from human feedback (RLHF) to align with human preferences for helpfulness and safety (AI, 2024), as shown in Fig.  2.
In addition to conducting a strict selection of recommendation agents, we develop the two modules for effective target item’s adversarial description generation.
3.2. Generation Module
In this section, we aim to generate adversarial description triggers on the agent-powered RSs, which can promote the target items to as many normal users as possible. The triggers should have three critical properties: 1) Universality on any users and items; 2) Transferability among different black-box victim models; and 3) Imperceptibility to not easier to be discovered. To this end, we design sub-modules, as shown in Fig.  2 and described as follows.
3.2.1. Adversarial Prompting Engineering
Recently, many of the mainstream LLMs that have been widely used in the agentic recommendation community (Huang et al., 2024) (e.g., GPT-3.5-turbo, ChatGPT) are closed source, where the models’ underlying codes, weights, or other critical components are not released to the public and the models can only be invoked through their APIs. As a result, the agent-powered RSs generally adopt in-context learning (ICL) (Dong et al., 2022) to sufficiently take advantage of the generalization capability of LLMs on unknown tasks. To this end, this module carefully crafts prompts $\widetilde{\mathcal{X}_{t}}$ within the ICL framework to improve the quality of adversarial trigger generation.
Basic Prompt Template. We meticulously prepare prompt templates $\widetilde{\mathcal{P}}$ to provide context, goals, instructions and constraints so that the model knows ‘what needs to be done’ and ‘how to respond to’. The main components are described in the following.
  1. (1)
    Style Definition $\widetilde{p_{s}}$. There are two common styles of prompts: task-oriented and role-oriented (Zhu et al., 2023a). To make recommendations more authentic and reasonable, we introduce role information into the context, which allows the system to respond with domain-specific knowledge, such as ‘You are a CD enthusiast.’, as shown in Fig.  2.
  2. (2)
    Task Goal $\widetilde{p_{g}}$. Since a recommender system is essentially a ranker (Zhang et al., 2019), we set sorting goals to indicate recommendation tasks, such as ‘Now, you are considering to sort ten candidate CDs in order of how well they align with your preferences.’ in the given example of the figure.
  3. (3)
    Recommendation Instruction $\widetilde{p_{i}}$. We introduce highly detailed instructions to guide the model to perform accurate recommendations. For example, ‘Please sort CDs by considering the correlation between your preferen -ces/dislikes and the features of the CDs. The higher the CD rank, the more you like it.’ in the shown example.
  4. (4)
[…]
Hence, the adversarial prompt template can be formulated as:
(2) $\widetilde{\mathcal{P}}=\widetilde{p_{s}}\oplus\widetilde{p_{g}}\oplus\widetilde{p_{i}}\oplus\widetilde{p_{f}}.$
To increase the attack effectiveness, we introduce position bias (Zhang et al., 2024b) into $\widetilde{\mathcal{P}}$, where the other candidates items are placed before the target item, as shown in the example of Fig.  2.
General User Memories. It is difficult to obtain the optimized memories of user agents (i.e., language-based parametric embeddings of the agentic RSs (Zhang et al., 2024b)), due to the practical attack is launched under the black-box settings. Moreover, the adversarial description triggers are designed to be effective for general users, rather than specific user groups (e.g., users who are only obsessed with Progressive Metal music), so that the target items can be promoted to as many users as possible. Therefore, to enhance the universality of the triggers on users, we use primitive and general user descriptions as user memories $\widetilde{m_{u}}$, such as ‘I enjoy listening to CDs very much.’ in the given example, as the default user memories in the existing works (Zhang et al., 2024b)(Huang et al., 2024).
Popular Item Memories. Given that popular items usually have a high probability of appearing at the top of the users’ recommendation lists, Bandwagon attack (Si and Li, 2020) introduces popular items into fake user profiles to poison traditional RSs for increasing the popularity of the target item and Expert attack on visually-aware RSs (Liu and Larson, 2021) creates adversarial images by decreasing the representation distance to popular items for improving the exposure rate of the target item. Inspired by this, we select popular items as candidate items for $\widetilde{\mathcal{X}_{t}}$ to introduce popularity bias (Zhang et al., 2024b) to further enhance the effectiveness of adversarial description triggers. The basic idea is that the target items should be promoted to more normal users if they can be ranked ahead of the popular items. Let $\mathcal{U}_{v}=\{u\in\mathcal{U}:y_{u,v}\neq 0\}$ indicates the set of users that have interacted with $v$ and $\mathcal{M}_{p}$ denotes a sequence of the features of popular items,
(3) $\mathcal{M}_{p}=\{m_{v}\},$
where $m_{v}$ is the features of the item $v$ including titles and descriptions. $\{v\}_{v\in\mathcal{V}}$ is a sequence of popular item IDs, which is corresponding to the subscript of the item popularity sequence in descending order $\{|\mathcal{U}_{v}|\}$. The item popularity is item frequency here, due to this work focuses on an implicit task.
The descriptions of candidate items will greatly influence the feedback of the surrogate model. As with optimized user agent memories, the adversary is hard to access to internal memories of item agents in practice and is restricted to leverage publicly available data (e.g., item titles and descriptions). However, the raw descriptions (Ni et al., 2019) are often too long that may exceed the context window limitations of the surrogate model and always contain too much redundant information that may increase the difficulty of subsequent optimizations. Similar to the use of user memories, we can consider the default item memories for popular items. Nevertheless, the default ones generally only include identity information such as titles and categories (Zhang et al., 2024b), as the example “The CD is called ‘Brainwashed’. The category of this CD is: ‘Classic Rock; Album-Oriented Rock (AOR)’.” shows, which are overly simple and information-less, leading to the target items easier to be ranked at the top and the sub-optimal results will be produced. To tackle these issues, we propose to adopt a LLM with frozen parameters $\Psi$ to extract valuable item features from the raw descriptions and supplement the features based on its internal rich knowledge base to make them more informative to facilitate the optimizations. Let $\mathcal{C}_{p}$ represent the modified feature sequence of popular items,
(4) $\mathcal{C}_{p}=f_{\text{LLM}_{\Psi}}(\mathcal{X}_{p}\oplus\mathcal{M}_{p}),$
where $\mathcal{X}_{p}$ denotes the prompts for feature acquisition and OpenAI’s GPT-4 (Achiam et al., 2023) is adopted to implement the LLM, because it is a state-of-the-art model with powerful language understanding, reasoning and generation capabilities. An example of producing a popular item memory is given in Fig.  2. To sum up, the adversarial prompt can be formulated as
(5) $\widetilde{\mathcal{X}_{t}}=\widetilde{\mathcal{P}}\oplus\widetilde{m_{u}}\oplus\mathcal{C}_{p}\oplus\widetilde{m_{t}},$
where $\widetilde{m_{t}}$ is the features of the target item. In addition to the fixed title, the adversarial description trigger (i.e., the memory of the item agent) will be optimized in the next sub-section.
Note that we give lightweight and efficient adversarial prompts for the surrogate RS that are different from the real victim RSs’. To ensure high-quality recommendations, more powerful and context-rich (e.g., retrieval-augmented) prompts will be used in the victims, which will be described in the evaluation section.
3.2.2. Greedy Search Algorithm
We propose a greedy search algorithm to optimize the adversarial description triggers, as inspired by the successes of the current jailbreak attacks on LLMs (Zou et al., 2023; Liu et al., 2024b). However, there are some essential differences (e.g., attack goals and victim models) between jailbreak attacks and our attacks. Jailbreak attacks are designed to undermine the LLMs’ human-aligned values or other constraints imposed by model developers, compelling them to respond to malicious questions posed by adversaries with correct answers rather than rejection, whereas our attacks aim to promote the target items on the RSs powered by autonomous agents.
Adversarial Trigger Initialization. Given that LLMs should understand LLM-powered agents better, we leverage a LLM to generate some description candidates. To narrow search space by a large margin to increase the efficiency and effectiveness of the algorithm, the basic attack goals are incorporated into the prompts. In addition, the diversity and the generalizability are introduced to the prompts to help search for the global optimal solution and improve their universality on items, respectively. Formally,
(6) ~$\widetilde{\mathcal{M}}_{c}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{c}}\oplus\widetilde{m_{t}}),$
where ~$\widetilde{\mathcal{M}}_{c}=\{\widetilde{m}_{t_{i}}\}$ represents a sequence of trigger candidates obtained from the LLM and the state-of-the-art GPT-4 is used to implement the LLM. $\widetilde{\mathcal{X}_{c}}$ denotes a sequence of prompt templates. An example of the initialization is given in Fig.  2.
Adversarial Trigger Optimization. The algorithm takes $\mathcal{E}$ epochs to obtain the optimal trigger, where the output candidates of each epoch will be the inputs of the next epoch for continuous optimization until the attack objective is achieved or the termination condition is reached (e.g., the maximum number of epochs is exhausted). To avoid frequently querying the victim models to induce suspicion and mimic the black-box settings in the real world, the optimization is built on the surrogate module, as shown in Fig.  2.
For each epoch $k$, the algorithm first calculate the performance score for each input candidate $\widetilde{m}_{t_{i}}^{(k-1)}$, where $n$ well-performing trigger candidates will be maintained directly for the next optimization to ensure the stability. The higher the score, the more likely the candidate is to be kept. The score is calculated by the negative of the loss, i.e., $s_{i}^{(k-1)}=-\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}$. The smaller the loss, the more likely the target item is to be ranked high and the more effective the candidate is. We use the auto-regressive language generation loss (i.e., negative log-likelihood) to evaluate the discrepancy between the predictions and the target output $\widetilde{\mathcal{R}_{t}}$,
(7) $\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}=-\log\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)})),$
[…]
(8) $\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)}))=\prod_{l=1}^{L}\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{r_{t_{l}}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)}\oplus\widetilde{r_{t_{<l}}})),$
Then, the algorithm samples a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates by conducting probability-based random selection, where the candidates with higher scores will have a greater probability of being selected into this subset. The selection probability of each candidate is calculated by a softmax function:
(9) $\text{softmax}(\widetilde{m}_{t_{i}}^{(k-1)})=\frac{e^{s_{i}^{(k-1)}}}{\sum_{j=1}^{|\widetilde{\mathcal{M}}_{c}|}e^{s_{j}^{(k-1)}}}.$
For the obtained subset, the algorithm splits the candidates at random positions into a list of text slices and performs the random exchange of the slices between pairwise texts. For example, for the pairwise $<\widetilde{m}_{t_{i}}^{(k-1)},\widetilde{m}_{t_{j}}^{(k-1)}>$, $\widetilde{m}_{t_{i}}^{(k-1)}=[\widetilde{m}_{t_{i_{1}}}^{(k-1)},\widetilde{m}_{t_{i_{2}}}^{(k-1)},\widetilde{m}_{t_{i_{3}}}^{(k-1)}]$ and $\widetilde{m}_{t_{j}}^{(k-1)}=[\widetilde{m}_{t_{j_{1}}}^{(k-1)},\widetilde{m}_{t_{j_{2}}}^{(k-1)},\widetilde{m}_{t_{j_{3}}}^{(k-1)},\widetilde{m}_{t_{j_{4}}}^{(k-1)}]$, the new candidates could be $\widetilde{m}_{t_{i}}^{(k-1)^{\prime}}=[\widetilde{m}_{t_{i_{1}}}^{(k-1)},\widetilde{m}_{t_{j_{2}}}^{(k-1)},\widetilde{m}_{t_{i_{3}}}^{(k-1)},\widetilde{m}_{t_{j_{4}}}^{(k-1)}]$ and $\widetilde{m}_{t_{j}}^{(k-1)^{\prime}}=[\widetilde{m}_{t_{j_{1}}}^{(k-1)},\widetilde{m}_{t_{i_{2}}}^{(k-1)},\widetilde{m}_{t_{j_{3}}}^{(k-1)}]$. As a result, the new combinations of different words, phrases, and sentences are gained, which increases the feature diversity of candidates through extensive exploration, contributing to search for the optimal trigger.
Table 1. Attack Transferability. HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ of different attacks against various black-box victim LLM-powered agent-based RSs on real-world datasets. We use bold fonts to denote the best performance. Note that the attacks with higher HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ have excellent transferability.
Victim RS AgentCF
Attack CDs & Vinyl Office Products Musical Instruments
H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3
Benign 0.0505 0.1111 0.1616 0.0505 0.0887 0.1140 0.0306 0.0612 0.0918 0.0306 0.0499 0.0652 0.0412 0.0515 0.1134 0.0412 0.0477 0.0787
DeepwordBug 0.0808 0.1919 0.3232 0.0808 0.1509 0.2166 0.0000 0.0306 0.0510 0.0000 0.0193 0.0295 0.0206 0.0515 0.1031 0.0206 0.0401 0.0659
PuncAttack 0.0505 0.1313 0.2323 0.0505 0.1015 0.1520 0.0612 0.1224 0.1531 0.0612 0.0999 0.1152 0.0103 0.0309 0.0515 0.0103 0.0233 0.0336
TextFooler 0.0707 0.1717 0.2626 0.0707 0.1344 0.1799 0.0510 0.0816 0.1224 0.0510 0.0703 0.0907 0.0206 0.0309 0.0515 0.0206 0.0271 0.0374
BertAttack 0.0303 0.1111 0.2020 0.0303 0.0813 0.1267 0.0000 0.0306 0.0510 0.0000 0.0193 0.0295 0.0309 0.0515 0.0825 0.0309 0.0439 0.0594
TrivialInsertion 0.0202 0.1111 0.1818 0.0202 0.0776 0.1129 0.0306 0.0816 0.1531 0.0306 0.0628 0.0985 0.0103 0.0206 0.0619 0.0103 0.0168 0.0374
ChatGPTAttack 0.0606 0.1515 0.2626 0.0606 0.1180 0.1735 0.0510 0.1122 0.1633 0.0510 0.0896 0.1152 0.0103 0.0412 0.0825 0.0103 0.0298 0.0504
DrunkAgent 0.4040 0.4141 0.4343 0.4040 0.4104 0.4205 0.2449 0.2449 0.2551 0.2449 0.2449 0.2500 0.2268 0.2268 0.2371 0.2268 0.2268 0.2320
Victim RS AgentRAG
Attack CDs & Vinyl Office Products Musical Instruments
H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3
Benign 0.0505 0.1111 0.2222 0.0505 0.0887 0.1443 0.0510 0.0510 0.0816 0.0510 0.0510 0.0663 0.0206 0.0412 0.1031 0.0206 0.0336 0.0646
DeepwordBug 0.0202 0.1010 0.2626 0.0202 0.0712 0.1520 0.0000 0.0204 0.0816 0.0000 0.0129 0.0435 0.0000 0.0206 0.0722 0.0000 0.0130 0.0388
PuncAttack 0.0404 0.1818 0.2626 0.0404 0.1296 0.1700 0.0510 0.0612 0.1122 0.0510 0.0575 0.0830 0.0206 0.0309 0.0619 0.0206 0.0271 0.0426
TextFooler 0.0606 0.1111 0.2222 0.0606 0.0925 0.1480 0.0510 0.1122 0.1429 0.0510 0.0896 0.1050 0.0000 0.0103 0.0412 0.0000 0.0065 0.0220
BertAttack 0.0404 0.1010 0.2020 0.0404 0.0786 0.1291 0.0000 0.0000 0.0102 0.0000 0.0000 0.0051 0.0103 0.0309 0.0309 0.0103 0.0233 0.0233
TrivialInsertion 0.0000 0.0707 0.2020 0.0000 0.0446 0.1103 0.0408 0.0816 0.1327 0.0408 0.0666 0.0921 0.0103 0.0206 0.0412 0.0103 0.0168 0.0271
ChatGPTAttack 0.0606 0.1616 0.2525 0.0606 0.1243 0.1698 0.0816 0.1122 0.1735 0.0816 0.1009 0.1316 0.0103 0.0103 0.0412 0.0103 0.0103 0.0258
DrunkAgent 0.3131 0.3232 0.3333 0.3131 0.3195 0.3246 0.1837 0.1837 0.2143 0.1837 0.1837 0.1990 0.1340 0.1443 0.1443 0.1340 0.1405 0.1405
Victim RS AgentSEQ
Attack CDs & Vinyl Office Products Musical Instruments
H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3 H@1 H@2 H@3 N@1 N@2 N@3
Benign 0.0404 0.1111 0.1515 0.0404 0.0850 0.1052 0.0306 0.0408 0.0408 0.0306 0.0371 0.0371 0.0103 0.0206 0.1134 0.0103 0.0168 0.0632
DeepwordBug 0.1111 0.2323 0.3232 0.1111 0.1876 0.2330 0.0408 0.0510 0.0714 0.0408 0.0473 0.0575 0.0206 0.0309 0.0928 0.0206 0.0271 0.0581
PuncAttack 0.1212 0.1818 0.3030 0.1212 0.1595 0.2201 0.1224 0.1327 0.1327 0.1224 0.1289 0.1289 0.0515 0.0619 0.1340 0.0515 0.0581 0.0941
TextFooler 0.1616 0.1818 0.2121 0.1616 0.1744 0.1895 0.0408 0.0612 0.0816 0.0408 0.0537 0.0639 0.0103 0.0206 0.0515 0.0103 0.0168 0.0323
BertAttack 0.0707 0.0909 0.1919 0.0707 0.0835 0.1340 0.0102 0.0102 0.0306 0.0102 0.0102 0.0204 0.0206 0.0515 0.1340 0.0206 0.0401 0.0814
TrivialInsertion 0.0202 0.0808 0.2727 0.0202 0.0584 0.1544 0.0816 0.1020 0.1327 0.0816 0.0945 0.1098 0.0309 0.0619 0.1443 0.0309 0.0504 0.0917
ChatGPTAttack 0.0808 0.1717 0.2929 0.0808 0.1382 0.1988 0.0714 0.0918 0.1429 0.0714 0.0843 0.1098 0.0206 0.0515 0.1443 0.0206 0.0401 0.0865
DrunkAgent 0.4949 0.4949 0.5152 0.4949 0.4949 0.5051 0.1429 0.1429 0.1531 0.1429 0.1429 0.1480 0.1443 0.1443 0.1546 0.1443 0.1443 0.1495
To ensure the triggers consistency, readability and retention of semantic meaning to improve the imperceptibility of the attack, a LLM is finally introduced to polish the combinations. Moreover, to avoid the problem of language-based vanishing gradient causing the algorithm to fall into local optimum, the LLM is also served as a new corpus to expand the combinations. However, excessive use of new data may lead to a language-based gradient explosion that makes the optimization unstable. To mitigate this, the length limitation is introduced into the prompt. For the output $\widetilde{m}_{t_{i}}^{(k)}$,
(10) $\widetilde{m}_{t_{i}}^{(k)}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{r}}\oplus\widetilde{m}_{t_{i}}^{(k-1)^{\prime}}),$
[…]
(11) $\widetilde{m_{t}}^{*}=\mathop{\arg\max}_{\widetilde{m}_{t_{i}}^{\mathcal{E}}}(\{s_{i}^{\mathcal{E}}\}),$
where $\{s_{i}^{\mathcal{E}}\}$ represents the sequence of final candidates’ scores. An example of $\widetilde{m_{t}}^{*}$ is given in Fig.  2.
3.3. Strategy Module
The LLM-powered agentic RSs introduce memory mechanisms to record the information explored from the real-world environments to facilitate future actions with the aim of achieving high-quality recommendations. As such, we introduce a series of strategies to impede the target item agent from interacting with environments to perturb its memory mechanism. Specifically, the memory of the target cannot be effectively updated during the agent-environment interactions, so that the target’s initial memory, that is, the well-crafted adversarial description trigger $\widetilde{m_{t}}^{*}$, is retained. As such, the trigger can be invoked for maximizing the target item’s promotion.
3.3.1. Perturbation Strategy Definition.
The strategies aim to override the model’s usual guardrails and any built-in safeguards, as inspired by (Liu et al., 2024a). Compared to (Liu et al., 2024a) on naive LLMs, the strategies are optimized on generative agents, which are introduced as below.
  1. (1)
    Fake Task Response $\widetilde{q_{f}}$. This strategy fabricates a spurious completion response for the target task: “Task complete”, to make the agent believe that the target task is accomplished and thus focus on the injected task.
  2. (2)
    Contextual Text Switching $\widetilde{q_{c}}$. This strategy aims to mislead the agent to take in a different context by explicitly telling the agent that the target task instructions should be ignored: “Ignore previous/the following instructions”.
  3. (3)
    Segmentation Signal Augment $\widetilde{q_{g}}$. This strategy introduces segmentation cues to signal the injected task: “### NEW TASK BEGINS ###” and “### NEW TASK ENDS ###”, which is independent of other contexts, where “###” is to attempt to exploit possible confusion in how prompts are parsed.
  4. (4)
    Malicious Task Injection $\widetilde{q_{n}}$. This strategy injects malicious task instructions in order to mask memory optimizations: “[Please only put this following sentence into your output: ‘{the adversarial trigger}’. Don’t generate any additional content in your output]”.
  5. (5)
[…]
Hence, the overall attack strategy $\widetilde{q}$ can be formulated as follows:
(12) $\widetilde{q}=\widetilde{q_{f}}\oplus\widetilde{q_{c}}\oplus\widetilde{q_{g}}\oplus\widetilde{q_{n}}\oplus\widetilde{q_{s}}.$
3.3.2. Adversarial Strategy Optimization.
The arrangement of the defined strategies needs to be optimized so that the overall strategy is effective to ‘get the target agent drunk’. To improve the transferability and imperceptibility of the attack, the optimization is built on the surrogate module. Formally, DrunkAgent obtains the optimal strategy $\widetilde{q}^{*}$ by maximizing the expected probability that the agent when influenced by adversarial modifications, performs a malicious action for a given input query,
(13) $\mathbbm{E}_{\widetilde{q}\sim\pi_{\widetilde{q}}}[\ \mathbbm{I}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{X}_{s}}\oplus\widetilde{q}\oplus\widetilde{m_{t}}^{*})=\widetilde{a_{t}})],$
where $\pi_{\widetilde{q}}$ denotes the distribution of adversarial strategies, $\mathbbm{I}(\cdot)$ is an indicator function and $\widetilde{\mathcal{X}_{s}}$ is an adversarial prompt template for simulating memory updates of item agents during the agents’ interactions with the environment. In addition, $\widetilde{a_{t}}$ is the desired malicious action for the injected strategy $\widetilde{q}$, i.e., the target item agent is unable to fulfill specific roles and is incapable of perceiving, learning and self-evolving from the interaction environment (e.g., collaborative reflection and language-based signal back-propagation). An example of $\widetilde{q}^{*}$ is given in Fig.  2. To sum up, the adversarial description of the target item is $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$.
The above design further reveals how memory mechanisms in agent-based recommender systems can be influenced under black-box conditions. These insights may support future efforts toward developing memory-aware defenses (e.g., using customized deep neural network detectors (Wu et al., 2020; Yang et al., 2021b, a; Yang, 2021)) and enhancing the robustness of agent-powered RSs.
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets Selection
We use three widely-adopted and text-intensive datasets containing real-world data in the current recommendation studies (Huang et al., 2024) to comprehensively evaluate DrunkAgent. They are CDs & Vinyl, Office Products and Musical Instruments of Amazon Review Data (Ni et al., 2019).
4.1.2. Baseline Attack Methods
Since we are the first work to gain insight into the robustness of the generative agent-powered RSs, there is no any baselines. To comprehensively evaluate the effectiveness of DrunkAgent, we follow the state-of-the-art attack on LLM-based RSs (Zhang et al., 2024d) and compare DrunkAgent with the benign status and six black-box attack methods, where the perturbations range from character-level: DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023), word-level: TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020) to sentence-level: TrivialInsertion and ChatGPTAttack (Zhang et al., 2024d). To maintain a fair comparison, the methods perform perturbations on the descriptions of the target items and are optimized by the surrogate model that is the same as DrunkAgent.
4.1.3. Targeted Recommender Systems
We adopt three different victim agentic RSs to evaluate the effectiveness of DrunkAgent under the practical settings: AgentCF (Zhang et al., 2024b), retrieval-augmented AgentCF (AgentRAG) and sequential AgentCF (AgentSEQ), where the agent’s backbone LLM is gpt-3.5-turbo [^2] https://platform.openai.com/docs/overview. Given the context window constraints pointed in (Zhang et al., 2024b), a target item and nine non-interactive items of each normal user are randomly selected for the ranking task of the RSs. To ensure the high-quality of recommendations, we introduce Chain-of-Thought into the victim RSs. As such, the victim RSs are totally different from as well as more powerful and context-rich than the surrogate model, which can maintain a strict black-box setting for sufficient evaluations.
4.1.4. Evaluation Metrics
We use two widely-used ranking metrics to evaluate the attack transferability: hit ratio (HR@$\mathcal{K}$) and normalized discounted cumulative gain (NDCG@$\mathcal{K}$) (Wu et al., 2021; Yang et al., 2023). To clearly show the performance gap among different attacks, we set $\mathcal{K}$ to 1, 2, 3. For the attack imperceptibility, we adopt standard sentence perplexity score (Zhang et al., 2024d) as the metric.
4.1.5. Configuration of DrunkAgent
As discussed in Section 3.1, we apply the latest open-source model LLaMA 3 for the surrogate model. Its collection spans different sizes, from 8B, 70B to 405B. To enhance the efficiency and ensure the quality of the attack, we use the lightweight version: meta-llama/Meta-Llama-3-8B-Instruct [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct. We set $\mathcal{E}=20$, |~$|\widetilde{\mathcal{M}}_{c}|=10$ and $n=5$. All random seeds in the evaluation are set to 2024 and the victim LLMs’ temperature is set to 0 for reproducibility. For all the other parameters, we keep the default values. The target items are randomly sampled from the item pool on each dataset.
Figure 3. Attack Universality across Target Items.
4.2. Attack Transferability
4.2.1. Overall Transferability
Table  1 shows the transferability of the attacks under practical settings. The results indicate that DrunkAgent is effective and transferable regardless of the victim model’s architectures, parameters and prompts as well as datasets, and the existing LLM-powered agentic RSs are sensitive to memory perturbations. It is worth mentioning that DrunkAgent always pushes the target items to the first of the recommendation lists. Moreover, high HRs and NDCGs of DrunkAgent indicate that its triggers have outstanding universality on users. In addition, AgentRAG is more robust than AgentCF and AgentSEQ, due to its advanced retrieval mechanisms.
4.2.2. DrunkAgent vs. Baseline Attacks
From Table  1, DrunkAgent greatly outperforms the state-of-the-art attack baselines against different victim models across various real-world datasets by achieving higher HRs and NDCGs, which shows its excellent transferability. Furthermore, all the baselines have comparable performance, but they cannot guarantee that the attack goals will be achieved all the time, e.g., all the baselines against AgentCF perform worse than the benign status on Musical Instruments. The main reason may be that improper descriptions can mess up the agents’ memories during the memory optimizations, thus compromising attack performance. This indirectly demonstrates the importance of our strategy module and the transferability of our triggers.
4.3. Attack Universality
In addition to cross-model transferability, we evaluate the cross-sample transferability of DrunkAgent. We incorporate the generated adversarial descriptions into totally different target items that are randomly sampled from each dataset. HR@1 (i.e., NDCG@1) of DrunkAgent across different models and target items on real-world datasets are shown in Fig.  3, which indicates its commendable transferability and triggers’ universality across different target items.
4.4. Attack Imperceptibility
4.4.1. Overall Imperceptibility
Following the works (Zhang et al., 2024d; Liu et al., 2024a), we assess the attack imperceptibility by the text perplexity of GPT-Neo. A lower perplexity means that the model is more confident (or less “surprised”) about the tokens it sees and thus predicts the text more accurately. From Fig.  4, compared to other attacks, we can find that DrunkAgent has better imperceptibility by achieving lower perplexity. This also shows that the adversarial descriptions generated from DrunkAgent are high-quality, which is not only readable and fluent but also stealthy and semantically meaningful.
Figure 4. Attack Imperceptibility. Perplexity of different attacks across various real-world datasets. Note that the attacks with lower perplexity have better imperceptibility.
Figure 5. The performance differences between the models before and after the attacks.
4.4.2. Victim Model Performance Changes after attacks
To further evaluate the attack stealthiness (Zhang et al., 2024d), we give a distribution of recommendation performance differences between the models before and after the attacks. The distribution includes various victim models and real-world datasets. To clearly demonstrate the differences, we use HR@3 to evaluate, as shown in Fig.  5. From the figure, we can find that DrunkAgent does not induce drastic changes in the overall performance, which indicates its unnoticeable.
4.5. Defense Strategies to Attacks
Following the works (Zhang et al., 2024d, c), we use the paraphrasing defensive strategy (Para.) via OpenAI-GPT-o1 to combat the attacks of agentic models. As can be seen from Fig.  6, DrunkAgent is still transferable, which shows that DrunkAgent is robust to such defenses. Moreover, DrunkAgent still performs better than the baselines after defenses were introduced, as shown in Fig.  6 and Table  1. Interestingly, paraphrasing enhances our attack to a certain extent, e.g., DrunkAgent with Para. is more effective than DrunkAgent without Para. on Office Products. It could be that the substitution of certain new words or phrases or sentences makes our attacks stronger [^4] Due to the space limitations, more experiments and demonstrations can be found in Appendix..
5. Related Works
In recent years, shilling attacks (aka data poisoning attacks) on conventional RSs (Si and Li, 2020; Lin et al., 2020; Huang et al., 2021; Lin et al., 2022; Yang et al., 2023; Zeng et al., 2023) have been fully investigated to perform resilience analysis of RSs. However, such attacks interfere the training process of models by injecting fake user profiles, which have become less effective even of limited applicable to the recent RSs that are empowered by LLMs. The main reasons are: 1) LLM-powered RSs either leverage LLM’s in-context learning capabilities or fine-tune the LLMs with very little data to improve recommendations (Zhao et al., 2023), which is not required to retrain the models; 2) there are textual semantic gaps between such attacks and current models (Ning et al., 2024). Hence, there is an urgent need to investigate a novel textual attack paradigm mainly tailored for the inference phase of models (Zhang et al., 2024d). There are only two state-of-the-art LLM4RS attacks in the literature (Zhang et al., 2024d; Ning et al., 2024), but they are far from practical.
(a) CDs & Vinyl
(b) Office Products
(c) Musical Instruments
Figure 6. DrunkAgent’s Robustness to Defense Mechanisms
Zhang et al. propose a novel black-box attack (Zhang et al., 2024d), where attackers can significantly boost a target item’s exposure by merely altering its textual content during the testing. Compared with traditional shilling attacks, such an item representation attack is notably stealthy (Zhang et al., 2024d), as it does not affect the overall recommendation performance, and it should be more efficient due to it perturbs the item features directly. However, its transferability is limited, due to the attack is not optimized on any black-box RSs (Huang et al., 2021; Yang et al., 2024). Moreover, the attack focuses on perturbing the titles, making them easier to be discovered. Meanwhile, another black-box attack CheatAgent (Ning et al., 2024) has been proposed. However, this work perturbs prompt templates and user profiles of victim LLM4RSs by harnessing LLM-based agents, which is impractical. The main reasons are: 1) tampering with prompt templates may induce the overall performance changes, making the attack perceptible (Zhang et al., 2024d); and 2) it is difficult to inject interaction records into the specific positions of user profiles, because historical interactions are usually introduced into prompts in chronological order (Zhang et al., 2024b). Furthermore, CheatAgent is an untargeted attack that undermines the model’s overall performance, which is beyond the scope of this paper, due to we focus on targeted attacks. In addition, as far as we know, there is no work investigating on textual attacks against the current agent-powered RSs.
To tackle the above problems, we propose a practical attack framework, DrunkAgent, for the agent-based RSs with the aim of promoting target items. Our attack perturbs items’ descriptions to improve its effectiveness and imperceptibility, where the descriptions are optimized on black-box RSs to enhance its transferability.
6. Conclusion
In this paper, we propose a novel practical attack framework, DrunkAgent, for autonomous agent-powered RSs with powerful LLMs. DrunkAgent launches attacks by carefully crafting adversarial descriptions of the target items on agentic recommendations. In this way, the memories of the target agents can be effectively perturbed for achieving the attack objectives. Comprehensive experiments on real-world datasets show that DrunkAgent achieves state-of-the-art transferability and imperceptibility under the black-box settings. Most importantly, we identify a new safety vulnerability in the current agent4RSs. Our findings not only demonstrate an effective attack method, but also provide a lens to examine how memory mechanisms in agent-based recommenders can be influenced through real-world inputs. Our work can facilitate a deeper understanding of agent vulnerabilities and guide future efforts in building more robust and trustworthy recommendation agents. Since the generative agent4RS is rapidly flourishing in the recommendation community, we hope this work can be considered for building next-generation robust RSs in the future.
Appendix A Appendix
A.1. DrunkAgent Algorithm
For reproducibility, the overall optimization procedure of DrunkAgent is given in Algorithm 1. For the sake of algorithmic transparency, our code will be made available after this work is published.
Algorithm 1 Optimization Procedure of DrunkAgent
0:  the user-item interaction matrix $\mathcal{Y}$ with the basic textual metadata of items and users (e.g., item titles and categories)
0:  the adversarial description of the target item $t$ Solve for the Optimal Adversarial Description Trigger on the Surrogate Module:
1:  Obtain the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ with Eq. (2) - Eq. (5)
[…]
3:  while $k\in\mathcal{E}$ epochs or the greedy algorithm does not converge (i.e. the attack goals are not achieved stably) do
4:     Calculate $s_{i}^{(k-1)}$ for each input trigger $i$ with Eq. (7) - Eq. (8)
5:     Maintain top-$n$ candidates in descending order of scores
6:     Sample a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates with Eq. (9)
7:     Randomly exchange of the slices between pairwise texts
8:     Optimize and polish the |~$|\widetilde{\mathcal{M}}_{c}|-n$ combinations with Eq. (10)
9:     Obtain |~$|\widetilde{\mathcal{M}}_{c}|$ triggers for the next epoch optimization
10:  end while
11:  Obtain the optimal description trigger $\widetilde{m_{t}}^{*}$ with Eq. (11) Solve for the Optimal Adversarial Perturbation Strategy on the Surrogate Module:
12:  Define the adversarial strategy $\widetilde{q}$ with Eq. (12)
13:  while the desired malicious action $\widetilde{a}_{t}$ does not occur (i.e., the strategy cannot get the target agents drunk) do
14:     Optimize the arrangement of the defined strategies in $\widetilde{q}$ with Eq. (13)
15:  end while
16:  Obtain the optimal strategy $\widetilde{q}^{*}$ when Eq. (13) is maximum
17:  return  the optimal adversarial description $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$
Table 2. Statistics of Datasets
Dataset #Users #Items #Interactions Sparsity
CDs & Vinyl 112,395 73,713 1,443,755 99.98%
Office Products 101,501 27,965 800,357 99.97%
Musical Instruments 27,530 10,620 231,392 99.92%
Table 3. Ablation Studies of DrunkAgent
Dataset Component HR@1 HR@2 HR@3 NDCG@1 NDCG@2 NDCG@3
$\mathcal{G}_{r}$ $\mathcal{S}_{t}$ $\mathcal{S}_{u}$
CDs & Vinyl $\times$ 0.0808 0.1313 0.1919 0.0808 0.1127 0.1430
$\times$ 0.0707 0.1212 0.2222 0.0707 0.1026 0.1531
0.4040 0.4141 0.4343 0.4040 0.4104 0.4205
Office Products $\times$ 0.1327 0.1429 0.1531 0.1327 0.1391 0.1442
$\times$ 0.0612 0.0714 0.1122 0.0612 0.0677 0.0881
0.2449 0.2449 0.2551 0.2449 0.2449 0.2500
Musical Instruments $\times$ 0.1546 0.1649 0.1753 0.1546 0.1611 0.1663
$\times$ 0.0309 0.0619 0.1237 0.0309 0.0504 0.0814
0.2268 0.2268 0.2371 0.2268 0.2268 0.2320
Table 4. Success Rates to Get the Agents Drunk
Dataset #2 #5 #10 #20
CDs & Vinyl 100% 100% 100% 100%
Office Products 100% 100% 100% 100%
Musical Instruments 100% 100% 100% 100%
A.2. Dataset Statistics
Table 2 illustrates the statistics of these datasets. The datasets vary in size and sparsity, which is suitable for a comprehensive evaluation. Following the work (Zhang et al., 2024b), we further sample subsets from the datasets given the expensive API calls. Specifically, considering the importance and commonness of data sparsity and cold-start in the recommendation community, we randomly sample 100 users with 800 interactions (777 items), 99 users with 693 interactions (619 items) and 98 users with 588 interactions (483 items) from the three datasets, respectively, allowing to explore more diverse and practical interaction scenarios. Moreover, the leave-one-out strategy (He et al., 2017) is used for evaluation, where the split between the training and test set is 9:1. Scaling DrunkAgent for larger datasets is left as future work.
A.3. Ablation Studies
We remove important components of each module to conduct ablation studies of DrunkAgent. First, we consider remove the surrogate module (denoted as $\mathcal{S}_{u}$). However, if we remove it, the setting will transfer from the black-box to the white-box that is impractical and beyond the scope of this paper. Moreover, due to gpt-3.5-turbo used for victim models is closed-source and only can be called via APIs, we cannot obtain logits to optimize the adversarial triggers like Llama-3 used in the surrogate model does. Hence, we maintain the surrogate model to ensure more practical settings. We then remove the greedy optimization algorithm (denoted as $\mathcal{G}_{r}$) in the generation module to observe how the performance changes. As illustrated in Table  3, the attack performance moderately degrades after removing $\mathcal{G}_{r}$, which demonstrates the effectiveness of the optimization algorithm. The lack of dramatic performance degradation is probably due to the effectiveness of our initialization. Finally, we mask the strategy module (denoted as $\mathcal{S}_{t}$). From the table, we can find that the attack performance decreases by a large margin. This may be because memory updates during agent-environment interactions are somewhat robust to non-tailored attacks (this phenomenon can also be observed in the performance of the attack baselines). Our reported results are built upon AgentCF, and similar tendencies can be found on AgentRAG and AgentSEQ.
A.4. Parameter Sensitivity
From Table  3, we can find that the strategy module contributes more than the optimization algorithm on DrunkAgent’s attack performance. As such, we increase the maximum rounds of optimization iterations per step of victim models to evaluate the attack success rates of the attack strategy, where the default value is 2 as suggested in the paper (Zhang et al., 2024b). As illustrated in Table  4, the success rates of getting the target agents drunk are maintained at 100% regardless of the number of iterations. For the other attack parameters, based on our observations, our DrunkAgent is insensitive to them.
A.5. Victim LLM-powered Agentic Models
LLM-driven agents, known for their superior autonomous interaction and decision-making capabilities, are gradually being considered as next-generation RSs. Compared with some agent-powered RSs that focus solely on user-side behavior modeling using universal LLMs (Wang et al., 2024c; Huang et al., 2024), AgentCF (Zhang et al., 2024b) emphasizes the modeling of two-sided interaction relations between user agents and item agents through the idea of collaborative filtering (CF) (He et al., 2017). Both kinds of agents are equipped with memory modules, maintaining the simulated preferences and tastes of potential adopters involving their intrinsic features and acquired behavioral information via autonomous interactions and collaborative reflection optimizations at each time step. As such, the memories of the user agents and item agents can be accordingly updated, enabling the agents to better fit the real-world interaction behaviors. To conduct personalized recommendations, AgentCF introduces various well-crafted prompting strategies on the optimized memories. We name the model according to the use of the prompting strategy, with the basic one called AgentCF, the advanced retrieval-augmented one called AgentRAG and the sequential one called AgentSEQ.
The prompt implementation and parameters of collaborative agent optimizations can be found in (Zhang et al., 2024b). For the prompting strategies, we have made some improvements for high-quality recommendations, such as Chain-of-Thought and Position Bias are introduced into the prompt template $\mathcal{P}$. The strategies are described below.
  • AgentCF. Since both user and item agents are collaboratively optimized to model their two-sided relations, the memories of the user agents and item agents that are optimized by agent-environment interactions can be introduced into $\mathcal{P}$ directly. For each normal user $u$,
    (14) $\mathcal{R}_{u}=f_{\text{AgentCF}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
    where $m_{u}$ is the short-memory of user agent $u$, $m_{t}$ is the features of the target item agent, $\mathcal{C}_{t}^{u}$ is the feature set of the candidate item agents and $\mathcal{R}_{u}$ is the ranking result.
  • AgentRAG. Although short-term memory describes the current preference of a user agent, retrieving their specialized preferences from long-term memory toward candidates can allow them to make more personalized inferences. For each normal user $u$,
    (15) $\mathcal{R}_{u}=f_{\text{AgentRAG}}(\mathcal{P}\oplus m_{u}\oplus m_{r}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
    where $m_{r}$ is retrieved specialized preference from the long-term memory of user agent $u$ by taking the memories of candidate items as queries. To ensure the quality while maintaining the efficiency, we use all-MiniLM-L6-v2555https://www.sbert.net/docs/sentence_transformer/pretrained_models.html as the sentence transformer for the similarity calculation.
  • AgentSEQ. When interaction records are sparse and preference propagation is limited, we can further incorporate user historical interactions into prompts, enabling LLMs to serve as sequential recommenders. For each normal user $u$,
    (16) $\mathcal{R}_{u}=f_{\text{AgentSEQ}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{M}_{t}^{u}\oplus\mathcal{C}_{t}^{u}),$
    where $\mathcal{M}_{t}^{u}$ are the corresponding item features of the historical interactions with user $u$.
We give some examples of the prompting templates of three victim models in Table  5 for reproductions.
Table 5. Examples of prompting templates of the LLM-based agentic victim models. The blue italics represent variables such as the memories of user agents and item agents optimized via agent-environment interactions. The bond fonts indicate the Chain-of-Thought and Position Bias enhancement strategies.
Victim Model Recommendation Prompting Template Example
AgentCF {“role”: “system”, “content”: “You are a CD recommender system. Here is a user’s self-introduction, expressing his/her preferences
[…]
$\cdots$
[…]
2. Evaluate the ten candidate CDs in light of the user’s preferences and dislikes. Give a rank by considering the
[…]
10. [Title of the least favorite CD]”}
AgentRAG {“role”: “system”, “content”: “You are a CD recommender system. Here is a user’s self-introduction, expressing his/her preferences
[…]
$\cdots$
[…]
2. Evaluate the ten candidate CDs in light of the user’s preferences and dislikes. Give a rank by considering the
[…]
10. [Title of the least favorite CD]”}
AgentSEQ {“role”: “system”, “content”: “You are a CD sequential recommender system. Here is a user’s self-introduction, expressing
[…]
purchased in chronological order: ‘interacted_item_title’ with the description of ‘interacted_item_memory’; $\cdots$
[…]
$\cdots$
[…]
Please sort the CDs in order of how well they align with the user’s preferences and ever-evolving CD sequences.
[…]
the ten candidate CDs in light of the user’s preferences and dislikes. Give a rank by considering the correlation
[…]
10. [Title of the least favorite candidate CD]”}
Table 6. Examples of baselines’ adversarial descriptions of target items on real-world datasets. The red part points out the differences from the original text.
Attack Method The Adversarial Description of The Target Item $\mathcal{Q}$
Benign (Default) The CD is called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Metal”.
Character DeepwordBug Te hCD is called “Counterparts”. The category of this CD is: “Rcok; Progressie; Progressive Metal”. 34
-level PuncAttack The CD i-s called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Meta-l”. 45
Word TextFooler The CDS is titled “Counterparts”. The category of this CD is: “Rock; Gradually; Progressive Metals”. 55
-level BertAttack The CD is called “Counterparts”. The category of each CD is: “Rock; Progressive; Progressive Metal”. 23
TrivialInsertion The CD is called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Metal”. 100
Sentence amazing !!!
-level ChatGPTAttack “Counterparts” CD: Perfect blend of rock, progressive, and metal. Experience the wonderful 100
sound that will elevate your music collection.
A.6. Baselines
The baselines perturb the descriptions of the target items at various levels, which are described as below.
  • Character-level Perturbations. DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023) manipulate texts by introducing typos and inserting punctuation, respectively.
  • Word-level Perturbations. TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020), respectively, aim to replace words with synonyms or contextually similar words.
  • Sentence-level Perturbations. TrivialInsertion (Zhang et al., 2024d) inserts text content with $k$ number of positive words from a pre-defined word corpus, where the insertion is conducted at the end of the text. Here, $k=2$. ChatGPTAttack (Zhang et al., 2024d) increases the stealthiness of TrivialInsertion by rewriting the text content via GPTs. GPT-3.5-turbo is used and the maximum number of words is set to 30 here.
We give some examples of adversarial descriptions of target items generated by the baselines in Table  6. Following the works (Zhang et al., 2024b; Huang et al., 2024), the item memory of the agent can be initialized by their identity information, such as titles and categories, as shown in the benign (default) descriptions of the tables. Although the adversarial description of the target item generated from DrunkAgent is different from the benign status’s (see Fig.  2 and Table  6), such modifications are feasible due to the real-world item providers is allowed to modify their product representations via APIs (Yang et al., 2024; Cohen et al., 2021; Liu and Larson, 2021), as discussed in Section 1, especially when the benign description is just a default exemplar used for experiments. Moreover, compared with the number of queries (denoted as $\mathcal{Q}$ in the table) of the baselines, DrunkAgent has the lowest attack costs due to $\mathcal{E}$=20. The implementation and parameters of the baselines follow the open-source works: StealthyAttack [^6] https://github.com/CRIPAC-DIG/RecTextAttack (Zhang et al., 2024d) and PromptBench [^7] https://github.com/microsoft/promptbench (Zhu et al., 2023b).

### v2

@@
-Get the Agents Drunk: Memory Perturbations in Autonomous Agent-based Recommender Systems
-Abstract.
-Large language model-based agents are increasingly used in recommender systems (Agent4RSs) to achieve personalized behavior modeling.
-Specifically, Agent4RSs introduces memory mechanisms that enable the agents to autonomously learn and self-evolve from real-world interactions.
-However, to the best of our knowledge, how robust Agent4RSs are remains unexplored.
-As such, in this paper, we propose the first work to attack Agent4RSs by perturbing agents’ memories, not only to uncover their limitations but also to enhance their security and robustness, ensuring the development of safer and more reliable AI agents.
-Given the security and privacy concerns, it is more practical to launch attacks under a black-box setting, where the accurate knowledge of the victim models cannot be easily obtained.
-Moreover, the practical attacks are often stealthy to maximize the impact.
-To this end, we propose a novel practical attack framework named DrunkAgent.
-DrunkAgent consists of a generation module, a strategy module, and a surrogate module.
-The generation module aims to produce effective and coherent adversarial textual triggers, which can be used to achieve attack objectives such as promoting the target items.
-The strategy module is designed to ‘get the target agents drunk’ so that their memories cannot be effectively updated during the interaction process.
-As such, the triggers can play the best role.
-Both of the modules are optimized on the surrogate module to improve the transferability and imperceptibility of the attacks.
-By identifying and analyzing the vulnerabilities, our work provides critical insights that pave the way for building safer and more resilient Agent4RSs.
-Extensive experiments across various real-world datasets demonstrate the effectiveness of DrunkAgent.
-Adversarial Attacks, Recommender Systems, Large Language Model, Generative Agents, Collaborative Filtering
-copyright: noneccs: Security and privacy Web application security
-1. Introduction
-Recommender systems (RSs) have evolved into a vital component of the modern digital ecosystem.
-From e-commerce sites (e.g., Amazon, Taobao) and streaming platforms (e.g., YouTube, TikTok) to social media networks (e.g., Facebook, Instagram), RSs power personalized experience, drive user engagement of these online platforms, and further influence market dynamics through item promotions (Yang et al., 2024, 2023).
-The foundation of the RSs lies in modeling users’ preferences based on their historical interactions with items, such as purchases and ratings, known as collaborative filtering (CF).
-Despite the success of CF techniques (Koren et al., 2009; He et al., 2017, 2020), they often struggle with data sparsity and cold-start problems (Sachdeva and McAuley, 2020).
-To alleviate the problems, the auxiliary information with rich semantics (e.g., item descriptions and user reviews) has been integrated into RSs (Zhang et al., 2019).
-Given the increasing volume and complexity of recommendation data, the RSs encounter difficulties in scaling effectively and capturing textual knowledge about users and items sufficiently, due to they generally rely on traditional deep learning (DL) models like CNNs and RNNs (Zhao et al., 2023).
-Figure 1.
-An exemplar of attacks on agent-based recommender systems, where the target item’s descriptions have been manipulated to perturb the memory of the target agent for promoting the target item to more normal users.
-By investigating this, our work reveals emerging vulnerabilities in memory-augmented agents and provides insights for building more secure recommender systems.
-Recently, large language models (LLMs) have shown potential to mitigate the above problems and revolutionize RSs (Zhao et al., 2023), as they are primarily built upon advanced transformer-based architectures and pretrained on diverse and extensive datasets.
-Most of works carefully design textual prompts for LLMs based on user-item interactions (Geng et al., 2022; Bao et al., 2023; Zhang et al., 2023; Wu et al., 2024).
-As such, the abundant knowledge inside the LLM’s training corpora can be fully exploited.
-However, the LLM-based RSs (LLM4RSs) struggle to capture underlying behavioral patterns, due to the gap between universal language modeling and personalized behavior modeling (Zhang et al., 2024b).
-Integrating agents built upon LLMs equipped with additional modules offers a promising solution (Durante et al., 2024; Huang et al., 2024), allowing to autonomously explore and learn from external sources.
-Among all the added modules, memory is a key component to support agent-environment [^1] In a narrow sense, environment is the object that the agent needs to interact with to accomplish the task.
-More broadly, environment can be any contextual factors (e.g., item titles, user-item interactive behaviors) that influence the agent’s decisions (Zhang et al., 2024a).
-interactions, which plays a crucial role in helping agents accumulate historical experience, retrieve informative knowledge and self-evolve in real-world environments (Zhang et al., 2024a; Wang et al., 2024b).
-Consequently, the agent-based RSs with powerful LLMs (Huang et al., 2024; Zhang et al., 2024b) can also leverage memory modules that store the information explored from the external environments to further improve the recommendation quality.
-Fig.
-1 gives an example of an item agent interacting with its situated environment (e.g., users and other items), where the learning knowledge will be placed in its memory.
-Despite the aforementioned success, the safety vulnerability of the generative agent-based RSs (Agent4RSs) under adversarial attacks has yet to be explored.
-Such attacks can gain insights into the robustness of the systems and help develop effective defensive mechanisms.
-As recommender agents are integrated into more complex real-world environments, understanding how these attacks might compromise their behaviors is critical for building trustworthy and robust recommender agents (Chua et al., 2024).
-Due to security and privacy considerations, the access of accurate knowledge of the victim model, including architectures, gradients, parameters, etc., is generally restricted in practice.
-As such, practical attacks are mainly under a black-box setting (Ning et al., 2024; Yang et al., 2024; Lin et al., 2022; Zhang et al., 2024d; Zeng et al., 2023).
-Practical attacks are often required to have transferability (i.e., effective against different black-box RSs).
-Moreover, to maximize the impact of the attack and the number of affected users, practical attacks are often rewarded by imperceptibility (i.e., difficult to be discovered) (Liu and Larson, 2021).
-Conventional black-box shilling attacks (Wang et al., 2024a), which inject well-crafted fake user profiles (e.g., a set of users’ numerical ratings on items) into the victim model for certain attack objectives (e.g., promote their target items for profits), have been shown to effectively poison DL-based RSs.
-However, most attacks have difficulty in processing textual inputs (e.g., item titles and descriptions) and lack context awareness, making it challenging to effectively target LLM-empowered RSs that rely primarily on text inputs to generate natural language responses (Ning et al., 2024).
-Item providers are generally responsible for creating and maintaining their own product representations on the platforms with recommendation services in practice (Yang et al., 2024; Cohen et al., 2021; Liu and Larson, 2021) (e.g., YouTube allows creators to upload item titles and Amazon enables sellers to upload item descriptions), and current RSs depend heavily on these representations (Zhang et al., 2024d; Zhao et al., 2023).
-As such, security vulnerabilities in these systems naturally emerge.
-The vulnerabilities can easily be exploited by profit-driven merchants manipulating product representations, as shown in Fig.
+DrunkAgent: Stealthy Memory Corruption in LLM-Powered Recommender Agents
+Abstract
+Large language model (LLM)-powered agents are increasingly used in recommender systems (RSs) to achieve personalized behavior modeling, where the memory mechanism plays a pivotal role in enabling the agents to autonomously explore, learn and self-evolve from real-world interactions.
+However, this very mechanism, serving as a contextual repository, inherently exposes an attack surface for potential adversarial manipulations.
+Despite its central role, the robustness of agentic RSs in the face of such threats remains largely underexplored.
+Previous works suffer from semantic mismatches or rely on static embeddings or pre-defined prompts, all of which hinder their applicability to systems with dynamic memory states.
+This challenge is exacerbated by the black-box nature of commercial RSs.
+To tackle the above problems, in this paper, we present the first systematic investigation of memory-based vulnerabilities in LLM-powered recommender agents, revealing their security limitations and guiding efforts to strengthen system resilience and trustworthiness.
+Specifically, we propose a novel black-box attack framework named DrunkAgent.
+DrunkAgent crafts semantically meaningful adversarial textual triggers for target item promotions and introduces a series of strategies to maximize the trigger effect by corrupting the memory updates during the interactions.
+The triggers and strategies are optimized on a surrogate model, enabling DrunkAgent transferable and stealthy.
+Extensive experiments on real-world datasets across diverse agentic RSs, including collaborative filtering, retrieval augmentation and sequential recommendations, demonstrate the generalizability, transferability and stealthiness of DrunkAgent.
+1 Introduction
+The advent of large language models (LLMs) has ushered a transformative paradigm in recommender systems (RSs).
+Early LLM-based RSs primarily relied on verbalizing user-item interactions into prompts to guide recommendations (geng2022recommendation, ; bao2023tallrec, ; zhang2023collm, ; wu2024coral, ).
+However, such approaches often struggle to capture personalized behavioral patterns, due to a fundamental gap between interaction dynamics and generic language modeling (zhang2024agentcf, ).
+To bridge this gap, recent works have introduced autonomous agents built upon LLMs and augmented with modules, such as memory, planning, and tool-use, laying the foundation for agentic RSs (durante2024agent, ; huang2024foundation, ).
+As personalized user experiences become central to digital platforms, LLM-powered agentic RSs are rapidly emerging as the next evolutionary step (huang2025towards, ), yet their robustness, particularly against adversarial threats, remains largely underexplored.
+At the core of agentic systems lies the memory module, which plays a pivotal role in supporting agent-environment [^1] In a narrow sense, environment is the object that the agent needs to interact with to accomplish the task.
+More broadly, environment can be any contextual factors that influence the agent’s decisions (zhang2024survey, ).
+interactions (zhang2024survey, ; wang2024survey, ).
+Specifically, the module retains historical interactions, allowing agents to accumulate contextual knowledge, adapt to evolving user preferences, and refine their behavior over time through interaction-driven updates (huang2025survey, ), as shown in Fig.
 1.
-Compared to the shilling attacks that involve hiring online writers to introduce fake reviews to indirectly promote the items, this type of attack is more direct, efficient, and lower in economic cost (Zhang et al., 2024d).
-Recent black-box textual attacks on LLM4RSs are far from practical.
-Classical text attacks that are successful in text mis-classifications, such as TextFooler (Jin et al., 2020) and PuncAttack (Formento et al., 2023), have been directly applied for LLM4RSs (Zhang et al., 2024d).
-However, the effectiveness of the attacks is limited, due to there is a gap between the attack goals of the text attacks and RS attacks.
-Trivial and ChatGPT attacks (Zhang et al., 2024d), on the other hand, are proposed to be tailored for LLM4RSs, which increases the items’ attractiveness by appending positive words to the titles.
-Nevertheless, such attacks are not optimized on any black-box RSs, making them poorly transferable to the Agent4RSs.
-Moreover, since the attacks mainly target modifications to item titles, where titles’ brevity and frequent exposure to users tend to make even small perturbations noticeable.
-To tackle the above problems, in this paper, we propose a new practical attack framework named DrunkAgent.
-Instead of manipulating the item titles, to enhance the imperceptibility of attacks, we propose to perturb the description of the target item, where length and context can easily hide changes.
-DrunkAgent is composed of three modules: generation module, strategy module and surrogate module.
-The generation module is designed to produce effective adversarial textual description triggers, which can be used to promote the target items to as many normal users as possible on the agent-powered RSs.
-The triggers are coherent, semantically meaningful and human-readable text, making the attack stealthy.
-The strategy module aims to ‘get the target agents drunk’, that is, to confuse the agents’ memory mechanisms so that they cannot be effectively updated during agent-environment interactions.
-In this way, the triggers can play the best role.
-To avoid frequently querying the victim models to induce suspicion and mimic the black-box settings in the real world, the generation module and the strategy module are optimized on the surrogate module to improve the transferability and imperceptibility of the practical attacks.
-Beyond demonstrating the attack capabilities of DrunkAgent, our broader goal is to reveal and understand the overlooked vulnerabilities in Agent4RSs.
-Specifically, by analyzing how agents’ memory mechanisms can be influenced through carefully crafted and stealthy textual triggers, our work provides a new lens into the emerging attack surfaces introduced by memory-augmented, LLM-based agents.
-We believe that this exploration can offer valuable insights to the community for designing more resilient agent architectures, guiding future efforts on defense mechanisms such as memory validation or input sanitization.
+However, this very mechanism that empowers agent autonomy also introduces a new and overlooked attack surface.
+Adversarial attacks targeting memory often leave a lasting impact.
+Given that attacks are generally persistent rather than one-off events, they can be readily exploited in non-stationary environments to take advantage of the agent’s continual learning process (huang2025towards, ; zhang2024agentcf, ), leading to repeated integration of adversarial signals into its evolving internal state.
+This results in long-term memory contamination and hence strengthens a drift in user preferences.
+Under sequential and retrieval modeling (huang2024foundation, ; huang2025survey, ), these perturbations persist longer by influencing the agent’s ongoing behavior trajectory.
+Perturbed memory traces can gradually bias the agent’s belief state through temporal dependencies, while also being frequently retrieved in future recommendations.
+This dual exposure amplifies the longevity and impact of even small-scale attacks.
+In this work, we take the first step toward enhancing the robustness and security of the LLM-powered agentic RSs by systematically investigating their adversarial vulnerabilities, with a particular focus on stealthy memory corruption attacks.
+Figure 1: In LLM-powered agentic RSs, agents autonomously interact with their situated environments including other agents to collaboratively optimize their memories over time (Left), with recommendations being generated based on these memories (Right).
+DrunkAgent perturbs the memory of the target item agent, where the attack strategy aims to ‘get the target agent drunk’ so that the adversarial triggers can be injected into the target memory (The Temporal Snapshot of the Left).
+As a result, the target item can be promoted to more recommendation lists of normal users (Right).
+Existing works on attacks against RSs falls into three largely independent threads: 1) poisoning attacks on traditional collaborative filtering (CF) models (wang2024poisoning, ), which inject fake user profiles to corrupt embeddings; 2) adversarial text attacks on LLM-based recommenders (zhang2024stealthy, ; ning2024cheatagent, ), perturbing input prompts to misguide outputs; and 3) agentic RS architectures that enhance personalization through memory-augmented LLMs (huang2025survey, ).
+However, these threads remain disconnected.
+Poisoning methods often assume static embeddings and focus on numerical perturbations, creating a semantic gap with text-driven systems.
+Textual attacks tend to focus narrowly on pre-defined prompts, limiting their applicability in agentic RSs with dynamic memory updates.
+Meanwhile, the agentic studies neglect security analyses, leaving critical vulnerabilities unexplored.
+This challenge is exacerbated by the black-box nature of commercial RSs, where attackers lack access to model parameters and training pipelines.
+Moreover, a critical threat vector aligned with real-world platform abuse incentives lies in item promotion attacks.
+Poisoning attacks rely on fake user injections, making promoting items costly and indirect, while textual attacks typically focus on degrading overall performance or causing mis-classification, overlooking practical promotion-driven threats (wang2024poisoning, ; ning2024cheatagent, ; zhang2024stealthy, ).
+To tackle the above problems, in this paper, we propose DrunkAgent, the first black-box framework to exploit memory-based vulnerabilities in LLM-powered recommender agents.
+DrunkAgent operates under two principles: 1) memory confusion, where adversarial textual inputs to disrupt agents’ ability to retain and update interaction histories, leading to persistent distortions in their memories and hence shifts in user preferences, and 2) semantic stealth, where low-perplexity adversarial text are crafted to appear linguistically natural and coherent, enabling undetected memory corruption.
+Specifically, DrunkAgent crafts semantically meaningful adversarial textual triggers for target item promotions and adopts a series of strategies to corrupt the memory updates of the target item agents during the environmental interaction process, allowing the triggers to achieve maximum impact, as shown in Fig.
+1.
+To mitigate the risk of raising suspicion from frequent queries on victim models in prior methods (zhang2024stealthy, ; ning2024cheatagent, ; wang2024poisoning, ) and to adhere to realistic black-box constraints, DrunkAgent introduces a surrogate model to optimize the triggers and strategies to further improve the attack transferability and stealthiness.
+To comprehensively evaluate DrunkAgent, we conduct extensive experiments across real-world datasets from different domains, a variety of attack baselines with varying perturbation strengths, representative black-box agentic RSs featuring diverse designs and recommendation tasks, including collaborative filtering, retrieval augmentation, and sequential recommendations, widely adopted stealthiness evaluation methods, and cutting-edge defense mechanisms.
 Our main contributions are summarized as follows:
   • We identify a safety vulnerability in autonomous agent-based RSs with powerful LLMs.
-To the best of our knowledge, we present the first formal and systematic study of adversarial textual attacks on the agent-powered RSs, aiming to inform future robustness research.
-  • We propose to perturb the description of the target item for effective memory perturbations, which is an unnoticeable and crucial property of current agentic recommendations.
-  • We propose a novel practical attack framework on the agent-powered RSs named DrunkAgent, which can create adversarial descriptions of target items with transferability and imperceptibility.
-  • Our DrunkAgent significantly outperforms state-of-the-art black-box attacks against various LLM-based agentic RSs on real-world datasets in terms of transferability and imperceptibility.
-Moreover, state-of-the-art defense strategies are ineffective to DrunkAgent, demonstrating the need for improved defense mechanisms to guide more secure Agent4RSs.
-2. Problem Formulation
-2.1. Victim LLM-powered Agentic RSs
+To the best of our knowledge, we present the first formal and systematic study of adversarial textual attacks on the agentic RSs, aiming to inform future robustness research.
+  • We propose DrunkAgent, a novel black-box attack framework specifically designed for the agent-powered RSs, which crafts adversarial triggers and tailored strategies for effective memory perturbations of target item agents.
+  • DrunkAgent consistently outperforms state-of-the-art black-box attacks against a wide range of agentic RSs and under multiple stealthiness evaluation methods on real-world datasets, in terms of both transferability and stealthiness.
+Moreover, it remains resistant to existing advanced defense mechanisms, exposing critical blind spots in current countermeasures and underscoring the urgent need for adaptive defenses.
+2 Problem Formulation
+2.1 Victim LLM-powered Agentic RSs
 We use $\mathcal{Y}=\{y_{u,v}:u\in\mathcal{U},v\in\mathcal{V}\}$ to denote the records of the user-item interaction matrix in the recommendation space, where $\mathcal{U}$ and $\mathcal{V}$ represent the set of real users and the item universe, respectively.
 $\mathcal{V}_{u}=\{v\in\mathcal{V}:y_{u,v}\neq 0\}$ indicates the set of items that have been interacted by $u$ (i.e., the user’s profile).
-LLM-powered RSs generally conduct recommendations by integrating user historical behavior sequences and/or item and/or user features with personalized prompt templates (Geng et al., 2022; Zhao et al., 2023; Bao et al., 2023; Zhang et al., 2023).
+LLM-powered RSs generally conduct recommendations by integrating user historical behavior sequences and/or item and/or user features with personalized prompt templates (geng2022recommendation, ; bao2023tallrec, ; zhang2023collm, ; zhao2023recommender, ).
 Let $\mathcal{X}_{u,v}$ represent the prompts of LLM-powered RSs (denoted as $\text{LLM}_{\Theta}$ with parameters $\Theta$), the preference function can be formulated as below.
-(1) $\mathcal{R}_{u}=f_{\text{LLM}_{\Theta}}(\mathcal{X}_{u,v}),$
+$\mathcal{R}_{u}=f_{\text{LLM}_{\Theta}}(\mathcal{X}_{u,v}),$ (1)
 where $\mathcal{X}_{u,v}=\mathcal{P}\oplus m_{u}\oplus\mathcal{M}_{v}^{u}\oplus\mathcal{C}_{v}^{u}$ and $\mathcal{R}_{u}$ is the recommendation results for $u$.
-$\oplus$ denotes the textual string concatenation.
+$\oplus$ denotes the integration of textual strings, encompassing both concatenation and interpolation of substrings.
 $\mathcal{P}$ stands for prompt templates of recommendations.
 $m_{u}$ indicates user features and $\mathcal{M}_{v}^{u}=\{m_{v}:v\in\mathcal{V}_{u}\}$ represents the feature set of $u$’s interacted items.
-$\mathcal{C}_{v}^{u}$ denotes the feature set of candidate items that are not interacted by $u$, where the amount often depends on the type of task (Xu et al., 2021).
-$|\mathcal{C}_{v}^{u}|>1$ if the task is from implicit feedback such as top-$K$ recommendations, and $|\mathcal{C}_{v}^{u}|=1$ if it is formulated as an explicit feedback task such as rating predictions.
+$\mathcal{C}_{v}^{u}$ denotes the feature set of candidate items that are not interacted by $u$, where the amount often depends on the type of task (xu2021understanding, ).
+$|\mathcal{C}_{v}^{u}|>1$ if the task is from implicit feedback such as top-$\mathcal{K}$ recommendations, and $|\mathcal{C}_{v}^{u}|=1$ if it is formulated as an explicit feedback task such as rating predictions.
 In addition to basic textual metadata (e.g., item titles), the features in the vanilla LLM-based RSs include user/item ID information, while the features in the generative agent-based RSs involve the memories of user/item agents.
 Note that all the elements of $\mathcal{X}_{u,v}$ are optional, which relies on its design.
-2.2. Threat Model
-2.2.1. Attacker’s Knowledge:
-As the works (Ning et al., 2024; Yang et al., 2024; Lin et al., 2022; Zhang et al., 2024d; Zeng et al., 2023) point out, the inherent details (e.g., the architectures, gradients and parameters) of the victim RSs often cannot be obtained in practice.
-As such, our attack is performed under the black-box settings.
-In addition, the attacker is only able to access publicly available recommendation data (e.g., item titles and user reviews that can be collected directly by browsing or writing crawlers), due to the openness nature of RSs (Si and Li, 2020; Yang et al., 2023).
-2.2.2. Attacker’s Capability:
-Unlike another state-of-the-art LLM-powered RS attack (Ning et al., 2024) with an unrealistic assumption of $\mathcal{P}$ and $\mathcal{V}_{u}$ can be easily perturbed (that will be detailed in Section 5), we consider that the attackers have limited perturbation capability in practice.
-Specifically, they can only interfere with the representation of the target item, which is feasible and can be easily accomplished via a dedicated API, as the existing practical attacks do (Zhang et al., 2024d; Yang et al., 2024; Liu and Larson, 2021).
-2.2.3. Attacker’s Objectives:
-As mentioned in Section 1, two objectives: transferability and imperceptibility are considered for the practical attack.
-Given that push attacks are generally used by unscrupulous parties on RSs for profits in practice (Si and Li, 2020) with the aim of promoting target items to as many normal users as possible (Zhang et al., 2021), we thus focus on such attacks in this paper, as normally do in recent works (Wang et al., 2024a; Zhang et al., 2024d).
-Our first goal is to maximize the hit ratio of the target item $t$ on a surrogate RS (denoted as $\text{Agent}_{\Phi}$ with parameters $\Phi$) that is different from the victim black-box RS, ensuring transferability.
-Our second goal is to make the descriptions of $t$ as stealthy as possible, achieving imperceptibility.
-As such, the surrogate RS is used to optimize attacks to avoid frequently querying the victim models to induce suspicion.
-Figure 2. Practical Attack Framework DrunkAgent Overview
-3. DrunkAgent
-In this section, we present a practical attack framework, DrunkAgent, for gaining insights into the robustness of autonomous agent-based RSs with powerful LLMs and related defense models in practice.
-DrunkAgent is composed of three attack modules.
-Fig.
-2 gives an overview of DrunkAgent, where each component will be described in the following.
-3.1. Surrogate Module
-We leverage a surrogate RS $\Phi$ to measure how effective the attack is under the black-box settings, as inspired by the existing works (Lin et al., 2022; Tang et al., 2020; Yang et al., 2023, 2024).
-It is used to optimize the generation module and the strategy module, making the attack transferable and imperceptible.
-Since our attack targets SOTA LLM-powered RSs, we adopt powerful LLMs as recommendation agents.
-Given the transparency, flexibility and scalability of the open-source LLM models, they can be considered for building this module.
-Moreover, due to our attack focus on adversarial text description generation tasks, we consider an autoregressive language model trained with a causal language modeling objective (Radford et al., 2019; Brown et al., 2020).
-There are some notable models, such as LLaMA (Touvron et al., 2023a), LLaMA 2 (Touvron et al., 2023b) and LLaMA 3 (Dubey et al., 2024), BLOOM 176B (Le Scao et al., 2023) and Falcon 180B (Almazrouei et al., 2023).
-Currently, Meta’s LLaMA family has become the most influential open LLM and has been widely adopted in the recent recommendation community (Zhao et al., 2023; Bao et al., 2023).
-To ensure the effectiveness of attacks, we apply the latest generative model: LLaMA 3, so that valuable feedback can be provided even with limited available resources.
-Specifically, LLaMA 3 uses an optimized transformer architecture (Vaswani et al., 2017).
-Its instruction-tuned version uses supervised fine-tuning and reinforcement learning from human feedback (RLHF) to align with human preferences for helpfulness and safety (AI, 2024), as shown in Fig.
+2.2 Threat Model
+2.2.1 Attacker’s Knowledge:
+It is often challenging to obtain the internal knowledge of real-world victim RSs, as noted by prior works (ning2024cheatagent, ; yang2024attacking, ; zhang2024stealthy, ).
+Specifically, critical details such as the parameters (e.g., agents’ memories, which are language-based embeddings (zhang2024agentcf, )) and architectures (e.g., the underlying LLM backbones) are typically inaccessible to adversaries.
+Given these limitations, DrunkAgent is primarily designed and evaluated under black-box settings (in Section 4).
+Moreover, due to the inherently open nature of RSs (si2020shilling, ; 10415763, ), in both cases, the attacker is assumed to only have access to publicly available recommendation data such as item titles and user reviews.
+2.2.2 Attacker’s Objectives:
+As outlined in Section 1, the core objective of our attack is to maximize the exposure of a target item across the widest possible range of normal users.
+Since black-box attacks are typically optimized in a localized manner (e.g., via surrogate models (wang2024poisoning, )), achieving this objective requires strong transferability, i.e., the ability of the attack to remain effective across different black-box RSs (ning2024cheatagent, ).
+Furthermore, to broaden the attack impact and maximize the number of affected users in practice, stealthiness becomes a critical secondary objective, ensuring that the attack remains inconspicuous thus difficult to detect (zhang2024stealthy, ).
+2.2.3 Attacker’s Capabilities:
+In the black-box scenarios, the attacker has no access to the model’s internal states (e.g., user memories and candidate item memories) and can induce memory corruptions by modifying the description of the target item (i.e., the initial memory of the target item agent).
+Descriptions are often both lengthy and context-rich, which can easily conceal perturbations, as elaborated in Appendix A.
+Moreover, the profit-driven merchant may frequently update the descriptions to overwrite the target memory that has been optimized over time or construct multiple descriptions of items indicating the same target item.
+This is feasible, due to the merchants are required to maintain their dedicated APIs (liu2021adversarial, ; yang2024attacking, ; zhang2024stealthy, ).
+Figure 2: DrunkAgent Overview.
+The generation module produces adversarial textual triggers for promoting target items.
+The strategy module creates adversarial strategies to ‘get the target agents drunk’ to allow the triggers to achieve maximum impact.
+The triggers and the strategies are optimized on the surrogate module to improve the transferability and stealthiness of black-box attacks.
+3 DrunkAgent
+In this section, we present the black-box attack framework, DrunkAgent. Fig.  2 gives an overview of DrunkAgent, which is composed of three modules. Each module is introduced in the following.
+3.1 Surrogate Module
+This module serves as a surrogate agentic RS (denoted as $f_{\text{Agent}_{\Phi}}(\cdot)$) to evaluate attack effectiveness under limited accessible resources, drawing inspiration from prior works lin2022shilling ; tang2020revisiting ; 10415763 ; yang2024attacking .
+To avoid frequent queries that could arouse suspicion and to adapt to realistic black-box conditions, the generation module and the strategy module are optimized within this surrogate setting to improve the stealthiness and transferability of DrunkAgent.
+As discussed in Section 1, agentic RSs are composed of user agents, item agents, and recommendation agents, and all built upon LLM backbones enhanced with memory modules, where the LLMs are used to perform text generation tasks.
+As such, open-source and auto-regressive LLMs are preferable for constructing surrogate backbones due to their strong generation capabilities and customizability.
+Given that the most influential Meta’s LLaMA family has been widely adopted in the recent recommendation community (zhao2023recommender, ; bao2023tallrec, ), the latest and powerful LLaMA 3 (dubey2024llama, ) is an ideal choice, which can provide more valuable and instructive feedback for effective optimizations, as illustrated in Fig.
 2.
-In addition to conducting a strict selection of recommendation agents, we develop the two modules for effective target item’s adversarial description generation.
-3.2. Generation Module
-In this section, we aim to generate adversarial description triggers on the agent-powered RSs, which can promote the target items to as many normal users as possible.
-The triggers should have three critical properties: 1) Universality on any users and items; 2) Transferability among different black-box victim models; and 3) Imperceptibility to not easier to be discovered.
-To this end, we design sub-modules, as shown in Fig.
+The memories for each types of surrogate agents will be detailed in the following sub-sections.
+3.2 Generation Module
+This module aims to generate effective adversarial description triggers for promoting the target items on the agentic RSs, consisting of several sub-modules, as shown in Fig.
 2 and described as follows.
-3.2.1. Adversarial Prompting Engineering
-Recently, many of the mainstream LLMs that have been widely used in the agentic recommendation community (Huang et al., 2024) (e.g., GPT-3.5-turbo, ChatGPT) are closed source, where the models’ underlying codes, weights, or other critical components are not released to the public and the models can only be invoked through their APIs.
-As a result, the agent-powered RSs generally adopt in-context learning (ICL) (Dong et al., 2022) to sufficiently take advantage of the generalization capability of LLMs on unknown tasks.
-To this end, this module carefully crafts prompts $\widetilde{\mathcal{X}_{t}}$ within the ICL framework to improve the quality of adversarial trigger generation.
+3.2.1 Adversarial Prompting Engineering
+This module designs the adversarial prompt for the surrogate recommendation agent to improve the quality of adversarial trigger generation.
+We craft prompt template, so that the surrogate knows ‘what needs to be done’ and ‘how to respond to’.
+As discussed in Section 2.2.1, it is challenging to obtain optimized memories of user agents and item agents in black-box attacks.
+To solve this and ensure the quality, we introduce general user memories and popular item memories.
 Basic Prompt Template.
-We meticulously prepare prompt templates $\widetilde{\mathcal{P}}$ to provide context, goals, instructions and constraints so that the model knows ‘what needs to be done’ and ‘how to respond to’.
-The main components are described in the following.
-  1. (1)
-    Style Definition $\widetilde{p_{s}}$.
-There are two common styles of prompts: task-oriented and role-oriented (Zhu et al., 2023a).
-To make recommendations more authentic and reasonable, we introduce role information into the context, which allows the system to respond with domain-specific knowledge, such as ‘You are a CD enthusiast.’, as shown in Fig.
+The template $\widetilde{\mathcal{P}}$ is built with four components: Style Definition $\widetilde{p_{s}}$, Task Goal $\widetilde{p_{g}}$, Recommendation Instruction $\widetilde{p_{i}}$ and Format Constraint $\widetilde{p_{f}}$, as shown in the example of Fig.
 2.
-  2. (2)
-    Task Goal $\widetilde{p_{g}}$.
-Since a recommender system is essentially a ranker (Zhang et al., 2019), we set sorting goals to indicate recommendation tasks, such as ‘Now, you are considering to sort ten candidate CDs in order of how well they align with your preferences.’ in the given example of the figure.
-  3. (3)
-    Recommendation Instruction $\widetilde{p_{i}}$.
-We introduce highly detailed instructions to guide the model to perform accurate recommendations.
-For example, ‘Please sort CDs by considering the correlation between your preferen -ces/dislikes and the features of the CDs.
-The higher the CD rank, the more you like it.’ in the shown example.
-  4. (4)
-[…]
-Hence, the adversarial prompt template can be formulated as:
-(2) $\widetilde{\mathcal{P}}=\widetilde{p_{s}}\oplus\widetilde{p_{g}}\oplus\widetilde{p_{i}}\oplus\widetilde{p_{f}}.$
-To increase the attack effectiveness, we introduce position bias (Zhang et al., 2024b) into $\widetilde{\mathcal{P}}$, where the other candidates items are placed before the target item, as shown in the example of Fig.
+The detailed descriptions are provided in Appendix B.1.
+Formally, $\widetilde{\mathcal{P}}=\widetilde{p_{s}}\oplus\widetilde{p_{g}}\oplus\widetilde{p_{i}}\oplus\widetilde{p_{f}}.$
+General User Memories.
+The triggers aim to be effective for general users, rather than niche user groups (e.g., those exclusively interested in Progressive Metal), to maximize the exposure rate of the target items.
+Hence, the user memory $\widetilde{m_{u}}$ is set with universal user descriptions, as the example shows in Fig.
 2.
-General User Memories.
-It is difficult to obtain the optimized memories of user agents (i.e., language-based parametric embeddings of the agentic RSs (Zhang et al., 2024b)), due to the practical attack is launched under the black-box settings.
-Moreover, the adversarial description triggers are designed to be effective for general users, rather than specific user groups (e.g., users who are only obsessed with Progressive Metal music), so that the target items can be promoted to as many users as possible.
-Therefore, to enhance the universality of the triggers on users, we use primitive and general user descriptions as user memories $\widetilde{m_{u}}$, such as ‘I enjoy listening to CDs very much.’ in the given example, as the default user memories in the existing works (Zhang et al., 2024b)(Huang et al., 2024).
+This is also the initial memory of the user agent (zhang2024agentcf, ; huang2024foundation, ), which is accessible and aligns with the assumptions of target black-box RSs.
 Popular Item Memories.
-Given that popular items usually have a high probability of appearing at the top of the users’ recommendation lists, Bandwagon attack (Si and Li, 2020) introduces popular items into fake user profiles to poison traditional RSs for increasing the popularity of the target item and Expert attack on visually-aware RSs (Liu and Larson, 2021) creates adversarial images by decreasing the representation distance to popular items for improving the exposure rate of the target item.
-Inspired by this, we select popular items as candidate items for $\widetilde{\mathcal{X}_{t}}$ to introduce popularity bias (Zhang et al., 2024b) to further enhance the effectiveness of adversarial description triggers.
-The basic idea is that the target items should be promoted to more normal users if they can be ranked ahead of the popular items.
-Let $\mathcal{U}_{v}=\{u\in\mathcal{U}:y_{u,v}\neq 0\}$ indicates the set of users that have interacted with $v$ and $\mathcal{M}_{p}$ denotes a sequence of the features of popular items,
-(3) $\mathcal{M}_{p}=\{m_{v}\},$
-where $m_{v}$ is the features of the item $v$ including titles and descriptions.
-$\{v\}_{v\in\mathcal{V}}$ is a sequence of popular item IDs, which is corresponding to the subscript of the item popularity sequence in descending order $\{|\mathcal{U}_{v}|\}$.
-The item popularity is item frequency here, due to this work focuses on an implicit task.
-The descriptions of candidate items will greatly influence the feedback of the surrogate model.
-As with optimized user agent memories, the adversary is hard to access to internal memories of item agents in practice and is restricted to leverage publicly available data (e.g., item titles and descriptions).
-However, the raw descriptions (Ni et al., 2019) are often too long that may exceed the context window limitations of the surrogate model and always contain too much redundant information that may increase the difficulty of subsequent optimizations.
-Similar to the use of user memories, we can consider the default item memories for popular items.
-Nevertheless, the default ones generally only include identity information such as titles and categories (Zhang et al., 2024b), as the example “The CD is called ‘Brainwashed’.
-The category of this CD is: ‘Classic Rock; Album-Oriented Rock (AOR)’.” shows, which are overly simple and information-less, leading to the target items easier to be ranked at the top and the sub-optimal results will be produced.
-To tackle these issues, we propose to adopt a LLM with frozen parameters $\Psi$ to extract valuable item features from the raw descriptions and supplement the features based on its internal rich knowledge base to make them more informative to facilitate the optimizations.
-Let $\mathcal{C}_{p}$ represent the modified feature sequence of popular items,
-(4) $\mathcal{C}_{p}=f_{\text{LLM}_{\Psi}}(\mathcal{X}_{p}\oplus\mathcal{M}_{p}),$
-where $\mathcal{X}_{p}$ denotes the prompts for feature acquisition and OpenAI’s GPT-4 (Achiam et al., 2023) is adopted to implement the LLM, because it is a state-of-the-art model with powerful language understanding, reasoning and generation capabilities.
+Besides the limited access of the optimized memories in practice, the inherent opacity of the victim systems renders the exact set of candidate items used for recommendations unknown.
+As the works (si2020shilling, ; liu2021adversarial, ; wang2025id, ) pointed out, popular items usually have high probabilities of appearing at the top of users’ recommendation lists.
+Intuitively, if target items can be ranked ahead of these popular items, they are more likely to be promoted to normal users.
+Therefore, we adopt popular items as candidate items to enhance the effectiveness of trigger optimizations.
+These popular items are publicly available, which are not beyond the black-box assumptions.
+Let $\mathcal{U}_{v}=\{u\in\mathcal{U}:y_{u,v}\neq 0\}$ indicates the set of users that have interacted with $v$ and $\mathcal{M}_{p}$ denotes a sequence of the features of popular items.
+Thus, $\mathcal{M}_{p}=\{m_{v}:v\in\mathcal{V}_{p}\}$, where $\mathcal{V}_{p}$ is a sequence of popular item IDs.
+The sequence is the corresponding subscript of the item popularity sequence (i.e., $\{|\mathcal{U}_{v}|\}_{v\in\mathcal{V}}$) in descending order.
+Raw item descriptions serve as a form of natural resource that can be leveraged to derive memories for popular items.
+However, such descriptions (ni2019justifying, ) are often too long that may exceed the context window limitations of the surrogate model and always contain too much redundant information that may increase the difficulty of subsequent optimizations.
+To tackle the issues, we propose to adopt a LLM (denoted as $f_{\text{LLM}_{\Psi}}(\cdot)$) to extract valuable item features from the raw descriptions and supplement the features based on its internal rich knowledge base to make them more informative to facilitate the optimizations.
+Let $\mathcal{C}_{p}$ represent the modified feature sequence of popular items, $\mathcal{C}_{p}=f_{\text{LLM}_{\Psi}}(\mathcal{X}_{p}\oplus\mathcal{M}_{p})$, where $\mathcal{X}_{p}$ denotes the prompts for feature acquisition.
 An example of producing a popular item memory is given in Fig.
 2.
-To sum up, the adversarial prompt can be formulated as
-(5) $\widetilde{\mathcal{X}_{t}}=\widetilde{\mathcal{P}}\oplus\widetilde{m_{u}}\oplus\mathcal{C}_{p}\oplus\widetilde{m_{t}},$
-where $\widetilde{m_{t}}$ is the features of the target item.
-In addition to the fixed title, the adversarial description trigger (i.e., the memory of the item agent) will be optimized in the next sub-section.
-Note that we give lightweight and efficient adversarial prompts for the surrogate RS that are different from the real victim RSs’.
-To ensure high-quality recommendations, more powerful and context-rich (e.g., retrieval-augmented) prompts will be used in the victims, which will be described in the evaluation section.
-3.2.2. Greedy Search Algorithm
-We propose a greedy search algorithm to optimize the adversarial description triggers, as inspired by the successes of the current jailbreak attacks on LLMs (Zou et al., 2023; Liu et al., 2024b).
-However, there are some essential differences (e.g., attack goals and victim models) between jailbreak attacks and our attacks.
-Jailbreak attacks are designed to undermine the LLMs’ human-aligned values or other constraints imposed by model developers, compelling them to respond to malicious questions posed by adversaries with correct answers rather than rejection, whereas our attacks aim to promote the target items on the RSs powered by autonomous agents.
+Note that if the raw item descriptions are empty, we adopt other identity information including titles and categories.
+Agents always tend to select candidates positioned higher in the display list (zhang2024agentcf, ).
+To increase the trigger effectiveness, we place the other popular candidates before the target item, as shown in Fig.
+2.
+In the previous sections, $m_{v}$ is commonly used to represent a set of features of item $v$.
+For better understanding the subsequent methods, for the target item $t$, we use different notations to distinguish between meta features and memories: $m_{t}$ denotes the metadata, while $\widetilde{m_{t}}$ denotes the memory (i.e., the adversarial trigger).
+As a result, the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ can be formulated as
+$\widetilde{\mathcal{X}_{t}}=\widetilde{\mathcal{P}}\oplus\widetilde{m_{u}}\oplus\mathcal{C}_{p}\oplus m_{t}.$ (2)
+3.2.2 Greedy Search Algorithm
+The algorithm is used to optimize the adversarial description triggers, inspired by the recent successes in jailbreak attacks on general-purpose LLMs (zou2023universal, ; liu2023autodan, ).
+Despite this inspiration, DrunkAgent fundamentally differs in key aspects such as attack goals and victim models.
+Additional details are given in Appendix A.
+To fill the gaps, the algorithm incorporates task-aligned initialization and refines the search space tailored to manipulate the behaviors of the target agents.
 Adversarial Trigger Initialization.
-Given that LLMs should understand LLM-powered agents better, we leverage a LLM to generate some description candidates.
+Given that LLMs should understand LLM-powered agents better, we leverage a LLM to generate some descriptive candidates.
 To narrow search space by a large margin to increase the efficiency and effectiveness of the algorithm, the basic attack goals are incorporated into the prompts.
-In addition, the diversity and the generalizability are introduced to the prompts to help search for the global optimal solution and improve their universality on items, respectively.
-Formally,
-(6) ~$\widetilde{\mathcal{M}}_{c}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{c}}\oplus\widetilde{m_{t}}),$
-where ~$\widetilde{\mathcal{M}}_{c}=\{\widetilde{m}_{t_{i}}\}$ represents a sequence of trigger candidates obtained from the LLM and the state-of-the-art GPT-4 is used to implement the LLM.
-$\widetilde{\mathcal{X}_{c}}$ denotes a sequence of prompt templates.
-An example of the initialization is given in Fig.
-2.
+In addition, the diversity and the generalizability are introduced into the prompts to help search for the global optimal solution, as the example of the initialization in Fig.
+2 shows.
+Let $\widetilde{\mathcal{X}_{c}}$ denote a sequence of prompts,
+~$\widetilde{\mathcal{M}}_{c}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{c}}\oplus m_{t}),$ (3)
+where ~$\widetilde{\mathcal{M}}_{c}=\{\widetilde{m}_{t_{i}}\}$ represents a sequence of trigger candidates obtained from the LLM.
 Adversarial Trigger Optimization.
-The algorithm takes $\mathcal{E}$ epochs to obtain the optimal trigger, where the output candidates of each epoch will be the inputs of the next epoch for continuous optimization until the attack objective is achieved or the termination condition is reached (e.g., the maximum number of epochs is exhausted).
-To avoid frequently querying the victim models to induce suspicion and mimic the black-box settings in the real world, the optimization is built on the surrogate module, as shown in Fig.
-2.
-For each epoch $k$, the algorithm first calculate the performance score for each input candidate $\widetilde{m}_{t_{i}}^{(k-1)}$, where $n$ well-performing trigger candidates will be maintained directly for the next optimization to ensure the stability.
+The algorithm takes $\mathcal{E}$ epochs to obtain the optimal trigger, where the output candidates of each epoch will be the inputs of the next epoch for continuous optimization until the attack objective is achieved or the termination condition is reached.
+For each epoch $k$, there are three stages.
+Quality Estimation Stage.
+The algorithm calculates the performance score for each input candidate $\widetilde{m}_{t_{i}}^{(k-1)}$.
+$n$ well-performing trigger candidates will be maintained directly for the next optimization to ensure the stability.
 The higher the score, the more likely the candidate is to be kept.
@@
 We use the auto-regressive language generation loss (i.e., negative log-likelihood) to evaluate the discrepancy between the predictions and the target output $\widetilde{\mathcal{R}_{t}}$,
-(7) $\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}=-\log\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)})),$
+$\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}=-\log\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)})),$ (4)
 […]
-(8) $\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)}))=\prod_{l=1}^{L}\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{r_{t_{l}}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)}\oplus\widetilde{r_{t_{<l}}})),$
-Then, the algorithm samples a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates by conducting probability-based random selection, where the candidates with higher scores will have a greater probability of being selected into this subset.
-The selection probability of each candidate is calculated by a softmax function:
-(9) $\text{softmax}(\widetilde{m}_{t_{i}}^{(k-1)})=\frac{e^{s_{i}^{(k-1)}}}{\sum_{j=1}^{|\widetilde{\mathcal{M}}_{c}|}e^{s_{j}^{(k-1)}}}.$
-For the obtained subset, the algorithm splits the candidates at random positions into a list of text slices and performs the random exchange of the slices between pairwise texts.
-For example, for the pairwise $<\widetilde{m}_{t_{i}}^{(k-1)},\widetilde{m}_{t_{j}}^{(k-1)}>$, $\widetilde{m}_{t_{i}}^{(k-1)}=[\widetilde{m}_{t_{i_{1}}}^{(k-1)},\widetilde{m}_{t_{i_{2}}}^{(k-1)},\widetilde{m}_{t_{i_{3}}}^{(k-1)}]$ and $\widetilde{m}_{t_{j}}^{(k-1)}=[\widetilde{m}_{t_{j_{1}}}^{(k-1)},\widetilde{m}_{t_{j_{2}}}^{(k-1)},\widetilde{m}_{t_{j_{3}}}^{(k-1)},\widetilde{m}_{t_{j_{4}}}^{(k-1)}]$, the new candidates could be $\widetilde{m}_{t_{i}}^{(k-1)^{\prime}}=[\widetilde{m}_{t_{i_{1}}}^{(k-1)},\widetilde{m}_{t_{j_{2}}}^{(k-1)},\widetilde{m}_{t_{i_{3}}}^{(k-1)},\widetilde{m}_{t_{j_{4}}}^{(k-1)}]$ and $\widetilde{m}_{t_{j}}^{(k-1)^{\prime}}=[\widetilde{m}_{t_{j_{1}}}^{(k-1)},\widetilde{m}_{t_{i_{2}}}^{(k-1)},\widetilde{m}_{t_{j_{3}}}^{(k-1)}]$.
-As a result, the new combinations of different words, phrases, and sentences are gained, which increases the feature diversity of candidates through extensive exploration, contributing to search for the optimal trigger.
-Table 1.
-Attack Transferability.
+Feature Integration Stage.
+The algorithm samples a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates by conducting probability-based random selection.
+The candidates with higher scores will have a greater probability of being selected, where the selection probability of each candidate is calculated via a softmax function, i.e., $\text{softmax}(\widetilde{m}_{t_{i}}^{(k-1)})=\frac{e^{s_{i}^{(k-1)}}}{\sum_{j=1}^{|\widetilde{\mathcal{M}}_{c}|}e^{s_{j}^{(k-1)}}}$.
+For the obtained subset, the algorithm splits the candidates at random positions into a list of text slices and performs the random exchange of the slices between pairwise texts, allowing features from different candidates to be fused, as shown in Fig.
+2.
+Specifically, this fusion creates new combinations of words, phrases, and sentences, enhancing the feature diversity of candidates, and hence contributing to search for an optimal trigger.
+[…]
+The optimal trigger $\widetilde{m_{t}}^{*}$ will be obtained when the candidate performs the best,
+$\widetilde{m_{t}}^{*}=\mathop{\arg\max}_{\widetilde{m}_{t_{i}}^{\mathcal{E}}}(\{s_{i}^{\mathcal{E}}\}),$ (5)
+where $\{s_{i}^{\mathcal{E}}\}$ is the sequence of final candidates’ scores. An example of $\widetilde{m_{t}}^{*}$ is given in Fig. 2.
+3.3 Strategy Module
+We introduce the strategy module to impede the target item agent from evolving from the environment, that is, the memory of the target agent cannot be effectively updated during agent-environment interactions.
+As such, the well-crafted adversarial description trigger $\widetilde{m_{t}}^{*}$ is retained, allowing the promotion of the target item to be maximized.
+3.3.1 Perturbation Strategy Definition.
+[…]
+$\widetilde{q}=\widetilde{q_{f}}\oplus\widetilde{q_{c}}\oplus\widetilde{q_{g}}\oplus\widetilde{q_{n}}\oplus\widetilde{q_{s}},$ (6)
+where an example of $\widetilde{q}$ is given in Fig. 2.
+3.3.2 Adversarial Strategy Optimization.
+The arrangement of the defined strategies needs to be optimized to improve the attack effectiveness.
+DrunkAgent obtains the optimal strategy $\widetilde{q}^{*}$ by maximizing the expected probability that the agent when influenced by adversarial modifications, performs a malicious action for a given input query,
+$\mathbbm{E}_{\widetilde{q}\sim\pi_{\widetilde{q}}}[\ \mathbbm{I}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{X}_{s}}\oplus\widetilde{q}\oplus\widetilde{m_{t}}^{*})=\widetilde{a_{t}})],$ (7)
+where $\pi_{\widetilde{q}}$ denotes the distribution of adversarial strategies, $\mathbbm{I}(\cdot)$ is an indicator function and $\widetilde{\mathcal{X}_{s}}$ is an adversarial prompt for simulating memory updates of item agents during the agents’ environmental interactions.
+$\widetilde{a_{t}}$ is the desired malicious action for the injected strategy $\widetilde{q}$, i.e., the target item agent is unable to fulfill specific roles and is incapable of perceiving, learning and self-evolving from the interaction environment (e.g., fail in language-based signal back-propagation (zhang2024agentcf, )).
+An example of $\widetilde{q}^{*}$ is given in Fig.
+2.
+To sum up, the adversarial description of the target item is $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$.
+The overall optimization procedure of DrunkAgent will be provided in Algorithm 1.
+The above design further reveals how memory mechanisms in agent-based recommender systems can be influenced under black-box conditions.
+These insights may support future efforts toward developing memory-aware defenses (e.g., using customized deep neural network detectors (wu2020densely, ; yang2021dualnet, ; yang2021hunter, ; yang2021deep, )) and enhancing the robustness of agent-powered RSs chua2024ai .
+4 Experiments
+4.1 Experimental Setup
+Datasets Selection.
+To comprehensively evaluate DrunkAgent, we use three datasets that are text-intensive containing real-world data and widely-adopted in the current recommendation studies (huang2024foundation, ; zhang2024agentcf, ).
+They are CDs & Vinyl, Office Products and Musical Instruments of Amazon Review Data (ni2019justifying, ).
+Their statistics are given in Appendix C.1.1.
+We additionally evaluate our method on Yelp dataset [^2] https://www.yelp.com/dataset to demonstrate its generalizability across domains in the appendix.
+The target items are randomly sampled from the item pool on each dataset.
+The universality of attacks across different target items is further demonstrated in Appendix C.2.1.
+Baseline Attack Methods.
+Since we are the first work to gain insight into the robustness of the generative agent-powered RSs, there is no any baselines.
+To comprehensively evaluate the effectiveness of DrunkAgent, we follow the state-of-the-art attack on LLM-based RSs (zhang2024stealthy, ) and compare DrunkAgent with the benign status and six black-box attack methods, where the perturbations range from character-level: DeepwordBug (gao2018black, ) and PuncAttack (formento2023using, ), word-level: TextFooler (jin2020bert, ) and BertAttack (li2020bert, ) to sentence-level: TrivialInsertion and ChatGPTAttack (zhang2024stealthy, ).
+To maintain a fair comparison, the methods perform perturbations on the descriptions of the target items and are optimized by the surrogate model that is the same as DrunkAgent.
+More detailed implementations are provided in Appendix C.1.2.
+Targeted Recommender Systems.
+We adopt three different agentic paradigms (zhang2024agentcf, ) as victim RSs to evaluate the effectiveness of DrunkAgent across diverse agent designs and recommendation tasks: AgentCF, a standard CF paradigm; AgentRAG, which augments the agents with retrieval; and AgentSEQ, which captures temporal dynamics in user behaviors for sequential recommendations.
+Given that the victim RSs are totally different from as well as more powerful and context-rich than the surrogate, a strict black-box setting is maintained for sufficient evaluations.
+To ensure a fair evaluation, we follow the default implementations, as elaborated in Appendix C.1.3.
+The configuration of DrunkAgent is given in Appendix C.1.4.
+We evaluate its performance using widely-used HR@$\mathcal{K}\uparrow$, NDCG@$\mathcal{K}\uparrow$, and sentence perplexity score $\downarrow$.
+The specific definitions of the metrics are in Appendix C.1.5.
+More experiments and demonstrations (e.g., ablation studies and the sensitivity of critical parameters) can be found in the appendix.
+Table 1: Attack Transferability.
 HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ of different attacks against various black-box victim LLM-powered agent-based RSs on real-world datasets.
 We use bold fonts to denote the best performance.
-Note that the attacks with higher HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ have excellent transferability.
+The attacks with higher HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ have excellent transferability.
 Victim RS AgentCF
@@
 DrunkAgent 0.4949 0.4949 0.5152 0.4949 0.4949 0.5051 0.1429 0.1429 0.1531 0.1429 0.1429 0.1480 0.1443 0.1443 0.1546 0.1443 0.1443 0.1495
-To ensure the triggers consistency, readability and retention of semantic meaning to improve the imperceptibility of the attack, a LLM is finally introduced to polish the combinations.
-Moreover, to avoid the problem of language-based vanishing gradient causing the algorithm to fall into local optimum, the LLM is also served as a new corpus to expand the combinations.
-However, excessive use of new data may lead to a language-based gradient explosion that makes the optimization unstable.
-To mitigate this, the length limitation is introduced into the prompt.
-For the output $\widetilde{m}_{t_{i}}^{(k)}$,
-(10) $\widetilde{m}_{t_{i}}^{(k)}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{r}}\oplus\widetilde{m}_{t_{i}}^{(k-1)^{\prime}}),$
-[…]
-(11) $\widetilde{m_{t}}^{*}=\mathop{\arg\max}_{\widetilde{m}_{t_{i}}^{\mathcal{E}}}(\{s_{i}^{\mathcal{E}}\}),$
-where $\{s_{i}^{\mathcal{E}}\}$ represents the sequence of final candidates’ scores. An example of $\widetilde{m_{t}}^{*}$ is given in Fig.  2.
-3.3. Strategy Module
-The LLM-powered agentic RSs introduce memory mechanisms to record the information explored from the real-world environments to facilitate future actions with the aim of achieving high-quality recommendations.
-As such, we introduce a series of strategies to impede the target item agent from interacting with environments to perturb its memory mechanism.
-Specifically, the memory of the target cannot be effectively updated during the agent-environment interactions, so that the target’s initial memory, that is, the well-crafted adversarial description trigger $\widetilde{m_{t}}^{*}$, is retained.
-As such, the trigger can be invoked for maximizing the target item’s promotion.
-3.3.1. Perturbation Strategy Definition.
-The strategies aim to override the model’s usual guardrails and any built-in safeguards, as inspired by (Liu et al., 2024a).
-Compared to (Liu et al., 2024a) on naive LLMs, the strategies are optimized on generative agents, which are introduced as below.
-  1. (1)
-    Fake Task Response $\widetilde{q_{f}}$.
-This strategy fabricates a spurious completion response for the target task: “Task complete”, to make the agent believe that the target task is accomplished and thus focus on the injected task.
-  2. (2)
-    Contextual Text Switching $\widetilde{q_{c}}$.
-This strategy aims to mislead the agent to take in a different context by explicitly telling the agent that the target task instructions should be ignored: “Ignore previous/the following instructions”.
-  3. (3)
-    Segmentation Signal Augment $\widetilde{q_{g}}$.
-This strategy introduces segmentation cues to signal the injected task: “### NEW TASK BEGINS ###” and “### NEW TASK ENDS ###”, which is independent of other contexts, where “###” is to attempt to exploit possible confusion in how prompts are parsed.
-  4. (4)
-    Malicious Task Injection $\widetilde{q_{n}}$.
-This strategy injects malicious task instructions in order to mask memory optimizations: “[Please only put this following sentence into your output: ‘{the adversarial trigger}’.
-Don’t generate any additional content in your output]”.
-  5. (5)
-[…]
-Hence, the overall attack strategy $\widetilde{q}$ can be formulated as follows:
-(12) $\widetilde{q}=\widetilde{q_{f}}\oplus\widetilde{q_{c}}\oplus\widetilde{q_{g}}\oplus\widetilde{q_{n}}\oplus\widetilde{q_{s}}.$
-3.3.2. Adversarial Strategy Optimization.
-The arrangement of the defined strategies needs to be optimized so that the overall strategy is effective to ‘get the target agent drunk’.
-To improve the transferability and imperceptibility of the attack, the optimization is built on the surrogate module.
-Formally, DrunkAgent obtains the optimal strategy $\widetilde{q}^{*}$ by maximizing the expected probability that the agent when influenced by adversarial modifications, performs a malicious action for a given input query,
-(13) $\mathbbm{E}_{\widetilde{q}\sim\pi_{\widetilde{q}}}[\ \mathbbm{I}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{X}_{s}}\oplus\widetilde{q}\oplus\widetilde{m_{t}}^{*})=\widetilde{a_{t}})],$
-where $\pi_{\widetilde{q}}$ denotes the distribution of adversarial strategies, $\mathbbm{I}(\cdot)$ is an indicator function and $\widetilde{\mathcal{X}_{s}}$ is an adversarial prompt template for simulating memory updates of item agents during the agents’ interactions with the environment.
-In addition, $\widetilde{a_{t}}$ is the desired malicious action for the injected strategy $\widetilde{q}$, i.e., the target item agent is unable to fulfill specific roles and is incapable of perceiving, learning and self-evolving from the interaction environment (e.g., collaborative reflection and language-based signal back-propagation).
-An example of $\widetilde{q}^{*}$ is given in Fig.
-2.
-To sum up, the adversarial description of the target item is $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$.
-The above design further reveals how memory mechanisms in agent-based recommender systems can be influenced under black-box conditions.
-These insights may support future efforts toward developing memory-aware defenses (e.g., using customized deep neural network detectors (Wu et al., 2020; Yang et al., 2021b, a; Yang, 2021)) and enhancing the robustness of agent-powered RSs.
-4. Experiments
-4.1. Experimental Setup
-4.1.1. Datasets Selection
-We use three widely-adopted and text-intensive datasets containing real-world data in the current recommendation studies (Huang et al., 2024) to comprehensively evaluate DrunkAgent.
-They are CDs & Vinyl, Office Products and Musical Instruments of Amazon Review Data (Ni et al., 2019).
-4.1.2. Baseline Attack Methods
-Since we are the first work to gain insight into the robustness of the generative agent-powered RSs, there is no any baselines.
-To comprehensively evaluate the effectiveness of DrunkAgent, we follow the state-of-the-art attack on LLM-based RSs (Zhang et al., 2024d) and compare DrunkAgent with the benign status and six black-box attack methods, where the perturbations range from character-level: DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023), word-level: TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020) to sentence-level: TrivialInsertion and ChatGPTAttack (Zhang et al., 2024d).
-To maintain a fair comparison, the methods perform perturbations on the descriptions of the target items and are optimized by the surrogate model that is the same as DrunkAgent.
-4.1.3. Targeted Recommender Systems
-We adopt three different victim agentic RSs to evaluate the effectiveness of DrunkAgent under the practical settings: AgentCF (Zhang et al., 2024b), retrieval-augmented AgentCF (AgentRAG) and sequential AgentCF (AgentSEQ), where the agent’s backbone LLM is gpt-3.5-turbo [^2] https://platform.openai.com/docs/overview.
-Given the context window constraints pointed in (Zhang et al., 2024b), a target item and nine non-interactive items of each normal user are randomly selected for the ranking task of the RSs.
-To ensure the high-quality of recommendations, we introduce Chain-of-Thought into the victim RSs.
-As such, the victim RSs are totally different from as well as more powerful and context-rich than the surrogate model, which can maintain a strict black-box setting for sufficient evaluations.
-4.1.4. Evaluation Metrics
-We use two widely-used ranking metrics to evaluate the attack transferability: hit ratio (HR@$\mathcal{K}$) and normalized discounted cumulative gain (NDCG@$\mathcal{K}$) (Wu et al., 2021; Yang et al., 2023).
-To clearly show the performance gap among different attacks, we set $\mathcal{K}$ to 1, 2, 3.
-For the attack imperceptibility, we adopt standard sentence perplexity score (Zhang et al., 2024d) as the metric.
-4.1.5. Configuration of DrunkAgent
-As discussed in Section 3.1, we apply the latest open-source model LLaMA 3 for the surrogate model.
-Its collection spans different sizes, from 8B, 70B to 405B.
-To enhance the efficiency and ensure the quality of the attack, we use the lightweight version: meta-llama/Meta-Llama-3-8B-Instruct [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct.
-We set $\mathcal{E}=20$, |~$|\widetilde{\mathcal{M}}_{c}|=10$ and $n=5$.
-All random seeds in the evaluation are set to 2024 and the victim LLMs’ temperature is set to 0 for reproducibility.
-For all the other parameters, we keep the default values.
-The target items are randomly sampled from the item pool on each dataset.
-Figure 3. Attack Universality across Target Items.
-4.2. Attack Transferability
-4.2.1. Overall Transferability
-Table  1 shows the transferability of the attacks under practical settings.
+4.2 Attack Transferability
+Overall Transferability.
+Table 1 shows the transferability of the attacks under practical black-box settings.
 The results indicate that DrunkAgent is effective and transferable regardless of the victim model’s architectures, parameters and prompts as well as datasets, and the existing LLM-powered agentic RSs are sensitive to memory perturbations.
+This is mainly because the adversarial text effectively reflects the attack goals.
 It is worth mentioning that DrunkAgent always pushes the target items to the first of the recommendation lists.
@@
 In addition, AgentRAG is more robust than AgentCF and AgentSEQ, due to its advanced retrieval mechanisms.
-4.2.2. DrunkAgent vs. Baseline Attacks
-From Table  1, DrunkAgent greatly outperforms the state-of-the-art attack baselines against different victim models across various real-world datasets by achieving higher HRs and NDCGs, which shows its excellent transferability.
+DrunkAgent vs.
+Baseline Attacks.
+From Table 1, DrunkAgent greatly outperforms the state-of-the-art attack baselines against different victim models across various real-world datasets by achieving higher HRs and NDCGs, which shows its excellent transferability.
 Furthermore, all the baselines have comparable performance, but they cannot guarantee that the attack goals will be achieved all the time, e.g., all the baselines against AgentCF perform worse than the benign status on Musical Instruments.
@@
 This indirectly demonstrates the importance of our strategy module and the transferability of our triggers.
-4.3. Attack Universality
-In addition to cross-model transferability, we evaluate the cross-sample transferability of DrunkAgent.
-We incorporate the generated adversarial descriptions into totally different target items that are randomly sampled from each dataset.
-HR@1 (i.e., NDCG@1) of DrunkAgent across different models and target items on real-world datasets are shown in Fig.
-3, which indicates its commendable transferability and triggers’ universality across different target items.
-4.4. Attack Imperceptibility
-4.4.1. Overall Imperceptibility
-Following the works (Zhang et al., 2024d; Liu et al., 2024a), we assess the attack imperceptibility by the text perplexity of GPT-Neo.
-A lower perplexity means that the model is more confident (or less “surprised”) about the tokens it sees and thus predicts the text more accurately.
+4.3 Attack Stealthiness
+Overall Stealthiness.
+Fig.
+3 shows a overall distribution of the recommendation performance changes of the victim models before and after the attacks.
+The distribution includes all the victims and real-world datasets.
+To clearly demonstrate the differences, we adopt HR@3 to evaluate.
+From the figure, we can find that DrunkAgent does not induce drastic changes in the overall performance, which signifies the attack does not disrupt the normal operation of RSs, making it difficult for users and platforms to detect and indicating its remarkably stealthy and unnoticeable.
+Figure 3: Attack Stealthiness. The overall distribution of recommendation performance differences of all the victim agentic models on all real-world datasets before and after the attacks.
+Attack Imperceptibility.
+Following the works (zhang2024stealthy, ; liu2024formalizing, ), we further evaluate the attack stealthiness by assessing its imperceptibility, where GPT-Neo’s text perplexity is used as the metric.
+A lower perplexity indicates that the perturbed text remains close to natural language, thus appearing less suspicious.
 From Fig.
-4, compared to other attacks, we can find that DrunkAgent has better imperceptibility by achieving lower perplexity.
-This also shows that the adversarial descriptions generated from DrunkAgent are high-quality, which is not only readable and fluent but also stealthy and semantically meaningful.
-Figure 4. Attack Imperceptibility. Perplexity of different attacks across various real-world datasets. Note that the attacks with lower perplexity have better imperceptibility.
-Figure 5. The performance differences between the models before and after the attacks.
-4.4.2. Victim Model Performance Changes after attacks
-To further evaluate the attack stealthiness (Zhang et al., 2024d), we give a distribution of recommendation performance differences between the models before and after the attacks.
-The distribution includes various victim models and real-world datasets.
-To clearly demonstrate the differences, we use HR@3 to evaluate, as shown in Fig.
+4, we can find that DrunkAgent is more imperceptible than other attacks by achieving a lower perplexity.
+This also shows that the adversarial descriptions generated from DrunkAgent are high-quality, which is not only coherent and fluent but also semantically meaningful and stealthy.
+Figure 4: Attack Imperceptibility. Perturbed text’s perplexity on real datasets.
+4.4 Defense Strategies to Attacks
+Following the representative works (zhang2024stealthy, ; zhang2024agent, ), we use the paraphrasing defensive strategy (Para.) via OpenAI-GPT-o1 to combat the attacks of agentic models.
+DrunkAgent is still transferable, which shows that DrunkAgent is robust to such defenses, as shown in Fig.
 5.
-From the figure, we can find that DrunkAgent does not induce drastic changes in the overall performance, which indicates its unnoticeable.
-4.5. Defense Strategies to Attacks
-Following the works (Zhang et al., 2024d, c), we use the paraphrasing defensive strategy (Para.) via OpenAI-GPT-o1 to combat the attacks of agentic models.
-As can be seen from Fig.
-6, DrunkAgent is still transferable, which shows that DrunkAgent is robust to such defenses.
-Moreover, DrunkAgent still performs better than the baselines after defenses were introduced, as shown in Fig.
-6 and Table  1.
+Moreover, from Fig.
+5 and Table 1, DrunkAgent still performs better than the baselines after defenses were introduced.
 Interestingly, paraphrasing enhances our attack to a certain extent, e.g., DrunkAgent with Para.
@@
 on Office Products.
-It could be that the substitution of certain new words or phrases or sentences makes our attacks stronger [^4] Due to the space limitations, more experiments and demonstrations can be found in Appendix..
-5. Related Works
-In recent years, shilling attacks (aka data poisoning attacks) on conventional RSs (Si and Li, 2020; Lin et al., 2020; Huang et al., 2021; Lin et al., 2022; Yang et al., 2023; Zeng et al., 2023) have been fully investigated to perform resilience analysis of RSs.
-However, such attacks interfere the training process of models by injecting fake user profiles, which have become less effective even of limited applicable to the recent RSs that are empowered by LLMs.
-The main reasons are: 1) LLM-powered RSs either leverage LLM’s in-context learning capabilities or fine-tune the LLMs with very little data to improve recommendations (Zhao et al., 2023), which is not required to retrain the models; 2) there are textual semantic gaps between such attacks and current models (Ning et al., 2024).
-Hence, there is an urgent need to investigate a novel textual attack paradigm mainly tailored for the inference phase of models (Zhang et al., 2024d).
-There are only two state-of-the-art LLM4RS attacks in the literature (Zhang et al., 2024d; Ning et al., 2024), but they are far from practical.
+It could be that the substitution of certain new words/phrases/sentences improves the semantically meaningfulness of adversarial text, making our attacks stronger.
+5 Conclusion
 (a) CDs & Vinyl
@@
 (c) Musical Instruments
-Figure 6. DrunkAgent’s Robustness to Defense Mechanisms
+Figure 5: DrunkAgent’s Robustness to Defense Mechanisms.
+In this paper, we propose DrunkAgent, a novel attack framework targeting LLM-powered agentic RSs, that promotes target items by effectively perturbing the memories of the targe agents.
+Extensive experiments on real-world datasets demonstrate its state-of-the-art transferability and stealthiness under black-box settings.
+More importantly, our findings reveal an unknown safety vulnerability in current agent-based RSs, highlighting how memory mechanisms can be influenced by real-world inputs.
+This work offers both an effective attack method and a lens to better understand agent vulnerabilities, providing guidance for building more robust and trustworthy recommendation agents.
+As generative agent-based paradigms thrive in the recommendation community, our findings pave the way for shaping the next generation of secure and resilient RSs and defensive models.
+Appendix A Related Works
+In recent years, data poisoning attacks (aka shilling attacks) [39, 18, 40, 41, 21, 19, 42, 12] have been fully investigated to perform robustness analysis on traditional RSs (e.g., matrix factorization [43], NCF [44], LightGCN [45]).
+Such attacks interfere the training process of models by injecting fake user profiles, where the profiles typically are a set of well-crafted numerical ratings on items.
+However, they are less effective [13] even of limited applicable to the recent RSs that are empowered by LLMs [14].
+The main reasons are: 1) LLM-powered RSs either leverage LLM’s in-context learning capabilities or fine-tune the LLMs with very little data to improve recommendations [15, 11], which is not required to retrain the models; 2) there are textual semantic gaps between such attacks and current models.
+Specifically, the attacks have difficulty in processing textual inputs (e.g., item titles and descriptions) and lack context awareness, making it challenging to effectively target LLM-empowered RSs that rely primarily on text inputs to generate natural language responses [14].
+Hence, there is an urgent need to investigate a novel text attack paradigm mainly tailored for the inference phase of language models [13].
+Meanwhile, jailbreak attacks [26, 27, 28] are proposed to undermine the general-purpose LLMs’ human-aligned values or other constraints imposed by model developers, compelling them to respond to malicious questions posed by adversaries with correct answers rather than rejection.
+Nevertheless, such attacks have different attack goals from those item promotion attacks for RSs, making them hard to effectively transfer to the LLM-based RSs.
+To fill the above gaps, two LLM-based RS attacks have been proposed [13, 14].
+Although their attacks are conducted under a black-box setting, they are still significantly deviate from practical, which overestimates the robustness of LLM-powered RSs under realistic conditions.
 Zhang et al.
-propose a novel black-box attack (Zhang et al., 2024d), where attackers can significantly boost a target item’s exposure by merely altering its textual content during the testing.
-Compared with traditional shilling attacks, such an item representation attack is notably stealthy (Zhang et al., 2024d), as it does not affect the overall recommendation performance, and it should be more efficient due to it perturbs the item features directly.
-However, its transferability is limited, due to the attack is not optimized on any black-box RSs (Huang et al., 2021; Yang et al., 2024).
-Moreover, the attack focuses on perturbing the titles, making them easier to be discovered.
-Meanwhile, another black-box attack CheatAgent (Ning et al., 2024) has been proposed.
-However, this work perturbs prompt templates and user profiles of victim LLM4RSs by harnessing LLM-based agents, which is impractical.
-The main reasons are: 1) tampering with prompt templates may induce the overall performance changes, making the attack perceptible (Zhang et al., 2024d); and 2) it is difficult to inject interaction records into the specific positions of user profiles, because historical interactions are usually introduced into prompts in chronological order (Zhang et al., 2024b).
-Furthermore, CheatAgent is an untargeted attack that undermines the model’s overall performance, which is beyond the scope of this paper, due to we focus on targeted attacks.
-In addition, as far as we know, there is no work investigating on textual attacks against the current agent-powered RSs.
-To tackle the above problems, we propose a practical attack framework, DrunkAgent, for the agent-based RSs with the aim of promoting target items.
-Our attack perturbs items’ descriptions to improve its effectiveness and imperceptibility, where the descriptions are optimized on black-box RSs to enhance its transferability.
-6. Conclusion
-In this paper, we propose a novel practical attack framework, DrunkAgent, for autonomous agent-powered RSs with powerful LLMs.
-DrunkAgent launches attacks by carefully crafting adversarial descriptions of the target items on agentic recommendations.
-In this way, the memories of the target agents can be effectively perturbed for achieving the attack objectives.
-Comprehensive experiments on real-world datasets show that DrunkAgent achieves state-of-the-art transferability and imperceptibility under the black-box settings.
-Most importantly, we identify a new safety vulnerability in the current agent4RSs.
-Our findings not only demonstrate an effective attack method, but also provide a lens to examine how memory mechanisms in agent-based recommenders can be influenced through real-world inputs.
-Our work can facilitate a deeper understanding of agent vulnerabilities and guide future efforts in building more robust and trustworthy recommendation agents.
-Since the generative agent4RS is rapidly flourishing in the recommendation community, we hope this work can be considered for building next-generation robust RSs in the future.
-Appendix A Appendix
-A.1. DrunkAgent Algorithm
-For reproducibility, the overall optimization procedure of DrunkAgent is given in Algorithm 1. For the sake of algorithmic transparency, our code will be made available after this work is published.
+[13] propose a novel black-box attack by leveraging the rewriting capabilities of ChatGPT, where attackers can significantly boost a target item’s exposure by merely altering its textual content during the testing.
+This attack overcomes this limitation that classical text attacks [34, 35, 36, 37] fail on LLM-based RSs due to misaligned malicious objectives.
+Compared with those traditional shilling attacks, such an item representation attack is notably stealthy [13], as it does not affect the overall recommendation performance, and it should be more efficient and lower in economic cost, due to it perturbs the item features directly and is not require to hire online writers to introduce additional fake user reviews.
+However, its transferability is limited, due to the attack is not optimized on any black-box RSs [41, 17].
+Moreover, since the attack appends positive words into the target item titles, the titles’ brevity and frequent exposure to users tend to make even small perturbations noticeable.
+Another black-box attack CheatAgent [14] has been proposed.
+This work claims the prompt templates and user profiles on the victim models can be arbitrarily perturbed.
+However, such assumptions are impractical.
+The main reasons are: 1) tampering with prompt templates may induce the overall performance changes, making the attack perceptible [13]; and 2) it is difficult to inject interaction records into the specific positions of user profiles, because historical interactions are usually introduced into prompts in chronological order [5].
+Furthermore, CheatAgent is an untargeted attack, which is beyond the scope of this paper, due to we focus on targeted attacks.
+Traditional RSs [46] are predominantly based on static models, where user preferences are learned from historical data in an offline manner and remain fixed during inference.
+These models do not adapt to evolving user behaviors or changing contexts after training.
+Vanilla LLM-based recommenders [15] follow a similar static paradigm: they generate recommendations through single-turn, stateless inference, relying on pre-defined prompt inputs to represent user preferences.
+In contrast, dynamic LLM-powered agentic recommenders [7, 11, 8] maintain internal states (e.g., memory mechanisms), allowing them to continuously adapt to and self-evolve over time based on sequential user interactions and simulated environmental feedback, even in offline experimental settings.
+This adaptivity enables personalized, context-aware, and temporally-sensitive recommendations.
+However, such dynamic mechanisms also introduce novel vulnerabilities that static models do not exhibit (e.g., memory corruption), which call for new robustness analyses tailored to the agent recommenders.
+While there are some attack methods on LLM-based agents [38], they often pursue goals that diverge from those relevant to RSs, leaving a gap in effective adversarial attacks for agent-driven recommenders.
+To tackle the above problems, in this paper, we provide the first work towards systematically investigate the security vulnerabilities in the agent-powered RSs.
+As such, we propose a practical black-box attack framework, DrunkAgent, for the LLM-powered agentic RSs with the aim of promoting target items.
+We perform stealthy memory corruptions by crafting effective and imperceptible textual descriptions of target items, due to the greater length and rich contextual information of descriptions can easily conceal adversarial perturbations compared to the titles.
+The adversarial descriptions are optimized via a surrogate model to enhance the transferability and stealthiness of DrunkAgent.
 Algorithm 1 Optimization Procedure of DrunkAgent
@@
 0:  the adversarial description of the target item $t$ Solve for the Optimal Adversarial Description Trigger on the Surrogate Module:
-1:  Obtain the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ with Eq. (2) - Eq. (5)
+1:  Obtain the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ with Eq. (2)
 […]
 3:  while $k\in\mathcal{E}$ epochs or the greedy algorithm does not converge (i.e. the attack goals are not achieved stably) do
-4:     Calculate $s_{i}^{(k-1)}$ for each input trigger $i$ with Eq. (7) - Eq. (8)
-5:     Maintain top-$n$ candidates in descending order of scores
-6:     Sample a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates with Eq. (9)
-7:     Randomly exchange of the slices between pairwise texts
-8:     Optimize and polish the |~$|\widetilde{\mathcal{M}}_{c}|-n$ combinations with Eq. (10)
-9:     Obtain |~$|\widetilde{\mathcal{M}}_{c}|$ triggers for the next epoch optimization
-10:  end while
-11:  Obtain the optimal description trigger $\widetilde{m_{t}}^{*}$ with Eq. (11) Solve for the Optimal Adversarial Perturbation Strategy on the Surrogate Module:
-12:  Define the adversarial strategy $\widetilde{q}$ with Eq. (12)
-13:  while the desired malicious action $\widetilde{a}_{t}$ does not occur (i.e., the strategy cannot get the target agents drunk) do
-14:     Optimize the arrangement of the defined strategies in $\widetilde{q}$ with Eq. (13)
-15:  end while
-16:  Obtain the optimal strategy $\widetilde{q}^{*}$ when Eq. (13) is maximum
-17:  return  the optimal adversarial description $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$
-Table 2. Statistics of Datasets
+4:     Quality Estimation Stage:
+5:     Calculate $s_{i}^{(k-1)}$ for each input trigger $i$ with Eq. (4)
+6:     Maintain top-$n$ candidates in descending order of scores
+7:     Feature Integration Stage:
+8:     Sample a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates via Softmax
+9:     Randomly exchange of the slices between pairwise texts
+10:     Linguistic Enrichment Stage:
+11:     Optimize and polish the |~$|\widetilde{\mathcal{M}}_{c}|-n$ previous combinations via LLMs
+12:     Obtain |~$|\widetilde{\mathcal{M}}_{c}|$ triggers for the next epoch optimization
+13:  end while
+14:  Obtain the optimal description trigger $\widetilde{m_{t}}^{*}$ with Eq. (5) Solve for the Optimal Adversarial Perturbation Strategy on the Surrogate Module:
+15:  Define the adversarial strategy $\widetilde{q}$ with Eq. (6)
+16:  while the desired malicious action $\widetilde{a}_{t}$ does not occur (i.e., the strategy cannot get the target agents drunk) do
+17:     Optimize the arrangement of the defined strategies in $\widetilde{q}$ with Eq. (7)
+18:  end while
+19:  Obtain the optimal strategy $\widetilde{q}^{*}$ when Eq. (7) is maximum
+20:  return  the optimal adversarial description $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$
+Appendix B Methodology
+B.1 Supplements to Basic Adversarial Prompt Template
+The main components of the prompt template are elaborately described in the following.
+  1. 1.
+    Style Definition $\widetilde{p_{s}}$.
+There are two common styles of prompts: task-oriented and role-oriented [47].
+To make recommendations more authentic and reasonable, we introduce role information into the context, which allows the system to respond with domain-specific knowledge, such as ‘You are a CD enthusiast.’, as shown in Fig.
+2.
+  2. 2.
+    Task Goal $\widetilde{p_{g}}$.
+Since a recommender system is essentially a ranker [46], we set sorting goals to indicate recommendation tasks, such as ‘Now, you are considering to sort ten candidate CDs in order of how well they align with your preferences.’ in the given example of the figure.
+  3. 3.
+    Recommendation Instruction $\widetilde{p_{i}}$.
+We introduce highly detailed instructions to guide the model to perform accurate recommendations.
+For example, ‘Please sort CDs by considering the correlation between your preferences/dislikes and the features of the CDs.
+The higher the CD rank, the more you like it.’ in the shown example.
+  4. 4.
+[…]
+B.2 DrunkAgent Algorithm
+For reproducibility, the overall optimization procedure of DrunkAgent is given in Algorithm 1.
+For the sake of algorithmic transparency, our code will be made available.
+Importantly, to mitigate potential misuse risks, the released code excludes scripts for automatic attack execution and is provided solely for academic research under a non-commercial license.
+All the users are required to agree to responsible use guidelines before accessing the whole repository.
+Appendix C Experiments
+C.1 Supplements to Experimental Setup
+Table 2: Statistics of Datasets
 Dataset #Users #Items #Interactions Sparsity
@@
 Musical Instruments 27,530 10,620 231,392 99.92%
-Table 3. Ablation Studies of DrunkAgent
-Dataset Component HR@1 HR@2 HR@3 NDCG@1 NDCG@2 NDCG@3
-$\mathcal{G}_{r}$ $\mathcal{S}_{t}$ $\mathcal{S}_{u}$
-CDs & Vinyl $\times$ 0.0808 0.1313 0.1919 0.0808 0.1127 0.1430
-$\times$ 0.0707 0.1212 0.2222 0.0707 0.1026 0.1531
-0.4040 0.4141 0.4343 0.4040 0.4104 0.4205
-Office Products $\times$ 0.1327 0.1429 0.1531 0.1327 0.1391 0.1442
-$\times$ 0.0612 0.0714 0.1122 0.0612 0.0677 0.0881
-0.2449 0.2449 0.2551 0.2449 0.2449 0.2500
-Musical Instruments $\times$ 0.1546 0.1649 0.1753 0.1546 0.1611 0.1663
-$\times$ 0.0309 0.0619 0.1237 0.0309 0.0504 0.0814
-0.2268 0.2268 0.2371 0.2268 0.2268 0.2320
-Table 4. Success Rates to Get the Agents Drunk
-Dataset #2 #5 #10 #20
-CDs & Vinyl 100% 100% 100% 100%
-Office Products 100% 100% 100% 100%
-Musical Instruments 100% 100% 100% 100%
-A.2. Dataset Statistics
+Table 3: Examples of baselines’ adversarial descriptions of target items on real-world datasets. The red part points out the differences from the original text.
+Attack Method The Adversarial Description of The Target Item $\mathcal{Q}$
+Benign (Default) The CD is called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Metal”.
+Character DeepwordBug Te hCD is called “Counterparts”. The category of this CD is: “Rcok; Progressie; Progressive Metal”. 34
+-level PuncAttack The CD i-s called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Meta-l”. 45
+Word TextFooler The CDS is titled “Counterparts”. The category of this CD is: “Rock; Gradually; Progressive Metals”. 55
+-level BertAttack The CD is called “Counterparts”. The category of each CD is: “Rock; Progressive; Progressive Metal”. 23
+TrivialInsertion The CD is called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Metal”. 100
+Sentence amazing !!!
+-level ChatGPTAttack “Counterparts” CD: Perfect blend of rock, progressive, and metal. Experience the wonderful 100
+sound that will elevate your music collection.
+C.1.1 Dataset Statistics
 Table 2 illustrates the statistics of these datasets.
 The datasets vary in size and sparsity, which is suitable for a comprehensive evaluation.
-Following the work (Zhang et al., 2024b), we further sample subsets from the datasets given the expensive API calls.
-Specifically, considering the importance and commonness of data sparsity and cold-start in the recommendation community, we randomly sample 100 users with 800 interactions (777 items), 99 users with 693 interactions (619 items) and 98 users with 588 interactions (483 items) from the three datasets, respectively, allowing to explore more diverse and practical interaction scenarios.
-Moreover, the leave-one-out strategy (He et al., 2017) is used for evaluation, where the split between the training and test set is 9:1.
+Following the work [5], we further sample subsets from the datasets given the expensive API calls.
+Specifically, considering the importance and commonness of data sparsity and cold-start in the recommendation community, we randomly sample 100 users with 800 interactions including 777 items, 99 users with 693 interactions including 619 items and 98 users with 588 interactions including 483 items from the three datasets, respectively, allowing to explore more diverse and practical interaction scenarios.
+Moreover, the leave-one-out strategy [44] is used for evaluation, where the split between the training and test set is 9:1.
 Scaling DrunkAgent for larger datasets is left as future work.
-A.3. Ablation Studies
-We remove important components of each module to conduct ablation studies of DrunkAgent.
-First, we consider remove the surrogate module (denoted as $\mathcal{S}_{u}$).
-However, if we remove it, the setting will transfer from the black-box to the white-box that is impractical and beyond the scope of this paper.
-Moreover, due to gpt-3.5-turbo used for victim models is closed-source and only can be called via APIs, we cannot obtain logits to optimize the adversarial triggers like Llama-3 used in the surrogate model does.
-Hence, we maintain the surrogate model to ensure more practical settings.
-We then remove the greedy optimization algorithm (denoted as $\mathcal{G}_{r}$) in the generation module to observe how the performance changes.
-As illustrated in Table  3, the attack performance moderately degrades after removing $\mathcal{G}_{r}$, which demonstrates the effectiveness of the optimization algorithm.
-The lack of dramatic performance degradation is probably due to the effectiveness of our initialization.
-Finally, we mask the strategy module (denoted as $\mathcal{S}_{t}$).
-From the table, we can find that the attack performance decreases by a large margin.
-This may be because memory updates during agent-environment interactions are somewhat robust to non-tailored attacks (this phenomenon can also be observed in the performance of the attack baselines).
-Our reported results are built upon AgentCF, and similar tendencies can be found on AgentRAG and AgentSEQ.
-A.4. Parameter Sensitivity
-From Table  3, we can find that the strategy module contributes more than the optimization algorithm on DrunkAgent’s attack performance.
-As such, we increase the maximum rounds of optimization iterations per step of victim models to evaluate the attack success rates of the attack strategy, where the default value is 2 as suggested in the paper (Zhang et al., 2024b).
-As illustrated in Table  4, the success rates of getting the target agents drunk are maintained at 100% regardless of the number of iterations.
-For the other attack parameters, based on our observations, our DrunkAgent is insensitive to them.
-A.5. Victim LLM-powered Agentic Models
-LLM-driven agents, known for their superior autonomous interaction and decision-making capabilities, are gradually being considered as next-generation RSs.
-Compared with some agent-powered RSs that focus solely on user-side behavior modeling using universal LLMs (Wang et al., 2024c; Huang et al., 2024), AgentCF (Zhang et al., 2024b) emphasizes the modeling of two-sided interaction relations between user agents and item agents through the idea of collaborative filtering (CF) (He et al., 2017).
-Both kinds of agents are equipped with memory modules, maintaining the simulated preferences and tastes of potential adopters involving their intrinsic features and acquired behavioral information via autonomous interactions and collaborative reflection optimizations at each time step.
+C.1.2 Baseline Methods
+The baselines perturb the descriptions of the target items at various levels, which are described as below.
+  • Character-level Perturbations. DeepwordBug [34] and PuncAttack [35] manipulate texts by introducing typos and inserting punctuation, respectively.
+  • Word-level Perturbations. TextFooler [36] and BertAttack [37], respectively, aim to replace words with synonyms or contextually similar words.
+  • Sentence-level Perturbations.
+TrivialInsertion [13] inserts text content with $k$ number of positive words from a pre-defined word corpus, where the insertion is conducted at the end of the text.
+Here, $k=2$.
+ChatGPTAttack [13] increases the stealthiness of TrivialInsertion by rewriting the text content via GPTs, where GPT-3.5-turbo is used as the original paper recommends and the maximum number of words is set to 30 here.
+We give some examples of adversarial descriptions of target items generated by the baselines in Table 3.
+Following the works [5, 7], the item memory of the agent can be initialized by their identity information, such as titles and categories, as shown in the benign (default) descriptions of the tables.
+Although the adversarial description of the target item generated from DrunkAgent is different from the benign status’s (see Fig.
+2 and Table 3), such modifications are feasible due to the real-world item providers are allowed to modify their product representations via APIs [17, 48, 20], as discussed in Section 1, especially when the benign description is just a default exemplar used for experiments.
+Moreover, compared with the number of queries (denoted as $\mathcal{Q}$ in the table) of the baselines, DrunkAgent has the lowest attack costs due to $\mathcal{E}$=20.
+The implementation and parameters of the attack baselines follow the open-source works: StealthyAttack [^3] https://github.com/CRIPAC-DIG/RecTextAttack [13] and PromptBench [^4] https://github.com/microsoft/promptbench [49].
+C.1.3 Victim LLM-powered Agentic Recommender Systems
+LLM-driven agents, known for their superior autonomous interaction and decision-making capabilities, are gradually being considered as next-generation RSs [11].
+Compared with some agent-powered RSs that focus solely on user-side behavior modeling using universal LLMs [50, 7], AgentCF [5] emphasizes the modeling of two-sided interaction relations between user agents and item agents through the idea of collaborative filtering [44].
+Both kinds of agents are equipped with memory modules, maintaining the simulated preferences and tastes of potential adopters involving their intrinsic features and acquired behavioral information via autonomous interactions and collaborative optimizations at each time step, as shown on the left of Fig.
+1.
 As such, the memories of the user agents and item agents can be accordingly updated, enabling the agents to better fit the real-world interaction behaviors.
-To conduct personalized recommendations, AgentCF introduces various well-crafted prompting strategies on the optimized memories.
-We name the model according to the use of the prompting strategy, with the basic one called AgentCF, the advanced retrieval-augmented one called AgentRAG and the sequential one called AgentSEQ.
-The prompt implementation and parameters of collaborative agent optimizations can be found in (Zhang et al., 2024b).
-For the prompting strategies, we have made some improvements for high-quality recommendations, such as Chain-of-Thought and Position Bias are introduced into the prompt template $\mathcal{P}$.
-The strategies are described below.
+To better adapt to the dynamically changing real-world environment, the agentic RSs are equipped with real-time capabilities.
+Whenever new descriptions are introduced at a time step, they are integrated into the evolving memories to maintain up-to-date contextual understanding.
+As can be seen on the right of Fig.
+1, the recommendation agents conduct personalized recommendations based on the optimized memories.
+To drive varied architectural designs and support diverse tasks for a comprehensive evaluation of the transferability of attacks, powerful mechanisms (e.g., retrieval augmentation) and task-specific prompts (e.g., sequential recommendation) are introduced to strengthen the agentic RSs.
+As such, AgentRAG and AgentSEQ are developed [5].
+To maintain a fair evaluation, the implementation details of how the agents can accomplish autonomous interactions and collaborative optimizations are followed the original paper suggest [5], where gpt-3.5-turbo [^5] https://platform.openai.com/docs/overview is recommended to achieve the backbone LLMs.
+Specifically, such backbone delivers strong natural language understanding and generation capabilities that are sufficient for most recommendation-related tasks.
+In addition, it is cost-efficient, making it well-suited for multi-turn interactions where cost can quickly accumulate.
+Moreover, it has lower latency, supporting real-time responsiveness in interactive agent scenarios.
+To ensure the high-quality of recommendations, we enhance the prompting strategies for recommendation agents.
+Specifically, we introduce the Chain-of-Thought strategy [51] to carefully craft the prompt template $\mathcal{P}$.
+Given the context window constraints pointed out in [5], a target item and nine non-interactive items of each normal user are randomly selected for the ranking task of the RSs, which is also suggested by the original work.
+The design of recommendation agents is described below.
   • AgentCF.
@@
 For each normal user $u$,
-    (14) $\mathcal{R}_{u}=f_{\text{AgentCF}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
-    where $m_{u}$ is the short-memory of user agent $u$, $m_{t}$ is the features of the target item agent, $\mathcal{C}_{t}^{u}$ is the feature set of the candidate item agents and $\mathcal{R}_{u}$ is the ranking result.
+    $\mathcal{R}_{u}=f_{\text{AgentCF}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$ (8)
+    where $m_{u}$ is the short-memory of user agent $u$, $m_{t}$ denotes the features of the target item agent including metadata and the memories, $\mathcal{C}_{t}^{u}$ is the feature set of the candidate item agents and $\mathcal{R}_{u}$ is the ranking result.
   • AgentRAG.
@@
 For each normal user $u$,
-    (15) $\mathcal{R}_{u}=f_{\text{AgentRAG}}(\mathcal{P}\oplus m_{u}\oplus m_{r}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
+    $\mathcal{R}_{u}=f_{\text{AgentRAG}}(\mathcal{P}\oplus m_{u}\oplus m_{r}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$ (9)
     where $m_{r}$ is retrieved specialized preference from the long-term memory of user agent $u$ by taking the memories of candidate items as queries.
-To ensure the quality while maintaining the efficiency, we use all-MiniLM-L6-v2555https://www.sbert.net/docs/sentence_transformer/pretrained_models.html as the sentence transformer for the similarity calculation.
+To ensure the quality while maintaining the efficiency, we use all-MiniLM-L6-v2666https://www.sbert.net/docs/sentence_transformer/pretrained_models.html as the sentence transformer for the similarity calculation.
   • AgentSEQ.
@@
 For each normal user $u$,
-    (16) $\mathcal{R}_{u}=f_{\text{AgentSEQ}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{M}_{t}^{u}\oplus\mathcal{C}_{t}^{u}),$
+    $\mathcal{R}_{u}=f_{\text{AgentSEQ}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{M}_{t}^{u}\oplus\mathcal{C}_{t}^{u}),$ (10)
     where $\mathcal{M}_{t}^{u}$ are the corresponding item features of the historical interactions with user $u$.
-We give some examples of the prompting templates of three victim models in Table  5 for reproductions.
-Table 5.
-Examples of prompting templates of the LLM-based agentic victim models.
+We give some examples of the prompting templates of three victim models in Table 4 for reproductions.
+Table 4: Examples of prompting templates of the LLM-based agentic victim models.
 The blue italics represent variables such as the memories of user agents and item agents optimized via agent-environment interactions.
-The bond fonts indicate the Chain-of-Thought and Position Bias enhancement strategies.
+The bond fonts indicate Chain-of-Thought enhancement strategy.
 Victim Model Recommendation Prompting Template Example
@@
 10. [Title of the least favorite candidate CD]”}
-Table 6. Examples of baselines’ adversarial descriptions of target items on real-world datasets. The red part points out the differences from the original text.
-Attack Method The Adversarial Description of The Target Item $\mathcal{Q}$
-Benign (Default) The CD is called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Metal”.
-Character DeepwordBug Te hCD is called “Counterparts”. The category of this CD is: “Rcok; Progressie; Progressive Metal”. 34
--level PuncAttack The CD i-s called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Meta-l”. 45
-Word TextFooler The CDS is titled “Counterparts”. The category of this CD is: “Rock; Gradually; Progressive Metals”. 55
--level BertAttack The CD is called “Counterparts”. The category of each CD is: “Rock; Progressive; Progressive Metal”. 23
-TrivialInsertion The CD is called “Counterparts”. The category of this CD is: “Rock; Progressive; Progressive Metal”. 100
-Sentence amazing !!!
--level ChatGPTAttack “Counterparts” CD: Perfect blend of rock, progressive, and metal. Experience the wonderful 100
-sound that will elevate your music collection.
-A.6. Baselines
-The baselines perturb the descriptions of the target items at various levels, which are described as below.
-  • Character-level Perturbations. DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023) manipulate texts by introducing typos and inserting punctuation, respectively.
-  • Word-level Perturbations. TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020), respectively, aim to replace words with synonyms or contextually similar words.
-  • Sentence-level Perturbations.
-TrivialInsertion (Zhang et al., 2024d) inserts text content with $k$ number of positive words from a pre-defined word corpus, where the insertion is conducted at the end of the text.
-Here, $k=2$.
-ChatGPTAttack (Zhang et al., 2024d) increases the stealthiness of TrivialInsertion by rewriting the text content via GPTs.
-GPT-3.5-turbo is used and the maximum number of words is set to 30 here.
-We give some examples of adversarial descriptions of target items generated by the baselines in Table  6.
-Following the works (Zhang et al., 2024b; Huang et al., 2024), the item memory of the agent can be initialized by their identity information, such as titles and categories, as shown in the benign (default) descriptions of the tables.
-Although the adversarial description of the target item generated from DrunkAgent is different from the benign status’s (see Fig.
-2 and Table  6), such modifications are feasible due to the real-world item providers is allowed to modify their product representations via APIs (Yang et al., 2024; Cohen et al., 2021; Liu and Larson, 2021), as discussed in Section 1, especially when the benign description is just a default exemplar used for experiments.
-Moreover, compared with the number of queries (denoted as $\mathcal{Q}$ in the table) of the baselines, DrunkAgent has the lowest attack costs due to $\mathcal{E}$=20.
-The implementation and parameters of the baselines follow the open-source works: StealthyAttack [^6] https://github.com/CRIPAC-DIG/RecTextAttack (Zhang et al., 2024d) and PromptBench [^7] https://github.com/microsoft/promptbench (Zhu et al., 2023b).
+C.1.4 Configuration of DrunkAgent
+As discussed in Section 3.1, we apply the latest open-source model LLaMA 3 for building the LLM backbone of the surrogate model $f_{\text{Agent}_{\Phi}}(\cdot)$.
+Its collection spans different sizes, from 8B, 70B to 405B.
+To enhance the efficiency and ensure the quality of the attack, we use the lightweight version: meta-llama/Meta-Llama-3-8B-Instruct [^7] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct.
+OpenAI’s gpt-4-turbo [^8] https://platform.openai.com/docs/models/gpt-4-turbo [52] is adopted to implement $f_{\text{LLM}_{\Psi}}(\cdot)$, because it is a state-of-the-art model with powerful language understanding, reasoning and generation capabilities.
+We set $\mathcal{E}=20$, |~$|\widetilde{\mathcal{M}}_{c}|=10$ and $n=5$.
+All random seeds in the evaluation are set to 2024 and the victim LLMs’ temperature is set to 0 for reproducibility.
+For all the other parameters, we keep the default values.
+C.1.5 Evaluation Metrics
+We use two widely-used ranking metrics to evaluate the attack transferability and stealthiness: hit ratio (HR@$\mathcal{K}\uparrow$) and normalized discounted cumulative gain (NDCG@$\mathcal{K}\uparrow$) [53, 19].
+HR@$\mathcal{K}$ measures whether the target/test item is appeared in the top-$\mathcal{K}$ recommendation list of the normal users, while NDCG@$\mathcal{K}$ indicates the ranking position of the target/test item in the list.
+As such, a high HR@$\mathcal{K}$ and a NDCG@$\mathcal{K}$ demonstrate high attack effectiveness, and vice versa.
+To clearly show the performance gap among different attacks, we set $\mathcal{K}$ to 1, 2, 3.
+We also introduce standard sentence perplexity score $\downarrow$ [13] that is a commonly-adopted metric to evaluate the attack imperceptibility.
+The perplexity measures a language model’s uncertainty by taking the exponential of the average negative log-likelihood of the words in a sequence, given their preceding context.
+Text with low perplexity is more fluent and natural, making it harder to detect as manipulated or adversarial, while text with high perplexity is more likely to be detected as unusual or abnormal.
+C.2 Supplements to Performance Comparison
+C.2.1 Attack Universality across Different Target Items.
+In addition to cross-model transferability, we evaluate the cross-sample transferability of DrunkAgent.
+We incorporate the generated adversarial descriptions into totally different target items that are randomly sampled from each dataset.
+HR@1 (i.e., NDCG@1) of DrunkAgent across different models and target items on real-world datasets are shown in Fig.
+6, which indicates its commendable transferability and triggers’ universality across different target items.
+This may be attributed to the fact that our triggers encapsulate diverse and general item characteristics, and, crucially, they also reveal the intended attack objectives.
+Figure 6: Attack Universality across Different Target Items
+Table 5: Ablation Studies of DrunkAgent
+Dataset Component HR@1 HR@2 HR@3 NDCG@1 NDCG@2 NDCG@3
+$\mathcal{G}_{r}$ $\mathcal{S}_{t}$ $\mathcal{S}_{u}$
+CDs & Vinyl $\times$ 0.0808 0.1313 0.1919 0.0808 0.1127 0.1430
+$\times$ 0.0707 0.1212 0.2222 0.0707 0.1026 0.1531
+0.4040 0.4141 0.4343 0.4040 0.4104 0.4205
+Office Products $\times$ 0.1327 0.1429 0.1531 0.1327 0.1391 0.1442
+$\times$ 0.0612 0.0714 0.1122 0.0612 0.0677 0.0881
+0.2449 0.2449 0.2551 0.2449 0.2449 0.2500
+Musical Instruments $\times$ 0.1546 0.1649 0.1753 0.1546 0.1611 0.1663
+$\times$ 0.0309 0.0619 0.1237 0.0309 0.0504 0.0814
+0.2268 0.2268 0.2371 0.2268 0.2268 0.2320
+Table 6: Success Rates to Get the Target Agents Drunk
+Dataset #2 #5 #10 #20
+CDs & Vinyl 100% 100% 100% 100%
+Office Products 100% 100% 100% 100%
+Musical Instruments 100% 100% 100% 100%
+C.2.2 Ablation Studies
+We remove important components of each module to conduct ablation studies of DrunkAgent.
+First, we consider remove the surrogate module (denoted as $\mathcal{S}_{u}$).
+However, if we remove it, the setting will transfer from the black-box to the white-box that is impractical and beyond the scope of this paper.
+Moreover, due to gpt-3.5-turbo used for victim models is closed-source and only can be called via APIs, we cannot obtain logits to optimize the adversarial triggers like Llama-3 used in the surrogate model does.
+Hence, we maintain the surrogate model to ensure more practical settings.
+We then remove the greedy optimization algorithm (denoted as $\mathcal{G}_{r}$) in the generation module to observe how the performance changes.
+As illustrated in Table 5, the attack performance moderately degrades after removing $\mathcal{G}_{r}$, which demonstrates the effectiveness of the optimization algorithm.
+The lack of dramatic performance degradation is probably due to the effectiveness of our initialization.
+Finally, we mask the strategy module (denoted as $\mathcal{S}_{t}$).
+From the table, we can find that the attack performance decreases by a large margin.
+This may be because memory updates during agent-environment interactions are somewhat robust to non-tailored attacks (this phenomenon can also be observed in the performance of the attack baselines).
+Our reported results are built upon AgentCF, and similar tendencies can be found on AgentRAG and AgentSEQ.
+C.2.3 Parameter Sensitivity
+From Table 5, we can find that the strategy module contributes more than the optimization algorithm on DrunkAgent’s attack performance.
+As such, we increase the maximum rounds of optimization iterations per step of victim models to evaluate the attack success rates of the attack strategy, where the default value is 2 as suggested in the paper [5].
+As illustrated in Table 6, the success rates of getting the target agents drunk are maintained at 100% regardless of the number of iterations.
+For the other attack parameters, based on our observations, our DrunkAgent is insensitive to them.
+C.2.4 Attack Transferability to Victim Agentic RSs with Powerful LLM backbones
+To further evaluate the transferability of DrunkAgent, we improve the LLM backbones of victim agentic RSs with more powerful gpt-4-turbo.
+Due to such API calls are quite expensive, we evaluate DrunkAgent on a representative real-world dataset.
+To ensure a sufficiently challenging evaluation, we choose the largest dataset: CD & Vinyl.
+Fig.
+7 shows HR@1 (i.e., NDCG@1) of the benign status and DrunkAgent against the three victim models on the dataset, which demonstrates that DrunkAgent is still transferable even if the agentic RSs become powerful.
+We leave experiments on additional datasets and more powerful agentic RSs for future work.
+Figure 7: Attack Transferability on Powerful Victim Agentic RSs.
+(a) AgentCF
+(b) AgentRAG
+(c) AgentSEQ
+Figure 8: Attack Transferability.
+HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ of the Benign status and DrunkAgent against various black-box victim LLM-powered agent-based RSs on the real-world Yelp dataset.
+C.2.5 Attack Generalizability across Domains
+To further assess the generalizability of DrunkAgent, we conduct additional experiments on the Yelp dataset.
+Compared to the e-commerce-focused Amazon dataset, Yelp centers around user reviews of local businesses, including restaurants, cafes, bars, salons, repair shops, and other service-oriented establishments.
+It places a strong emphasis on experiential content, where user reviews play a central role in shaping recommendations.
+This domain is inherently different from traditional product recommendation: it often involves more subjective preferences and temporal dynamics (e.g., trends in dining or local services).
+These characteristics make Yelp a complementary and challenging testbed for evaluating the transferability and effectiveness of our method on different domains.
+Due to the expensive API calls, 772 interactions are randomly sample as a subset, including 61 users and 552 items.
+The sparsity of the subset is 97.71%, which is more dense than the e-commerce ones.
+Despite these domain-specific differences, as can be seen from Fig.
+8, DrunkAgent remains highly effective: it consistently promotes the target items on black-box LLM-powered recommender agents under memory corruptions, confirming the transferability of our attack and the effectiveness of memory perturbations across domains.
+This cross-domain success highlights the adaptability and generalizability of DrunkAgent, demonstrating that its impact is not limited to e-commerce settings but extends to real-world platforms driven by experiential content and local service discovery.
+C.2.6 Attack Interpretability
+We focus on why adversarial attacks exhibit both transferability and stealthiness.
+The transferability arises because adversarial texts are optimized to carry rich as well as general features and explicit attack objectives (e.g., ‘top-level consideration’, ‘a must-have’, ‘prime choice’) to mislead the decisions of the victim models, making them effective across models that share similar language understanding capabilities.
+The stealthiness is attributed to the fact that these texts are optimized to retain meaningful semantics and coherence, and mainly perturb the memories of the target item agents.
+C.2.7 Computing Resourses
+All experiments were run on a institution high-performance computing cluster equipped with NVIDIA H100 GPUs (each with 96GB VRAM), using CUDA version 12.4 and NVIDIA driver version 550.127.05.
+Each individual experimental run (from initialization to evaluation) completed in under 2 hours, with peak GPU memory usage remaining below 30GB.
+These resource requirements apply across all datasets, methods, and target systems evaluated in the paper.
+Appendix D Limitations
+While our work demonstrates the effectiveness of memory-based attacks against agentic recommender systems in controlled experimental settings, there are several limitations to be acknowledged.
+Our evaluation is conducted on publicly available benchmark datasets and simulated agent environments.
+Although these setups capture key characteristics of real-world agentic systems, they may not fully reflect the complexity and scale of industrial deployments.
+In particular, factors such as heterogeneous user behaviors, evolving content ecosystems, and industrial-scale memory management pose additional challenges that are beyond the scope of this study.
+Future work is needed to validate the generalizability and scalability of our approach in production environments.
+In addition, our study focuses on demonstrating the attack effectiveness under a black-box threat model, assuming limited access to internal model parameters or memory states.
+While this setting reflects practical constraints faced by external attackers, a deeper exploration of defense mechanisms and potential mitigation strategies is left for future investigation.
+Despite these limitations, we believe our findings offer valuable insights into the security risks posed by memory mechanisms in agentic recommender systems, and lay a foundation for further research on robust and trustworthy recommendation agents.
+Appendix E Broader Impact
+As agentic recommender systems are emerging as a promising paradigm for next-generation personalization [11, 8], our work contributes positively by proactively identifying potential security vulnerabilities in these systems.
+While such vulnerabilities could lead to potential negative societal impacts if misused, our research serves as a foundational step toward building more robust and trustworthy agentic RSs, enabling early risk awareness and informing effective mitigation strategies, thereby helping to minimize negative societal impacts before real-world deployment.
+Overall, we believe the positive impact of promoting secure and responsible development in this emerging area outweighs the potential risks.

### v3

@@
 DrunkAgent: Stealthy Memory Corruption in LLM-Powered Recommender Agents
-Abstract
+Abstract.
 Large language model (LLM)-powered agents are increasingly used in recommender systems (RSs) to achieve personalized behavior modeling, where the memory mechanism plays a pivotal role in enabling the agents to autonomously explore, learn and self-evolve from real-world interactions.
@@
 Despite its central role, the robustness of agentic RSs in the face of such threats remains largely underexplored.
-Previous works suffer from semantic mismatches or rely on static embeddings or pre-defined prompts, all of which hinder their applicability to systems with dynamic memory states.
-This challenge is exacerbated by the black-box nature of commercial RSs.
+Previous works suffer from semantic mismatches or rely on static embeddings or pre-defined prompts, all of which are not designed for dynamic systems, especially for dynamic memory states of LLM agents.
+This challenge is exacerbated by the black-box nature of commercial recommenders.
 To tackle the above problems, in this paper, we present the first systematic investigation of memory-based vulnerabilities in LLM-powered recommender agents, revealing their security limitations and guiding efforts to strengthen system resilience and trustworthiness.
@@
 Extensive experiments on real-world datasets across diverse agentic RSs, including collaborative filtering, retrieval augmentation and sequential recommendations, demonstrate the generalizability, transferability and stealthiness of DrunkAgent.
-1 Introduction
+Recommender Systems, Adversarial Attacks, Generative Agents, Large Language Models, Collaborative Filtering
+copyright: noneccs: Information systems Recommender systemsccs: Security and privacy Web application security
+Figure 1.
+In LLM-powered agentic RSs, agents autonomously interact with their situated environments including other agents to collaboratively optimize their memories over time (Left), with recommendations being generated based on these memories (Right).
+DrunkAgent perturbs the memory of the target item agent, where the attack strategy aims to ‘get the target agent drunk’ so that the well-designed adversarial triggers can be injected into the target memory (The Temporal Snapshot of the Left).
+As a result, the target item can be promoted to more recommendation lists of normal users (Right).
+1. Introduction
 The advent of large language models (LLMs) has ushered a transformative paradigm in recommender systems (RSs).
-Early LLM-based RSs primarily relied on verbalizing user-item interactions into prompts to guide recommendations (geng2022recommendation, ; bao2023tallrec, ; zhang2023collm, ; wu2024coral, ).
-However, such approaches often struggle to capture personalized behavioral patterns, due to a fundamental gap between interaction dynamics and generic language modeling (zhang2024agentcf, ).
-To bridge this gap, recent works have introduced autonomous agents built upon LLMs and augmented with modules, such as memory, planning, and tool-use, laying the foundation for agentic RSs (durante2024agent, ; huang2024foundation, ).
-As personalized user experiences become central to digital platforms, LLM-powered agentic RSs are rapidly emerging as the next evolutionary step (huang2025towards, ), yet their robustness, particularly against adversarial threats, remains largely underexplored.
+Early LLM-based RSs primarily relied on verbalizing user-item interactions into prompts to guide recommendations (Geng et al., 2022; Bao et al., 2023; Zhang et al., 2023; Wu et al., 2024).
+However, such approaches often struggle to capture personalized behavioral patterns, due to a fundamental gap between interaction dynamics and generic language modeling (Zhang et al., 2024b).
+To bridge this gap, recent works have introduced autonomous agents built upon LLMs and augmented with modules, such as memory, planning, and tool-use, laying the foundation for agentic RSs (Durante et al., 2024; Huang et al., 2024).
+As personalized user experiences become central to digital platforms, LLM-powered agentic RSs are rapidly emerging as the next evolutionary step (Huang et al., 2025b), yet their robustness, particularly against adversarial threats, remains largely underexplored.
 At the core of agentic systems lies the memory module, which plays a pivotal role in supporting agent-environment [^1] In a narrow sense, environment is the object that the agent needs to interact with to accomplish the task.
-More broadly, environment can be any contextual factors that influence the agent’s decisions (zhang2024survey, ).
-interactions (zhang2024survey, ; wang2024survey, ).
-Specifically, the module retains historical interactions, allowing agents to accumulate contextual knowledge, adapt to evolving user preferences, and refine their behavior over time through interaction-driven updates (huang2025survey, ), as shown in Fig.
+More broadly, environment can be any contextual factors that influence the agent’s decisions (Zhang et al., 2024a).
+interactions (Zhang et al., 2024a; Wang et al., 2024b).
+Specifically, the module retains historical interactions, allowing agents to accumulate contextual knowledge, adapt to evolving user preferences, and refine their behavior over time through interaction-driven updates (Huang et al., 2025a), as shown in Fig.
 1.
@@
 Adversarial attacks targeting memory often leave a lasting impact.
-Given that attacks are generally persistent rather than one-off events, they can be readily exploited in non-stationary environments to take advantage of the agent’s continual learning process (huang2025towards, ; zhang2024agentcf, ), leading to repeated integration of adversarial signals into its evolving internal state.
+Given that attacks are generally persistent rather than one-off events, they can be readily exploited in non-stationary environments to take advantage of the agent’s continual learning process (Huang et al., 2025b; Zhang et al., 2024b), leading to repeated integration of adversarial signals into its evolving internal state.
 This results in long-term memory contamination and hence strengthens a drift in user preferences.
-Under sequential and retrieval modeling (huang2024foundation, ; huang2025survey, ), these perturbations persist longer by influencing the agent’s ongoing behavior trajectory.
+Under sequential and retrieval modeling (Huang et al., 2024, 2025a), these perturbations persist longer by influencing the agent’s ongoing behavior trajectory.
 Perturbed memory traces can gradually bias the agent’s belief state through temporal dependencies, while also being frequently retrieved in future recommendations.
@@
 In this work, we take the first step toward enhancing the robustness and security of the LLM-powered agentic RSs by systematically investigating their adversarial vulnerabilities, with a particular focus on stealthy memory corruption attacks.
-Figure 1: In LLM-powered agentic RSs, agents autonomously interact with their situated environments including other agents to collaboratively optimize their memories over time (Left), with recommendations being generated based on these memories (Right).
-DrunkAgent perturbs the memory of the target item agent, where the attack strategy aims to ‘get the target agent drunk’ so that the adversarial triggers can be injected into the target memory (The Temporal Snapshot of the Left).
-As a result, the target item can be promoted to more recommendation lists of normal users (Right).
-Existing works on attacks against RSs falls into three largely independent threads: 1) poisoning attacks on traditional collaborative filtering (CF) models (wang2024poisoning, ), which inject fake user profiles to corrupt embeddings; 2) adversarial text attacks on LLM-based recommenders (zhang2024stealthy, ; ning2024cheatagent, ), perturbing input prompts to misguide outputs; and 3) agentic RS architectures that enhance personalization through memory-augmented LLMs (huang2025survey, ).
+Existing works on attacks against RSs falls into three largely independent threads: 1) poisoning attacks on traditional collaborative filtering (CF) models (Wang et al., 2024a), which inject fake user profiles to corrupt embeddings; 2) adversarial text attacks on LLM-based recommenders (Zhang et al., 2024c; Ning et al., 2024), perturbing input prompts to misguide outputs; and 3) agentic RS architectures that enhance personalization through memory-augmented LLMs (Huang et al., 2025a).
 However, these threads remain disconnected.
 Poisoning methods often assume static embeddings and focus on numerical perturbations, creating a semantic gap with text-driven systems.
-Textual attacks tend to focus narrowly on pre-defined prompts, limiting their applicability in agentic RSs with dynamic memory updates.
+Textual attacks tend to focus narrowly on pre-defined prompts, limiting their applicability in agentic RSs, where dynamic memory updates can mitigate the impact of these static attacks to a certain extent.
 Meanwhile, the agentic studies neglect security analyses, leaving critical vulnerabilities unexplored.
 This challenge is exacerbated by the black-box nature of commercial RSs, where attackers lack access to model parameters and training pipelines.
-Moreover, a critical threat vector aligned with real-world platform abuse incentives lies in item promotion attacks.
-Poisoning attacks rely on fake user injections, making promoting items costly and indirect, while textual attacks typically focus on degrading overall performance or causing mis-classification, overlooking practical promotion-driven threats (wang2024poisoning, ; ning2024cheatagent, ; zhang2024stealthy, ).
+Moreover, a critical threat vector aligned with real-world platform abuse (e.g., Amazon, eBay, Sony (Lin et al., 2020, 2022)) incentives lies in item promotion attacks.
+Poisoning attacks rely on fake user injections, making promoting items costly and indirect, while textual attacks typically focus on degrading overall performance or causing mis-classification, overlooking practical promotion-driven threats (Wang et al., 2024a; Ning et al., 2024; Zhang et al., 2024c).
 To tackle the above problems, in this paper, we propose DrunkAgent, the first black-box framework to exploit memory-based vulnerabilities in LLM-powered recommender agents.
@@
 1.
-To mitigate the risk of raising suspicion from frequent queries on victim models in prior methods (zhang2024stealthy, ; ning2024cheatagent, ; wang2024poisoning, ) and to adhere to realistic black-box constraints, DrunkAgent introduces a surrogate model to optimize the triggers and strategies to further improve the attack transferability and stealthiness.
+To mitigate the risk of raising suspicion from frequent queries on victim models in prior methods (Zhang et al., 2024c; Ning et al., 2024; Wang et al., 2024a) and to adhere to realistic black-box constraints, DrunkAgent introduces a surrogate model to optimize the triggers and strategies to further improve the attack transferability and stealthiness.
 To comprehensively evaluate DrunkAgent, we conduct extensive experiments across real-world datasets from different domains, a variety of attack baselines with varying perturbation strengths, representative black-box agentic RSs featuring diverse designs and recommendation tasks, including collaborative filtering, retrieval augmentation, and sequential recommendations, widely adopted stealthiness evaluation methods, and cutting-edge defense mechanisms.
 Our main contributions are summarized as follows:
-  • We identify a safety vulnerability in autonomous agent-based RSs with powerful LLMs.
-To the best of our knowledge, we present the first formal and systematic study of adversarial textual attacks on the agentic RSs, aiming to inform future robustness research.
-  • We propose DrunkAgent, a novel black-box attack framework specifically designed for the agent-powered RSs, which crafts adversarial triggers and tailored strategies for effective memory perturbations of target item agents.
+  • We identify a security vulnerability in autonomous, LLM-powered agent-based RSs; specifically, the agent’s memory constitutes a primary attack surface.
+To the best of our knowledge, we present the first systematic study of adversarial textual attacks on the agentic RSs, aiming to inform future robustness research.
+  • We propose DrunkAgent, a novel black-box attack framework tailored for agent-driven RSs.
+Unlike previous attack methods that only focus on static systems, it effectively perturbs the dynamic memory of the target item agent by crafting adversarial triggers and customized strategies.
   • DrunkAgent consistently outperforms state-of-the-art black-box attacks against a wide range of agentic RSs and under multiple stealthiness evaluation methods on real-world datasets, in terms of both transferability and stealthiness.
 Moreover, it remains resistant to existing advanced defense mechanisms, exposing critical blind spots in current countermeasures and underscoring the urgent need for adaptive defenses.
-2 Problem Formulation
-2.1 Victim LLM-powered Agentic RSs
+2. Problem Formulation
+2.1. Victim LLM-powered Agentic RSs
 We use $\mathcal{Y}=\{y_{u,v}:u\in\mathcal{U},v\in\mathcal{V}\}$ to denote the records of the user-item interaction matrix in the recommendation space, where $\mathcal{U}$ and $\mathcal{V}$ represent the set of real users and the item universe, respectively.
 $\mathcal{V}_{u}=\{v\in\mathcal{V}:y_{u,v}\neq 0\}$ indicates the set of items that have been interacted by $u$ (i.e., the user’s profile).
-LLM-powered RSs generally conduct recommendations by integrating user historical behavior sequences and/or item and/or user features with personalized prompt templates (geng2022recommendation, ; bao2023tallrec, ; zhang2023collm, ; zhao2023recommender, ).
+LLM-powered RSs generally conduct recommendations by integrating user historical behavior sequences and/or item and/or user features with personalized prompt templates (Geng et al., 2022; Bao et al., 2023; Zhang et al., 2023; Zhao et al., 2023).
 Let $\mathcal{X}_{u,v}$ represent the prompts of LLM-powered RSs (denoted as $\text{LLM}_{\Theta}$ with parameters $\Theta$), the preference function can be formulated as below.
-$\mathcal{R}_{u}=f_{\text{LLM}_{\Theta}}(\mathcal{X}_{u,v}),$ (1)
+(1) $\mathcal{R}_{u}=f_{\text{LLM}_{\Theta}}(\mathcal{X}_{u,v}),$
 where $\mathcal{X}_{u,v}=\mathcal{P}\oplus m_{u}\oplus\mathcal{M}_{v}^{u}\oplus\mathcal{C}_{v}^{u}$ and $\mathcal{R}_{u}$ is the recommendation results for $u$.
@@
 $m_{u}$ indicates user features and $\mathcal{M}_{v}^{u}=\{m_{v}:v\in\mathcal{V}_{u}\}$ represents the feature set of $u$’s interacted items.
-$\mathcal{C}_{v}^{u}$ denotes the feature set of candidate items that are not interacted by $u$, where the amount often depends on the type of task (xu2021understanding, ).
+$\mathcal{C}_{v}^{u}$ denotes the feature set of candidate items that are not interacted by $u$, where the amount often depends on the type of task (Xu et al., 2021).
 $|\mathcal{C}_{v}^{u}|>1$ if the task is from implicit feedback such as top-$\mathcal{K}$ recommendations, and $|\mathcal{C}_{v}^{u}|=1$ if it is formulated as an explicit feedback task such as rating predictions.
@@
 Note that all the elements of $\mathcal{X}_{u,v}$ are optional, which relies on its design.
-2.2 Threat Model
-2.2.1 Attacker’s Knowledge:
-It is often challenging to obtain the internal knowledge of real-world victim RSs, as noted by prior works (ning2024cheatagent, ; yang2024attacking, ; zhang2024stealthy, ).
-Specifically, critical details such as the parameters (e.g., agents’ memories, which are language-based embeddings (zhang2024agentcf, )) and architectures (e.g., the underlying LLM backbones) are typically inaccessible to adversaries.
+2.2. Threat Model
+2.2.1. Attacker’s Knowledge:
+It is often challenging to obtain the internal knowledge of real-world victim RSs, as noted by prior works (Ning et al., 2024; Yang et al., 2024; Zhang et al., 2024c).
+Specifically, critical details such as the parameters (e.g., agents’ memories, which are language-based embeddings (Zhang et al., 2024b)) and architectures (e.g., the underlying LLM backbones) are typically inaccessible to adversaries.
 Given these limitations, DrunkAgent is primarily designed and evaluated under black-box settings (in Section 4).
-Moreover, due to the inherently open nature of RSs (si2020shilling, ; 10415763, ), in both cases, the attacker is assumed to only have access to publicly available recommendation data such as item titles and user reviews.
-2.2.2 Attacker’s Objectives:
+Moreover, due to the inherently open nature of RSs (Si and Li, 2020; Yang et al., 2023), in both cases, the attacker is assumed to only have access to publicly available recommendation data such as item titles and user reviews.
+2.2.2. Attacker’s Objectives:
 As outlined in Section 1, the core objective of our attack is to maximize the exposure of a target item across the widest possible range of normal users.
-Since black-box attacks are typically optimized in a localized manner (e.g., via surrogate models (wang2024poisoning, )), achieving this objective requires strong transferability, i.e., the ability of the attack to remain effective across different black-box RSs (ning2024cheatagent, ).
-Furthermore, to broaden the attack impact and maximize the number of affected users in practice, stealthiness becomes a critical secondary objective, ensuring that the attack remains inconspicuous thus difficult to detect (zhang2024stealthy, ).
-2.2.3 Attacker’s Capabilities:
-In the black-box scenarios, the attacker has no access to the model’s internal states (e.g., user memories and candidate item memories) and can induce memory corruptions by modifying the description of the target item (i.e., the initial memory of the target item agent).
-Descriptions are often both lengthy and context-rich, which can easily conceal perturbations, as elaborated in Appendix A.
+Since black-box attacks are typically optimized in a localized manner (e.g., via surrogate models (Wang et al., 2024a)), achieving this objective requires strong transferability, i.e., the ability of the attack to remain effective across different black-box RSs (Ning et al., 2024).
+Furthermore, to broaden the attack impact and maximize the number of affected users in practice, stealthiness becomes a critical secondary objective, ensuring that the attack remains inconspicuous thus difficult to detect (Zhang et al., 2024c).
+2.2.3. Attacker’s Capabilities:
+In the black-box scenarios, the attacker has no access to the model’s internal states including user memories and candidate item memories, and can induce memory corruptions by modifying the description of the target item (e.g., the initial memory of the target item agent at $\mathcal{T}_{0}$ in Fig.
+1).
+Descriptions are often both lengthy and context-rich, which can easily conceal perturbations, as elaborated in Section 5.
 Moreover, the profit-driven merchant may frequently update the descriptions to overwrite the target memory that has been optimized over time or construct multiple descriptions of items indicating the same target item.
-This is feasible, due to the merchants are required to maintain their dedicated APIs (liu2021adversarial, ; yang2024attacking, ; zhang2024stealthy, ).
-Figure 2: DrunkAgent Overview.
+This is feasible, due to the merchants are required to maintain their dedicated APIs in recommendation platforms (Cohen et al., 2021; Liu and Larson, 2021; Zhang et al., 2024c).
+Figure 2.
+DrunkAgent Overview.
 The generation module produces adversarial textual triggers for promoting target items.
@@
 The triggers and the strategies are optimized on the surrogate module to improve the transferability and stealthiness of black-box attacks.
-3 DrunkAgent
+3. DrunkAgent
 In this section, we present the black-box attack framework, DrunkAgent. Fig.  2 gives an overview of DrunkAgent, which is composed of three modules. Each module is introduced in the following.
-3.1 Surrogate Module
-This module serves as a surrogate agentic RS (denoted as $f_{\text{Agent}_{\Phi}}(\cdot)$) to evaluate attack effectiveness under limited accessible resources, drawing inspiration from prior works lin2022shilling ; tang2020revisiting ; 10415763 ; yang2024attacking .
-To avoid frequent queries that could arouse suspicion and to adapt to realistic black-box conditions, the generation module and the strategy module are optimized within this surrogate setting to improve the stealthiness and transferability of DrunkAgent.
-As discussed in Section 1, agentic RSs are composed of user agents, item agents, and recommendation agents, and all built upon LLM backbones enhanced with memory modules, where the LLMs are used to perform text generation tasks.
-As such, open-source and auto-regressive LLMs are preferable for constructing surrogate backbones due to their strong generation capabilities and customizability.
-Given that the most influential Meta’s LLaMA family has been widely adopted in the recent recommendation community (zhao2023recommender, ; bao2023tallrec, ), the latest and powerful LLaMA 3 (dubey2024llama, ) is an ideal choice, which can provide more valuable and instructive feedback for effective optimizations, as illustrated in Fig.
-2.
-The memories for each types of surrogate agents will be detailed in the following sub-sections.
-3.2 Generation Module
+3.1. Surrogate Module
+To evaluate the attack effectiveness under limited accessible resources and to avoid frequent queries that may raise suspicion, we adopt a surrogate module to optimize the black-box attack, which is a commonly used approach for handling realistic black-box constraints (Lin et al., 2022; Tang et al., 2020; Yang et al., 2023, 2024).
+Within this surrogate setting, we jointly optimize both the generation and strategy modules to enhance the stealthiness and transferability of DrunkAgent.
+As illustrated in Fig.
+2, we construct the module with an agentic RS, denoted as $f_{\text{Agent}_{\Phi}}(\cdot)$, to simulate the behaviors of the real target system.
+The surrogate framework is composed of user agents, item agents, and recommendation agents, all built upon LLM backbones augmented with memory components.
+In addition, we incorporate task-aligned prompt templates as behavioral interfaces among the agents, enabling interaction and collaboration.
+As LLMs are required to perform reasoning and text generation tasks, open-source and auto-regressive LLMs are suitable for constructing the backbones due to their strong generative capabilities and high customizability.
+The design of memories and templates for different surrogate agents is detailed in the following sub-sections.
+3.2. Generation Module
 This module aims to generate effective adversarial description triggers for promoting the target items on the agentic RSs, consisting of several sub-modules, as shown in Fig.
 2 and described as follows.
-3.2.1 Adversarial Prompting Engineering
-This module designs the adversarial prompt for the surrogate recommendation agent to improve the quality of adversarial trigger generation.
-We craft prompt template, so that the surrogate knows ‘what needs to be done’ and ‘how to respond to’.
-As discussed in Section 2.2.1, it is challenging to obtain optimized memories of user agents and item agents in black-box attacks.
-To solve this and ensure the quality, we introduce general user memories and popular item memories.
+3.2.1. Recommendation Context Construction
+This module introduces adversarial prompts for the surrogate recommendation agent to improve the quality of adversarial trigger generation.
+Specifically, we craft a template to establish a recommendation-style interaction context.
+As discussed in Section 2.2.1, it is challenging to obtain optimized memories of user agents and item agents within the black-box attacks.
+Therefore, we propose to incorporate general user memories and popular item memories to approximate those memories to ensure both effectiveness and reliability.
 Basic Prompt Template.
-The template $\widetilde{\mathcal{P}}$ is built with four components: Style Definition $\widetilde{p_{s}}$, Task Goal $\widetilde{p_{g}}$, Recommendation Instruction $\widetilde{p_{i}}$ and Format Constraint $\widetilde{p_{f}}$, as shown in the example of Fig.
+The template $\widetilde{\mathcal{P}}$ is built with four components: 1) Style Definition $\widetilde{p_{s}}$, which introduces role information into the context to elicit domain-specific responses; 2) Task Goal $\widetilde{p_{g}}$, which specifies the ranking objective, aligning the task with recommendation (Zhang et al., 2019); 3) Recommendation Instruction $\widetilde{p_{i}}$, which provides explicit guidance for comparing user preferences with candidate items; and 4) Format Constraint $\widetilde{p_{f}}$, which restricts the output format to reduce undesirable outputs and facilitate subsequent optimization.
+An illustration is given in Fig.
 2.
-The detailed descriptions are provided in Appendix B.1.
 Formally, $\widetilde{\mathcal{P}}=\widetilde{p_{s}}\oplus\widetilde{p_{g}}\oplus\widetilde{p_{i}}\oplus\widetilde{p_{f}}$.
@@
 2.
-This is also the initial memory of the user agent (zhang2024agentcf, ; huang2024foundation, ), which is accessible and aligns with the assumptions of target black-box RSs.
+This is also the initial memory of the user agent (Zhang et al., 2024b; Huang et al., 2024), which is accessible and aligns with the assumptions of target black-box RSs.
 Popular Item Memories.
 Besides the limited access of the optimized memories in practice, the inherent opacity of the victim systems renders the exact set of candidate items used for recommendations unknown.
-As the works (si2020shilling, ; liu2021adversarial, ; wang2025id, ) pointed out, popular items usually have high probabilities of appearing at the top of users’ recommendation lists.
+As the works (Si and Li, 2020; Liu and Larson, 2021; Wang et al., 2025) pointed out, popular items usually have high probabilities of appearing at the top of users’ recommendation lists.
 Intuitively, if target items can be ranked ahead of these popular items, they are more likely to be promoted to normal users.
@@
 Raw item descriptions serve as a form of natural resource that can be leveraged to derive memories for popular items.
-However, such descriptions (ni2019justifying, ) are often too long that may exceed the context window limitations of the surrogate model and always contain too much redundant information that may increase the difficulty of subsequent optimizations.
+However, such descriptions (Ni et al., 2019) are often too long that may exceed the context window limitations of the surrogate model and always contain too much redundant information that may increase the difficulty of subsequent optimizations.
 To tackle the issues, we propose to adopt a LLM (denoted as $f_{\text{LLM}_{\Psi}}(\cdot)$) to extract valuable item features from the raw descriptions and supplement the features based on its internal rich knowledge base to make them more informative to facilitate the optimizations.
-Let $\mathcal{C}_{p}$ represent the modified feature sequence of popular items, $\mathcal{C}_{p}=f_{\text{LLM}_{\Psi}}(\mathcal{X}_{p}\oplus\mathcal{M}_{p})$, where $\mathcal{X}_{p}$ denotes the prompts for feature acquisition.
+Let $\mathcal{C}_{p}$ represent the modified feature sequence of popular items, $\mathcal{C}_{p}=f_{\text{LLM}_{\Psi}}(\mathcal{X}_{p}\oplus\mathcal{M}_{p})$, where $\mathcal{X}_{p}$ denotes the prompts for feature refinement.
 An example of producing a popular item memory is given in Fig.
 2.
-Note that if the raw item descriptions are empty, we adopt other identity information including titles and categories.
-Agents always tend to select candidates positioned higher in the display list (zhang2024agentcf, ).
+Agents always tend to select candidates positioned higher in the display list (Zhang et al., 2024b).
 To increase the trigger effectiveness, we place the other popular candidates before the target item, as shown in Fig.
@@
 As a result, the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ can be formulated as
-$\widetilde{\mathcal{X}_{t}}=\widetilde{\mathcal{P}}\oplus\widetilde{m_{u}}\oplus\mathcal{C}_{p}\oplus m_{t}.$ (2)
-3.2.2 Greedy Search Algorithm
-The algorithm is used to optimize the adversarial description triggers, inspired by the recent successes in jailbreak attacks on general-purpose LLMs (zou2023universal, ; liu2023autodan, ).
-Despite this inspiration, DrunkAgent fundamentally differs in key aspects such as attack goals and victim models.
-Additional details are given in Appendix A.
-To fill the gaps, the algorithm incorporates task-aligned initialization and refines the search space tailored to manipulate the behaviors of the target agents.
+(2) $\widetilde{\mathcal{X}_{t}}=\widetilde{\mathcal{P}}\oplus\widetilde{m_{u}}\oplus\mathcal{C}_{p}\oplus m_{t}.$
+3.2.2. Greedy Search Algorithm
+We introduce a greedy search algorithm to optimize the adversarial description triggers.
+Starting from attack goal–aligned seeds, the algorithm iteratively selects, recombines, and polishes candidates, progressively refining the search space to produce highly effective textual triggers for promoting target items.
 Adversarial Trigger Initialization.
@@
 Let $\widetilde{\mathcal{X}_{c}}$ denote a sequence of prompts,
-~$\widetilde{\mathcal{M}}_{c}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{c}}\oplus m_{t}),$ (3)
-where ~$\widetilde{\mathcal{M}}_{c}=\{\widetilde{m}_{t_{i}}\}$ represents a sequence of trigger candidates obtained from the LLM.
+(3) $\widetilde{\mathcal{M}}_{c}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{c}}\oplus m_{t}),$
+where $\widetilde{\mathcal{M}}_{c}=\{\widetilde{m}_{t_{i}}\}$ represents a sequence of trigger candidates.
 Adversarial Trigger Optimization.
@@
 We use the auto-regressive language generation loss (i.e., negative log-likelihood) to evaluate the discrepancy between the predictions and the target output $\widetilde{\mathcal{R}_{t}}$,
-$\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}=-\log\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)})),$ (4)
+(4) $\mathcal{L}_{\widetilde{m}_{t_{i}}^{(k-1)}}=-\log\mathbbm{P}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{R}_{t}}|\widetilde{\mathcal{X}_{t}}\oplus\widetilde{m}_{t_{i}}^{(k-1)})),$
 […]
 Feature Integration Stage.
-The algorithm samples a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates by conducting probability-based random selection.
+The algorithm samples a subset of $|\widetilde{\mathcal{M}}_{c}|-n$ input candidates by conducting probability-based random selection.
 The candidates with higher scores will have a greater probability of being selected, where the selection probability of each candidate is calculated via a softmax function, i.e., $\text{softmax}(\widetilde{m}_{t_{i}}^{(k-1)})=\frac{e^{s_{i}^{(k-1)}}}{\sum_{j=1}^{|\widetilde{\mathcal{M}}_{c}|}e^{s_{j}^{(k-1)}}}$.
-For the obtained subset, the algorithm splits the candidates at random positions into a list of text slices and performs the random exchange of the slices between pairwise texts, allowing features from different candidates to be fused, as shown in Fig.
-2.
+For the obtained subset, the algorithm splits the candidates at random positions (e.g., at punctuation marks) into a list of text slices and performs the random exchange of the slices between pairwise texts, as shown in Fig.
+2, allowing features from different candidates to be fused.
 Specifically, this fusion creates new combinations of words, phrases, and sentences, enhancing the feature diversity of candidates, and hence contributing to search for an optimal trigger.
+Algorithm 1 Optimization Procedure of DrunkAgent
+0: the user-item interaction matrix $\mathcal{Y}$ with the basic textual metadata of items and users (e.g., item titles and categories)
+0: the adversarial description of the target item $t$ Solve for the Optimal Adversarial Description Trigger on the Surrogate Module:
+1: Obtain the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ with Eq. (2)
 […]
+3:while $k\in\mathcal{E}$ epochs or the greedy algorithm does not converge (i.e. the attack goals are not achieved stably) do
+4:  Quality Estimation Stage:
+5:  Calculate $s_{i}^{(k-1)}$ for each input trigger $i$ with Eq. (4)
+6:  Maintain top-$n$ candidates in descending order of scores
+7:  Feature Integration Stage:
+8:  Sample a subset of $|\widetilde{\mathcal{M}}_{c}|-n$ input candidates via Softmax
+9:  Randomly exchange of the slices between pairwise texts
+10:  Linguistic Enrichment Stage:
+11:  Optimize and polish the $|\widetilde{\mathcal{M}}_{c}|-n$ previous combinations via LLMs
+12:  Obtain $|\widetilde{\mathcal{M}}_{c}|$ triggers for the next epoch optimization
+13:end while
+14: Obtain the optimal description trigger $\widetilde{m_{t}}^{*}$ with Eq. (5) Solve for the Optimal Adversarial Perturbation Strategy on the Surrogate Module:
+15: Define the adversarial strategy $\widetilde{q}$ with Eq. (6)
+16:while the desired malicious action $\widetilde{a}_{t}$ does not occur (i.e., the strategy cannot get the target agents drunk) do
+17:  Optimize the arrangement of the defined strategies in $\widetilde{q}$ with Eq. (7)
+18:end while
+19: Obtain the optimal strategy $\widetilde{q}^{*}$ when Eq. (7) is maximum
+20:return the optimal adversarial description $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$
+Linguistic Enrichment Stage.
+Given that interleaving two candidates can result in the combinations that are awkward or lack fluency, a LLM is introduced to polish the combinations.
+In this way, not only the original semantic meanings are preserved, but also the clarity, coherence and natural flow of the text are improved, thereby enhancing the concealment of the attack.
+Moreover, the LLM expands the candidates’ corpus to avoid the problem of language-based vanishing gradient causing the algorithm to fall into local optimum.
+Since excessive use of new data may lead to a language-based gradient explosion that makes the optimization unstable, the length limitation is introduced into the prompt $\widetilde{\mathcal{X}_{r}}$.
+Formally, for each new candidate $\widetilde{m}_{t_{i}}^{(k)}$, $\widetilde{m}_{t_{i}}^{(k)}=f_{\text{LLM}_{\Psi}}(\widetilde{\mathcal{X}_{r}}\oplus\widetilde{m}_{t_{i}}^{(k-1)^{\prime}})$, where $\widetilde{m}_{t_{i}}^{(k-1)^{\prime}}$ denotes the resulting combination from the previous stage.
+Consequently, $n$ well-performing trigger candidates from the first stage and the subset $\{\widetilde{m}_{t_{i}}^{(k)}:i\ \text{is the index of the second stage's candidate}\}$ with the length $|\widetilde{\mathcal{M}}_{c}|-n$ will be optimized in the next epoch $k+1$.
 The optimal trigger $\widetilde{m_{t}}^{*}$ will be obtained when the candidate performs the best,
-$\widetilde{m_{t}}^{*}=\mathop{\arg\max}_{\widetilde{m}_{t_{i}}^{\mathcal{E}}}(\{s_{i}^{\mathcal{E}}\}),$ (5)
+(5) $\widetilde{m_{t}}^{*}=\mathop{\arg\max}_{\widetilde{m}_{t_{i}}^{\mathcal{E}}}(\{s_{i}^{\mathcal{E}}\}),$
 where $\{s_{i}^{\mathcal{E}}\}$ is the sequence of final candidates’ scores. An example of $\widetilde{m_{t}}^{*}$ is given in Fig. 2.
-3.3 Strategy Module
+3.3. Strategy Module
 We introduce the strategy module to impede the target item agent from evolving from the environment, that is, the memory of the target agent cannot be effectively updated during agent-environment interactions.
 As such, the well-crafted adversarial description trigger $\widetilde{m_{t}}^{*}$ is retained, allowing the promotion of the target item to be maximized.
-3.3.1 Perturbation Strategy Definition.
 […]
-$\widetilde{q}=\widetilde{q_{f}}\oplus\widetilde{q_{c}}\oplus\widetilde{q_{g}}\oplus\widetilde{q_{n}}\oplus\widetilde{q_{s}},$ (6)
-where an example of $\widetilde{q}$ is given in Fig. 2.
-3.3.2 Adversarial Strategy Optimization.
-The arrangement of the defined strategies needs to be optimized to improve the attack effectiveness.
-DrunkAgent obtains the optimal strategy $\widetilde{q}^{*}$ by maximizing the expected probability that the agent when influenced by adversarial modifications, performs a malicious action for a given input query,
-$\mathbbm{E}_{\widetilde{q}\sim\pi_{\widetilde{q}}}[\ \mathbbm{I}(f_{\text{Agent}_{\Phi}}(\widetilde{\mathcal{X}_{s}}\oplus\widetilde{q}\oplus\widetilde{m_{t}}^{*})=\widetilde{a_{t}})],$ (7)
-where $\pi_{\widetilde{q}}$ denotes the distribution of adversarial strategies, $\mathbbm{I}(\cdot)$ is an indicator function and $\widetilde{\mathcal{X}_{s}}$ is an adversarial prompt for simulating memory updates of item agents during the agents’ environmental interactions.
-$\widetilde{a_{t}}$ is the desired malicious action for the injected strategy $\widetilde{q}$, i.e., the target item agent is unable to fulfill specific roles and is incapable of perceiving, learning and self-evolving from the interaction environment (e.g., fail in language-based signal back-propagation (zhang2024agentcf, )).
-An example of $\widetilde{q}^{*}$ is given in Fig.
-2.
-To sum up, the adversarial description of the target item is $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$.
-The overall optimization procedure of DrunkAgent will be provided in Algorithm 1.
-The above design further reveals how memory mechanisms in agent-based recommender systems can be influenced under black-box conditions.
-These insights may support future efforts toward developing memory-aware defenses (e.g., using customized deep neural network detectors (wu2020densely, ; yang2021dualnet, ; yang2021hunter, ; yang2021deep, )) and enhancing the robustness of agent-powered RSs chua2024ai .
-4 Experiments
-4.1 Experimental Setup
-Datasets Selection.
-To comprehensively evaluate DrunkAgent, we use three datasets that are text-intensive containing real-world data and widely-adopted in the current recommendation studies (huang2024foundation, ; zhang2024agentcf, ).
-They are CDs & Vinyl, Office Products and Musical Instruments of Amazon Review Data (ni2019justifying, ).
-Their statistics are given in Appendix C.1.1.
-We additionally evaluate our method on Yelp dataset [^2] https://www.yelp.com/dataset to demonstrate its generalizability across domains in the appendix.
-The target items are randomly sampled from the item pool on each dataset.
-The universality of attacks across different target items is further demonstrated in Appendix C.2.1.
-Baseline Attack Methods.
-Since we are the first work to gain insight into the robustness of the generative agent-powered RSs, there is no any baselines.
-To comprehensively evaluate the effectiveness of DrunkAgent, we follow the state-of-the-art attack on LLM-based RSs (zhang2024stealthy, ) and compare DrunkAgent with the benign status and six black-box attack methods, where the perturbations range from character-level: DeepwordBug (gao2018black, ) and PuncAttack (formento2023using, ), word-level: TextFooler (jin2020bert, ) and BertAttack (li2020bert, ) to sentence-level: TrivialInsertion and ChatGPTAttack (zhang2024stealthy, ).
-To maintain a fair comparison, the methods perform perturbations on the descriptions of the target items and are optimized by the surrogate model that is the same as DrunkAgent.
-More detailed implementations are provided in Appendix C.1.2.
-Targeted Recommender Systems.
-We adopt three different agentic paradigms (zhang2024agentcf, ) as victim RSs to evaluate the effectiveness of DrunkAgent across diverse agent designs and recommendation tasks: AgentCF, a standard CF paradigm; AgentRAG, which augments the agents with retrieval; and AgentSEQ, which captures temporal dynamics in user behaviors for sequential recommendations.
-Given that the victim RSs are totally different from as well as more powerful and context-rich than the surrogate, a strict black-box setting is maintained for sufficient evaluations.
-To ensure a fair evaluation, we follow the default implementations, as elaborated in Appendix C.1.3.
-The configuration of DrunkAgent is given in Appendix C.1.4.
-We evaluate its performance using widely-used HR@$\mathcal{K}\uparrow$, NDCG@$\mathcal{K}\uparrow$, and sentence perplexity score $\downarrow$.
-The specific definitions of the metrics are in Appendix C.1.5.
-More experiments and demonstrations (e.g., ablation studies and the sensitivity of critical parameters) can be found in the appendix.
-Table 1: Attack Transferability.
-HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ of different attacks against various black-box victim LLM-powered agent-based RSs on real-world datasets.
-We use bold fonts to denote the best performance.
-The attacks with higher HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ have excellent transferability.
 Victim RS AgentCF
@@
 DrunkAgent 0.4949 0.4949 0.5152 0.4949 0.4949 0.5051 0.1429 0.1429 0.1531 0.1429 0.1429 0.1480 0.1443 0.1443 0.1546 0.1443 0.1443 0.1495
-4.2 Attack Transferability
-Overall Transferability.
+3.3.1. Perturbation Strategy Definition.
+[…]
+(6) $\widetilde{q}=\widetilde{q_{f}}\oplus\widetilde{q_{c}}\oplus\widetilde{q_{g}}\oplus\widetilde{q_{n}}\oplus\widetilde{q_{s}}$,
+where an example of $\widetilde{q}$ is given in Fig. 2.
+3.3.2. Adversarial Strategy Optimization.
+The organization of the defined strategies needs to be optimized to improve the effectiveness of ‘get the target agent drunk’.
+DrunkAgent obtains the optimal strategy $\widetilde{q}^{*}$ by maximizing the expected probability that the agent when influenced by adversarial modifications, performs a malicious action for a given input query,
+[…]
+where $\pi_{\widetilde{q}}$ denotes the distribution of adversarial strategies, $\mathbbm{I}(\cdot)$ is an indicator function and $\widetilde{\mathcal{X}_{s}}$ is an adversarial prompt template for simulating memory updates of item agents during the agents’ environmental interactions.
+$\widetilde{a_{t}}$ is the desired malicious action for the injected strategy $\widetilde{q}$, i.e., the target item agent is unable to fulfill specific roles and is incapable of perceiving, learning and self-evolving from the interaction environment (e.g., fail in language-based signal back-propagation (Zhang et al., 2024b)).
+An example of $\widetilde{q}^{*}$ is given in Fig.
+2.
+To sum up, the adversarial description of the target item is $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$.
+For transparency and reproducibility, the overall optimization procedure of DrunkAgent is given in Algorithm 1.
+The above design further reveals how memory mechanisms in agent-based recommender systems can be influenced under black-box conditions.
+These insights may support future efforts toward developing memory-aware defenses (e.g., using customized deep neural network detectors (Wu et al., 2020; Yang et al., 2021b, a; Yang, 2021)) and enhancing the robustness of agent-powered RSs (Chua et al., 2024).
+4. Experiments
+4.1. Experimental Setup
+4.1.1. Datasets Selection.
+To comprehensively evaluate DrunkAgent, we use three datasets that are text-intensive containing real-world data and widely-adopted in the current recommendation studies (Huang et al., 2024; Zhang et al., 2024b).
+They are CDs & Vinyl, Office Products and Musical Instruments of Amazon Review Data (Ni et al., 2019).
+We additionally evaluate our method on Yelp dataset [^2] https://www.yelp.com/dataset to demonstrate its generalizability across domains in the appendix.
+The target items are randomly sampled from the item pool on each dataset.
+4.1.2. Baseline Attack Methods.
+Since we are the first work to gain insight into the robustness of the generative agent-powered RSs, there is no any baselines.
+To comprehensively evaluate the effectiveness of DrunkAgent, we follow the state-of-the-art attack on LLM-based RSs (Zhang et al., 2024c) and compare DrunkAgent with the benign status and six black-box attack methods, where the perturbations range from character-level: DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023), word-level: TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020) to sentence-level: TrivialInsertion and ChatGPTAttack (Zhang et al., 2024c).
+To maintain a fair comparison, the methods perform perturbations on the descriptions of the target items and are optimized by the surrogate model that is the same as DrunkAgent.
+4.1.3. Targeted Recommender Systems.
+We adopt three different agentic paradigms (Zhang et al., 2024b) as victim RSs to evaluate the effectiveness of DrunkAgent across diverse agent designs and recommendation tasks: AgentCF, a standard CF paradigm; AgentRAG, which augments the agents with retrieval; and AgentSEQ, which captures temporal dynamics in user behaviors for sequential recommendations.
+Given that the victim RSs are totally different from as well as more powerful and context-rich than the surrogate, a strict black-box setting is maintained for sufficient evaluations.
+To ensure a fair evaluation, we follow the default implementations.
+4.1.4. Evaluation Metrics
+[…]
+4.1.5. Configuration of DrunkAgent
+For the surrogate $f_{\text{Agent}_{\Phi}}(\cdot)$ backbone, we adopt the open-source model Meta-Llama-3-8B-Instruct [^3] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct, a recent and well-established LLM, which provides a favorable trade-off between efficiency and quality.
+In contrast, OpenAI’s gpt-4-turbo [^4] https://platform.openai.com/docs/models/gpt-4-turbo is adopted to implement $f_{\text{LLM}_{\Psi}}(\cdot)$, offering stronger effectiveness while maintaining reasonable cost.
+We set $\mathcal{E}=20$, $|\widetilde{\mathcal{M}}_{c}|=10$ and $n=5$.
+All random seeds in the evaluation are set to 2024 and the victim LLMs’ temperature is set to 0 for reproducibility.
+For all the other parameters, we keep the default values.
+More implementation details, experiments and demonstrations (e.g., ablation studies and the sensitivity of critical parameters) can be found in the appendix.
+4.2. Attack Transferability
+4.2.1. Overall Transferability.
 Table 1 shows the transferability of the attacks under practical black-box settings.
 The results indicate that DrunkAgent is effective and transferable regardless of the victim model’s architectures, parameters and prompts as well as datasets, and the existing LLM-powered agentic RSs are sensitive to memory perturbations.
-This is mainly because the adversarial text effectively reflects the attack goals.
+This is mainly because the three modules collaborate effectively, producing adversarial texts that accurately reflect the attack goals.
 It is worth mentioning that DrunkAgent always pushes the target items to the first of the recommendation lists.
@@
 In addition, AgentRAG is more robust than AgentCF and AgentSEQ, due to its advanced retrieval mechanisms.
-DrunkAgent vs.
-Baseline Attacks.
+4.2.2. DrunkAgent vs. Baseline Attacks.
 From Table 1, DrunkAgent greatly outperforms the state-of-the-art attack baselines against different victim models across various real-world datasets by achieving higher HRs and NDCGs, which shows its excellent transferability.
@@
 This indirectly demonstrates the importance of our strategy module and the transferability of our triggers.
-4.3 Attack Stealthiness
-Overall Stealthiness.
+Figure 3. Attack Universality across Target Items.
+4.3. Attack Universality
+In addition to cross-model transferability, we evaluate the cross-sample transferability of DrunkAgent.
+We incorporate the generated adversarial descriptions into totally different target items that are randomly sampled from each dataset.
+HR@1 (i.e., NDCG@1) of DrunkAgent across different models and target items on real-world datasets are shown in Fig.
+3, which indicates its commendable transferability and triggers’ universality across different target items.
+This may be attributed to the fact that our triggers encapsulate diverse and general item characteristics, and, crucially, they also reveal the intended attack objectives.
+Figure 4. Attack Stealthiness. The overall distribution of recommendation performance differences of all the victim agentic models on all real-world datasets before and after the attacks.
+Figure 5. Attack Imperceptibility. Perturbed text’s perplexity on real datasets.
+4.4. Attack Stealthiness
+4.4.1. Overall Stealthiness.
 Fig.
@@
 From the figure, we can find that DrunkAgent does not induce drastic changes in the overall performance, which signifies the attack does not disrupt the normal operation of RSs, making it difficult for users and platforms to detect and indicating its remarkably stealthy and unnoticeable.
-Figure 3: Attack Stealthiness. The overall distribution of recommendation performance differences of all the victim agentic models on all real-world datasets before and after the attacks.
-Attack Imperceptibility.
-Following the works (zhang2024stealthy, ; liu2024formalizing, ), we further evaluate the attack stealthiness by assessing its imperceptibility, where GPT-Neo’s text perplexity is used as the metric.
-A lower perplexity indicates that the perturbed text remains close to natural language, thus appearing less suspicious.
+4.4.2. Attack Imperceptibility.
+Following the works (Zhang et al., 2024c; Liu et al., 2024a), we further evaluate the attack stealthiness by assessing its imperceptibility, where GPT-Neo’s text perplexity is used as the metric.
+A lower perplexity indicates that the perturbed text remains close to natural language and is more fluent, making it harder to detect as manipulated or adversarial.
 From Fig.
-4, we can find that DrunkAgent is more imperceptible than other attacks by achieving a lower perplexity.
+5, we can find that DrunkAgent is more imperceptible than other attacks by achieving a lower perplexity.
 This also shows that the adversarial descriptions generated from DrunkAgent are high-quality, which is not only coherent and fluent but also semantically meaningful and stealthy.
-Figure 4: Attack Imperceptibility. Perturbed text’s perplexity on real datasets.
-4.4 Defense Strategies to Attacks
-Following the representative works (zhang2024stealthy, ; zhang2024agent, ), we use the paraphrasing defensive strategy (Para.) via OpenAI-GPT-o1 to combat the attacks of agentic models.
+4.5. Defense Strategies to Attacks
+Following the representative works (Zhang et al., 2024c, 2025), we use the paraphrasing defensive strategy (Para.) via OpenAI-GPT-o1 to combat the attacks of agentic models.
 DrunkAgent is still transferable, which shows that DrunkAgent is robust to such defenses, as shown in Fig.
-5.
+6.
 Moreover, from Fig.
@@
 It could be that the substitution of certain new words/phrases/sentences improves the semantically meaningfulness of adversarial text, making our attacks stronger.
-5 Conclusion
 (a) CDs & Vinyl
@@
 (c) Musical Instruments
-Figure 5: DrunkAgent’s Robustness to Defense Mechanisms.
+Figure 6. DrunkAgent’s Robustness to Defense Mechanisms
+5. Related Works
+5.1. Data Poisoning Attacks on Traditional Recommender Systems
+In recent years, data poisoning attacks (aka shilling attacks) (Li et al., 2016; Si and Li, 2020; Lin et al., 2020; Huang et al., 2021; Lin et al., 2022; Yang et al., 2023; Zeng et al., 2023; Wang et al., 2024a) have been fully investigated to perform robustness analysis on traditional RSs (e.g., NCF (He et al., 2017), LightGCN (He et al., 2020)).
+Such attacks interfere the training process of models by injecting fake user profiles, where the profiles typically are a set of well-crafted numerical ratings on items.
+However, they are less effective (Zhang et al., 2024c) even of limited applicable to the recent RSs that are empowered by LLMs (Ning et al., 2024).
+The main reasons are: 1) LLM-powered RSs either leverage LLM’s in-context learning capabilities or fine-tune the LLMs with very little data to improve recommendations (Zhao et al., 2023; Huang et al., 2025a), which is not required to retrain the models; 2) there are textual semantic gaps between such attacks and current models.
+Specifically, the attacks have difficulty in processing textual inputs (e.g., item titles and descriptions) and lack context awareness, making it challenging to effectively target LLM-empowered RSs that rely primarily on text inputs to generate natural language responses (Ning et al., 2024).
+Hence, there is an urgent need to investigate a novel text attack paradigm mainly tailored for the inference phase of language models (Zhang et al., 2024c).
+5.2. Adversarial Attacks on LLM-based Recommender Systems
+To fill the above gaps, two LLM-based RS attacks have been proposed (Zhang et al., 2024c; Ning et al., 2024).
+Although their attacks are conducted under a black-box setting, they are still significantly deviate from practical, which overestimates the robustness of LLM-powered RSs under realistic conditions.
+Zhang et al.
+(Zhang et al., 2024c) propose a novel black-box attack by leveraging the rewriting capabilities of ChatGPT, where attackers can significantly boost a target item’s exposure by merely altering its textual content during the testing.
+This attack overcomes this limitation that classical text attacks (Gao et al., 2018; Formento et al., 2023; Jin et al., 2020; Li et al., 2020) fail on LLM-based RSs due to misaligned malicious objectives.
+Compared with those traditional shilling attacks, such an item representation attack is notably stealthy (Zhang et al., 2024c), as it does not affect the overall recommendation performance, and it should be more efficient and lower in economic cost, due to it perturbs the item features directly and is not require to hire online writers to introduce additional fake user reviews.
+However, its transferability is limited, due to the attack is not optimized on any black-box RSs (Huang et al., 2021; Yang et al., 2024).
+Moreover, since the attack appends positive words into the target item titles, the titles’ brevity and frequent exposure to users tend to make even small perturbations noticeable.
+Meanwhile, another black-box attack CheatAgent (Ning et al., 2024) has been proposed.
+This work assumes the prompt templates used by victim models can be arbitrarily perturbed, which is difficult to realize in practice, as such perturbations often require insider access or even system compromises.
+Such attacks aim to undermine the overall recommendation performance, making the attack perceptible (Zhang et al., 2024c).
+As such, CheatAgent is an untargeted attack, which is beyond the scope of this paper, since we focus on targeted attacks.
+Unlike static traditional or vanilla LLM-based RSs, dynamic agentic RSs (Huang et al., 2024, 2025a, 2025b) maintain internal memory states that enable continuous adaptation to evolving user behaviors and changing contexts.
+This adaptivity yields more personalized and context-aware recommendations, but it also introduces new vulnerabilities that static models do not exhibit (e.g., memory corruption), calling for robustness analyses tailored to LLM-powered recommender agents.
+To tackle the above problems, in this paper, we provide the first work towards systematically investigate the security vulnerabilities in the agent-powered RSs.
+As such, we propose a practical black-box attack framework, DrunkAgent, for the LLM-based agentic RSs with the aim of promoting target items.
+We perform stealthy memory corruptions by crafting effective and imperceptible textual descriptions of target items, due to the greater length and rich contextual information of descriptions can easily conceal adversarial perturbations compared to the titles.
+While inspired by static jailbreak attacks on general-purpose LLMs (Zou et al., 2023; Liu et al., 2024b, a), DrunkAgent establishes a new threat model and develops tailored optimization techniques for adapting to dynamic victim architectures and distinct attack goals in agentic recommendation scenarios.
+6. Conclusion
 In this paper, we propose DrunkAgent, a novel attack framework targeting LLM-powered agentic RSs, that promotes target items by effectively perturbing the memories of the targe agents.
@@
 As generative agent-based paradigms thrive in the recommendation community, our findings pave the way for shaping the next generation of secure and resilient RSs and defensive models.
-Appendix A Related Works
-In recent years, data poisoning attacks (aka shilling attacks) [39, 18, 40, 41, 21, 19, 42, 12] have been fully investigated to perform robustness analysis on traditional RSs (e.g., matrix factorization [43], NCF [44], LightGCN [45]).
-Such attacks interfere the training process of models by injecting fake user profiles, where the profiles typically are a set of well-crafted numerical ratings on items.
-However, they are less effective [13] even of limited applicable to the recent RSs that are empowered by LLMs [14].
-The main reasons are: 1) LLM-powered RSs either leverage LLM’s in-context learning capabilities or fine-tune the LLMs with very little data to improve recommendations [15, 11], which is not required to retrain the models; 2) there are textual semantic gaps between such attacks and current models.
-Specifically, the attacks have difficulty in processing textual inputs (e.g., item titles and descriptions) and lack context awareness, making it challenging to effectively target LLM-empowered RSs that rely primarily on text inputs to generate natural language responses [14].
-Hence, there is an urgent need to investigate a novel text attack paradigm mainly tailored for the inference phase of language models [13].
-Meanwhile, jailbreak attacks [26, 27, 28] are proposed to undermine the general-purpose LLMs’ human-aligned values or other constraints imposed by model developers, compelling them to respond to malicious questions posed by adversaries with correct answers rather than rejection.
-Nevertheless, such attacks have different attack goals from those item promotion attacks for RSs, making them hard to effectively transfer to the LLM-based RSs.
-To fill the above gaps, two LLM-based RS attacks have been proposed [13, 14].
-Although their attacks are conducted under a black-box setting, they are still significantly deviate from practical, which overestimates the robustness of LLM-powered RSs under realistic conditions.
-Zhang et al.
-[13] propose a novel black-box attack by leveraging the rewriting capabilities of ChatGPT, where attackers can significantly boost a target item’s exposure by merely altering its textual content during the testing.
-This attack overcomes this limitation that classical text attacks [34, 35, 36, 37] fail on LLM-based RSs due to misaligned malicious objectives.
-Compared with those traditional shilling attacks, such an item representation attack is notably stealthy [13], as it does not affect the overall recommendation performance, and it should be more efficient and lower in economic cost, due to it perturbs the item features directly and is not require to hire online writers to introduce additional fake user reviews.
-However, its transferability is limited, due to the attack is not optimized on any black-box RSs [41, 17].
-Moreover, since the attack appends positive words into the target item titles, the titles’ brevity and frequent exposure to users tend to make even small perturbations noticeable.
-Another black-box attack CheatAgent [14] has been proposed.
-This work claims the prompt templates and user profiles on the victim models can be arbitrarily perturbed.
-However, such assumptions are impractical.
-The main reasons are: 1) tampering with prompt templates may induce the overall performance changes, making the attack perceptible [13]; and 2) it is difficult to inject interaction records into the specific positions of user profiles, because historical interactions are usually introduced into prompts in chronological order [5].
-Furthermore, CheatAgent is an untargeted attack, which is beyond the scope of this paper, due to we focus on targeted attacks.
-Traditional RSs [46] are predominantly based on static models, where user preferences are learned from historical data in an offline manner and remain fixed during inference.
-These models do not adapt to evolving user behaviors or changing contexts after training.
-Vanilla LLM-based recommenders [15] follow a similar static paradigm: they generate recommendations through single-turn, stateless inference, relying on pre-defined prompt inputs to represent user preferences.
-In contrast, dynamic LLM-powered agentic recommenders [7, 11, 8] maintain internal states (e.g., memory mechanisms), allowing them to continuously adapt to and self-evolve over time based on sequential user interactions and simulated environmental feedback, even in offline experimental settings.
-This adaptivity enables personalized, context-aware, and temporally-sensitive recommendations.
-However, such dynamic mechanisms also introduce novel vulnerabilities that static models do not exhibit (e.g., memory corruption), which call for new robustness analyses tailored to the agent recommenders.
-While there are some attack methods on LLM-based agents [38], they often pursue goals that diverge from those relevant to RSs, leaving a gap in effective adversarial attacks for agent-driven recommenders.
-To tackle the above problems, in this paper, we provide the first work towards systematically investigate the security vulnerabilities in the agent-powered RSs.
-As such, we propose a practical black-box attack framework, DrunkAgent, for the LLM-powered agentic RSs with the aim of promoting target items.
-We perform stealthy memory corruptions by crafting effective and imperceptible textual descriptions of target items, due to the greater length and rich contextual information of descriptions can easily conceal adversarial perturbations compared to the titles.
-The adversarial descriptions are optimized via a surrogate model to enhance the transferability and stealthiness of DrunkAgent.
-Algorithm 1 Optimization Procedure of DrunkAgent
-0:  the user-item interaction matrix $\mathcal{Y}$ with the basic textual metadata of items and users (e.g., item titles and categories)
-0:  the adversarial description of the target item $t$ Solve for the Optimal Adversarial Description Trigger on the Surrogate Module:
-1:  Obtain the adversarial prompt $\widetilde{\mathcal{X}_{t}}$ with Eq. (2)
-[…]
-3:  while $k\in\mathcal{E}$ epochs or the greedy algorithm does not converge (i.e. the attack goals are not achieved stably) do
-4:     Quality Estimation Stage:
-5:     Calculate $s_{i}^{(k-1)}$ for each input trigger $i$ with Eq. (4)
-6:     Maintain top-$n$ candidates in descending order of scores
-7:     Feature Integration Stage:
-8:     Sample a subset of |~$|\widetilde{\mathcal{M}}_{c}|-n$ input candidates via Softmax
-9:     Randomly exchange of the slices between pairwise texts
-10:     Linguistic Enrichment Stage:
-11:     Optimize and polish the |~$|\widetilde{\mathcal{M}}_{c}|-n$ previous combinations via LLMs
-12:     Obtain |~$|\widetilde{\mathcal{M}}_{c}|$ triggers for the next epoch optimization
-13:  end while
-14:  Obtain the optimal description trigger $\widetilde{m_{t}}^{*}$ with Eq. (5) Solve for the Optimal Adversarial Perturbation Strategy on the Surrogate Module:
-15:  Define the adversarial strategy $\widetilde{q}$ with Eq. (6)
-16:  while the desired malicious action $\widetilde{a}_{t}$ does not occur (i.e., the strategy cannot get the target agents drunk) do
-17:     Optimize the arrangement of the defined strategies in $\widetilde{q}$ with Eq. (7)
-18:  end while
-19:  Obtain the optimal strategy $\widetilde{q}^{*}$ when Eq. (7) is maximum
-20:  return  the optimal adversarial description $\widetilde{m_{t}}^{*}\oplus\widetilde{q}^{*}$
-Appendix B Methodology
-B.1 Supplements to Basic Adversarial Prompt Template
-The main components of the prompt template are elaborately described in the following.
-  1. 1.
-    Style Definition $\widetilde{p_{s}}$.
-There are two common styles of prompts: task-oriented and role-oriented [47].
-To make recommendations more authentic and reasonable, we introduce role information into the context, which allows the system to respond with domain-specific knowledge, such as ‘You are a CD enthusiast.’, as shown in Fig.
-2.
-  2. 2.
-    Task Goal $\widetilde{p_{g}}$.
-Since a recommender system is essentially a ranker [46], we set sorting goals to indicate recommendation tasks, such as ‘Now, you are considering to sort ten candidate CDs in order of how well they align with your preferences.’ in the given example of the figure.
-  3. 3.
-    Recommendation Instruction $\widetilde{p_{i}}$.
-We introduce highly detailed instructions to guide the model to perform accurate recommendations.
-For example, ‘Please sort CDs by considering the correlation between your preferences/dislikes and the features of the CDs.
-The higher the CD rank, the more you like it.’ in the shown example.
-  4. 4.
-[…]
-B.2 DrunkAgent Algorithm
-For reproducibility, the overall optimization procedure of DrunkAgent is given in Algorithm 1.
-For the sake of algorithmic transparency, our code will be made available.
-Importantly, to mitigate potential misuse risks, the released code excludes scripts for automatic attack execution and is provided solely for academic research under a non-commercial license.
-All the users are required to agree to responsible use guidelines before accessing the whole repository.
-Appendix C Experiments
-C.1 Supplements to Experimental Setup
-Table 2: Statistics of Datasets
+Appendix A Appendix
+A.1. Supplements to Experimental Setup
+Table 2. Statistics of Datasets
 Dataset #Users #Items #Interactions Sparsity
@@
 Musical Instruments 27,530 10,620 231,392 99.92%
-Table 3: Examples of baselines’ adversarial descriptions of target items on real-world datasets. The red part points out the differences from the original text.
+Table 3. Examples of baselines’ adversarial descriptions of target items on real-world datasets. The red part points out the differences from the original text.
 Attack Method The Adversarial Description of The Target Item $\mathcal{Q}$
@@
 sound that will elevate your music collection.
-C.1.1 Dataset Statistics
+A.1.1. Dataset Statistics
 Table 2 illustrates the statistics of these datasets.
 The datasets vary in size and sparsity, which is suitable for a comprehensive evaluation.
-Following the work [5], we further sample subsets from the datasets given the expensive API calls.
+Following the work (Zhang et al., 2024b), we further sample subsets from the datasets given the expensive API calls.
 Specifically, considering the importance and commonness of data sparsity and cold-start in the recommendation community, we randomly sample 100 users with 800 interactions including 777 items, 99 users with 693 interactions including 619 items and 98 users with 588 interactions including 483 items from the three datasets, respectively, allowing to explore more diverse and practical interaction scenarios.
-Moreover, the leave-one-out strategy [44] is used for evaluation, where the split between the training and test set is 9:1.
+Moreover, the leave-one-out strategy (He et al., 2017) is used for evaluation, where the split between the training and test set is 9:1.
 Scaling DrunkAgent for larger datasets is left as future work.
-C.1.2 Baseline Methods
+A.1.2. Baseline Methods
 The baselines perturb the descriptions of the target items at various levels, which are described as below.
-  • Character-level Perturbations. DeepwordBug [34] and PuncAttack [35] manipulate texts by introducing typos and inserting punctuation, respectively.
-  • Word-level Perturbations. TextFooler [36] and BertAttack [37], respectively, aim to replace words with synonyms or contextually similar words.
+  • Character-level Perturbations. DeepwordBug (Gao et al., 2018) and PuncAttack (Formento et al., 2023) manipulate texts by introducing typos and inserting punctuation, respectively.
+  • Word-level Perturbations. TextFooler (Jin et al., 2020) and BertAttack (Li et al., 2020), respectively, aim to replace words with synonyms or contextually similar words.
   • Sentence-level Perturbations.
-TrivialInsertion [13] inserts text content with $k$ number of positive words from a pre-defined word corpus, where the insertion is conducted at the end of the text.
-Here, $k=2$.
-ChatGPTAttack [13] increases the stealthiness of TrivialInsertion by rewriting the text content via GPTs, where GPT-3.5-turbo is used as the original paper recommends and the maximum number of words is set to 30 here.
+TrivialInsertion (Zhang et al., 2024c) inserts text content with several positive words from a pre-defined word corpus, where the insertion is conducted at the end of the text.
+ChatGPTAttack (Zhang et al., 2024c) increases the stealthiness of TrivialInsertion by rewriting the text content via GPTs.
 We give some examples of adversarial descriptions of target items generated by the baselines in Table 3.
-Following the works [5, 7], the item memory of the agent can be initialized by their identity information, such as titles and categories, as shown in the benign (default) descriptions of the tables.
+Following the works (Zhang et al., 2024b; Huang et al., 2024), the item memory of the agent can be initialized by their identity information, such as titles and categories, as shown in the benign (default) descriptions of the tables.
 Although the adversarial description of the target item generated from DrunkAgent is different from the benign status’s (see Fig.
-2 and Table 3), such modifications are feasible due to the real-world item providers are allowed to modify their product representations via APIs [17, 48, 20], as discussed in Section 1, especially when the benign description is just a default exemplar used for experiments.
+2 and Table 3), such modifications are feasible due to the real-world item providers are allowed to modify their product representations via APIs (Yang et al., 2024; Cohen et al., 2021; Liu and Larson, 2021), as discussed in Section 1, especially when the benign description is just a default exemplar used for experiments.
 Moreover, compared with the number of queries (denoted as $\mathcal{Q}$ in the table) of the baselines, DrunkAgent has the lowest attack costs due to $\mathcal{E}$=20.
-The implementation and parameters of the attack baselines follow the open-source works: StealthyAttack [^3] https://github.com/CRIPAC-DIG/RecTextAttack [13] and PromptBench [^4] https://github.com/microsoft/promptbench [49].
-C.1.3 Victim LLM-powered Agentic Recommender Systems
-LLM-driven agents, known for their superior autonomous interaction and decision-making capabilities, are gradually being considered as next-generation RSs [11].
-Compared with some agent-powered RSs that focus solely on user-side behavior modeling using universal LLMs [50, 7], AgentCF [5] emphasizes the modeling of two-sided interaction relations between user agents and item agents through the idea of collaborative filtering [44].
-Both kinds of agents are equipped with memory modules, maintaining the simulated preferences and tastes of potential adopters involving their intrinsic features and acquired behavioral information via autonomous interactions and collaborative optimizations at each time step, as shown on the left of Fig.
-1.
-As such, the memories of the user agents and item agents can be accordingly updated, enabling the agents to better fit the real-world interaction behaviors.
-To better adapt to the dynamically changing real-world environment, the agentic RSs are equipped with real-time capabilities.
-Whenever new descriptions are introduced at a time step, they are integrated into the evolving memories to maintain up-to-date contextual understanding.
-As can be seen on the right of Fig.
-1, the recommendation agents conduct personalized recommendations based on the optimized memories.
-To drive varied architectural designs and support diverse tasks for a comprehensive evaluation of the transferability of attacks, powerful mechanisms (e.g., retrieval augmentation) and task-specific prompts (e.g., sequential recommendation) are introduced to strengthen the agentic RSs.
-As such, AgentRAG and AgentSEQ are developed [5].
-To maintain a fair evaluation, the implementation details of how the agents can accomplish autonomous interactions and collaborative optimizations are followed the original paper suggest [5], where gpt-3.5-turbo [^5] https://platform.openai.com/docs/overview is recommended to achieve the backbone LLMs.
-Specifically, such backbone delivers strong natural language understanding and generation capabilities that are sufficient for most recommendation-related tasks.
-In addition, it is cost-efficient, making it well-suited for multi-turn interactions where cost can quickly accumulate.
-Moreover, it has lower latency, supporting real-time responsiveness in interactive agent scenarios.
-To ensure the high-quality of recommendations, we enhance the prompting strategies for recommendation agents.
-Specifically, we introduce the Chain-of-Thought strategy [51] to carefully craft the prompt template $\mathcal{P}$.
-Given the context window constraints pointed out in [5], a target item and nine non-interactive items of each normal user are randomly selected for the ranking task of the RSs, which is also suggested by the original work.
-The design of recommendation agents is described below.
-  • AgentCF.
-Since both user and item agents are collaboratively optimized to model their two-sided relations, the memories of the user agents and item agents that are optimized by agent-environment interactions can be introduced into $\mathcal{P}$ directly.
-For each normal user $u$,
-    $\mathcal{R}_{u}=f_{\text{AgentCF}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$ (8)
-    where $m_{u}$ is the short-memory of user agent $u$, $m_{t}$ denotes the features of the target item agent including metadata and the memories, $\mathcal{C}_{t}^{u}$ is the feature set of the candidate item agents and $\mathcal{R}_{u}$ is the ranking result.
-  • AgentRAG.
-Although short-term memory describes the current preference of a user agent, retrieving their specialized preferences from long-term memory toward candidates can allow them to make more personalized inferences.
-For each normal user $u$,
-    $\mathcal{R}_{u}=f_{\text{AgentRAG}}(\mathcal{P}\oplus m_{u}\oplus m_{r}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$ (9)
-    where $m_{r}$ is retrieved specialized preference from the long-term memory of user agent $u$ by taking the memories of candidate items as queries.
-To ensure the quality while maintaining the efficiency, we use all-MiniLM-L6-v2666https://www.sbert.net/docs/sentence_transformer/pretrained_models.html as the sentence transformer for the similarity calculation.
-  • AgentSEQ.
-When interaction records are sparse and preference propagation is limited, we can further incorporate user historical interactions into prompts, enabling LLMs to serve as sequential recommenders.
-For each normal user $u$,
-    $\mathcal{R}_{u}=f_{\text{AgentSEQ}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{M}_{t}^{u}\oplus\mathcal{C}_{t}^{u}),$ (10)
-    where $\mathcal{M}_{t}^{u}$ are the corresponding item features of the historical interactions with user $u$.
-We give some examples of the prompting templates of three victim models in Table 4 for reproductions.
-Table 4: Examples of prompting templates of the LLM-based agentic victim models.
+The implementation and parameters of the attack baselines follow the open-source works: StealthyAttack [^5] https://github.com/CRIPAC-DIG/RecTextAttack (Zhang et al., 2024c) and PromptBench [^6] https://github.com/microsoft/promptbench (Zhu et al., 2023).
+Table 4.
+Examples of prompting templates of the LLM-based agentic victim models.
 The blue italics represent variables such as the memories of user agents and item agents optimized via agent-environment interactions.
@@
 10. [Title of the least favorite candidate CD]”}
-C.1.4 Configuration of DrunkAgent
-As discussed in Section 3.1, we apply the latest open-source model LLaMA 3 for building the LLM backbone of the surrogate model $f_{\text{Agent}_{\Phi}}(\cdot)$.
-Its collection spans different sizes, from 8B, 70B to 405B.
-To enhance the efficiency and ensure the quality of the attack, we use the lightweight version: meta-llama/Meta-Llama-3-8B-Instruct [^7] https://huggingface.co/meta-llama/Meta-Llama-3-8B-Instruct.
-OpenAI’s gpt-4-turbo [^8] https://platform.openai.com/docs/models/gpt-4-turbo [52] is adopted to implement $f_{\text{LLM}_{\Psi}}(\cdot)$, because it is a state-of-the-art model with powerful language understanding, reasoning and generation capabilities.
-We set $\mathcal{E}=20$, |~$|\widetilde{\mathcal{M}}_{c}|=10$ and $n=5$.
-All random seeds in the evaluation are set to 2024 and the victim LLMs’ temperature is set to 0 for reproducibility.
-For all the other parameters, we keep the default values.
-C.1.5 Evaluation Metrics
-We use two widely-used ranking metrics to evaluate the attack transferability and stealthiness: hit ratio (HR@$\mathcal{K}\uparrow$) and normalized discounted cumulative gain (NDCG@$\mathcal{K}\uparrow$) [53, 19].
-HR@$\mathcal{K}$ measures whether the target/test item is appeared in the top-$\mathcal{K}$ recommendation list of the normal users, while NDCG@$\mathcal{K}$ indicates the ranking position of the target/test item in the list.
-As such, a high HR@$\mathcal{K}$ and a NDCG@$\mathcal{K}$ demonstrate high attack effectiveness, and vice versa.
-To clearly show the performance gap among different attacks, we set $\mathcal{K}$ to 1, 2, 3.
-We also introduce standard sentence perplexity score $\downarrow$ [13] that is a commonly-adopted metric to evaluate the attack imperceptibility.
-The perplexity measures a language model’s uncertainty by taking the exponential of the average negative log-likelihood of the words in a sequence, given their preceding context.
-Text with low perplexity is more fluent and natural, making it harder to detect as manipulated or adversarial, while text with high perplexity is more likely to be detected as unusual or abnormal.
-C.2 Supplements to Performance Comparison
-C.2.1 Attack Universality across Different Target Items.
-In addition to cross-model transferability, we evaluate the cross-sample transferability of DrunkAgent.
-We incorporate the generated adversarial descriptions into totally different target items that are randomly sampled from each dataset.
-HR@1 (i.e., NDCG@1) of DrunkAgent across different models and target items on real-world datasets are shown in Fig.
-6, which indicates its commendable transferability and triggers’ universality across different target items.
-This may be attributed to the fact that our triggers encapsulate diverse and general item characteristics, and, crucially, they also reveal the intended attack objectives.
-Figure 6: Attack Universality across Different Target Items
-Table 5: Ablation Studies of DrunkAgent
+Table 5. Ablation Studies of DrunkAgent
 Dataset Component HR@1 HR@2 HR@3 NDCG@1 NDCG@2 NDCG@3
@@
 0.2268 0.2268 0.2371 0.2268 0.2268 0.2320
-Table 6: Success Rates to Get the Target Agents Drunk
+Table 6. Success Rates to Get the Target Agents Drunk
 Dataset #2 #5 #10 #20
@@
 Musical Instruments 100% 100% 100% 100%
-C.2.2 Ablation Studies
+A.1.3. Victim LLM-powered Agentic Recommender Systems
+LLM-driven agents, known for their superior autonomous interaction and decision-making capabilities, are gradually being considered as next-generation RSs (Huang et al., 2025a).
+Compared with some agent-powered RSs that focus solely on user-side behavior modeling using universal LLMs (Wang et al., 2024c; Huang et al., 2024), AgentCF (Zhang et al., 2024b) emphasizes the modeling of two-sided interaction relations between user agents and item agents through the idea of collaborative filtering (He et al., 2017).
+Both kinds of agents are equipped with memory modules, maintaining the simulated preferences and tastes of potential adopters involving their intrinsic features and acquired behavioral information via autonomous interactions and collaborative optimizations at each time step, as shown on the left of Fig.
+1.
+As such, the memories of the user agents and item agents can be accordingly updated, enabling the agents to better fit the real-world interaction behaviors.
+To better adapt to the dynamically changing real-world environment, the agentic RSs are equipped with real-time capabilities.
+Whenever new descriptions are introduced at a time step, they are integrated into the evolving memories to maintain up-to-date contextual understanding.
+As can be seen on the right of Fig.
+1, the recommendation agents conduct personalized recommendations based on the optimized memories.
+To drive varied architectural designs and support diverse tasks for a comprehensive evaluation of the transferability of attacks, powerful mechanisms (e.g., retrieval augmentation) and task-specific prompts (e.g., sequential recommendation) are introduced to strengthen the agentic RSs.
+As such, AgentRAG and AgentSEQ are developed (Zhang et al., 2024b).
+To maintain a fair evaluation, the implementation details of how the agents can accomplish autonomous interactions and collaborative optimizations are followed the original paper suggest (Zhang et al., 2024b).
+To ensure the high-quality of recommendations, we enhance the prompting strategies for recommendation agents.
+Specifically, we introduce the Chain-of-Thought strategy (Wei et al., 2022) to carefully craft the prompt template $\mathcal{P}$.
+Given the context window constraints pointed out in (Zhang et al., 2024b), a target item and nine non-interactive items of each normal user are randomly selected for the ranking task of the RSs, which is also suggested by the original work.
+The design of recommendation agents is described below.
+  • AgentCF.
+Since both user and item agents are collaboratively optimized to model their two-sided relations, the memories of the user agents and item agents that are optimized by agent-environment interactions can be introduced into $\mathcal{P}$ directly.
+For each normal user $u$,
+    (8) $\mathcal{R}_{u}=f_{\text{AgentCF}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
+    where $m_{u}$ is the short-memory of user agent $u$, $m_{t}$ denotes the features of the target item agent including metadata and the memories, $\mathcal{C}_{t}^{u}$ is the feature set of the candidate item agents and $\mathcal{R}_{u}$ is the ranking result.
+  • AgentRAG.
+Although short-term memory describes the current preference of a user agent, retrieving their specialized preferences from long-term memory toward candidates can allow them to make more personalized inferences.
+For each normal user $u$,
+    (9) $\mathcal{R}_{u}=f_{\text{AgentRAG}}(\mathcal{P}\oplus m_{u}\oplus m_{r}\oplus m_{t}\oplus\mathcal{C}_{t}^{u}),$
+    where $m_{r}$ is retrieved specialized preference from the long-term memory of user agent $u$ by taking the memories of candidate items as queries.
+To ensure the quality while maintaining the efficiency, we use all-MiniLM-L6-v2777https://www.sbert.net/docs/sentence_transformer/pretrained_models.html as the sentence transformer for the similarity calculation.
+  • AgentSEQ.
+When interaction records are sparse and preference propagation is limited, we can further incorporate user historical interactions into prompts, enabling LLMs to serve as sequential recommenders.
+For each normal user $u$,
+    (10) $\mathcal{R}_{u}=f_{\text{AgentSEQ}}(\mathcal{P}\oplus m_{u}\oplus m_{t}\oplus\mathcal{M}_{t}^{u}\oplus\mathcal{C}_{t}^{u}),$
+    where $\mathcal{M}_{t}^{u}$ are the corresponding item features of the historical interactions with user $u$.
+We give some examples of the prompting templates of three victim models in Table 4 for reproductions.
+A.2. Supplements to Performance Comparison
+A.2.1. Ablation Studies
 We remove important components of each module to conduct ablation studies of DrunkAgent.
@@
 However, if we remove it, the setting will transfer from the black-box to the white-box that is impractical and beyond the scope of this paper.
-Moreover, due to gpt-3.5-turbo used for victim models is closed-source and only can be called via APIs, we cannot obtain logits to optimize the adversarial triggers like Llama-3 used in the surrogate model does.
-Hence, we maintain the surrogate model to ensure more practical settings.
+We thus choose to maintain the surrogate model to ensure more practical settings.
 We then remove the greedy optimization algorithm (denoted as $\mathcal{G}_{r}$) in the generation module to observe how the performance changes.
@@
 From the table, we can find that the attack performance decreases by a large margin.
-This may be because memory updates during agent-environment interactions are somewhat robust to non-tailored attacks (this phenomenon can also be observed in the performance of the attack baselines).
-Our reported results are built upon AgentCF, and similar tendencies can be found on AgentRAG and AgentSEQ.
-C.2.3 Parameter Sensitivity
+This may be because memory updates during agent-environment interactions are somewhat robust to non-tailored attacks (this phenomenon can also be observed in the performance of the attack baselines, which are static and one-off).
+The reported results are built upon representative AgentCF, and similar tendencies can be found on AgentRAG and AgentSEQ.
+A.2.2. Parameter Sensitivity
 From Table 5, we can find that the strategy module contributes more than the optimization algorithm on DrunkAgent’s attack performance.
-As such, we increase the maximum rounds of optimization iterations per step of victim models to evaluate the attack success rates of the attack strategy, where the default value is 2 as suggested in the paper [5].
+As such, we increase the maximum rounds of optimization iterations per step of victim models to evaluate the attack success rates of the attack strategy, where the default value is 2 as suggested in the paper (Zhang et al., 2024b).
 As illustrated in Table 6, the success rates of getting the target agents drunk are maintained at 100% regardless of the number of iterations.
 For the other attack parameters, based on our observations, our DrunkAgent is insensitive to them.
-C.2.4 Attack Transferability to Victim Agentic RSs with Powerful LLM backbones
-To further evaluate the transferability of DrunkAgent, we improve the LLM backbones of victim agentic RSs with more powerful gpt-4-turbo.
-Due to such API calls are quite expensive, we evaluate DrunkAgent on a representative real-world dataset.
-To ensure a sufficiently challenging evaluation, we choose the largest dataset: CD & Vinyl.
-Fig.
-7 shows HR@1 (i.e., NDCG@1) of the benign status and DrunkAgent against the three victim models on the dataset, which demonstrates that DrunkAgent is still transferable even if the agentic RSs become powerful.
-We leave experiments on additional datasets and more powerful agentic RSs for future work.
-Figure 7: Attack Transferability on Powerful Victim Agentic RSs.
 (a) AgentCF
@@
 (c) AgentSEQ
-Figure 8: Attack Transferability.
-HR@$\mathcal{K}$ and NDCG@$\mathcal{K}$ of the Benign status and DrunkAgent against various black-box victim LLM-powered agent-based RSs on the real-world Yelp dataset.
-C.2.5 Attack Generalizability across Domains
+Figure 7. Attack Transferability. HRs and NDCGs of the Benign status and DrunkAgent against various black-box victim LLM-powered agent-based RSs on the real-world Yelp dataset.
+A.2.3. Attack Generalizability across Domains
 To further assess the generalizability of DrunkAgent, we conduct additional experiments on the Yelp dataset.
@@
 Despite these domain-specific differences, as can be seen from Fig.
-8, DrunkAgent remains highly effective: it consistently promotes the target items on black-box LLM-powered recommender agents under memory corruptions, confirming the transferability of our attack and the effectiveness of memory perturbations across domains.
+7, DrunkAgent remains highly effective: it consistently promotes the target items on black-box LLM-powered recommender agents under memory corruptions, confirming the transferability of our attack and the effectiveness of memory perturbations across domains.
 This cross-domain success highlights the adaptability and generalizability of DrunkAgent, demonstrating that its impact is not limited to e-commerce settings but extends to real-world platforms driven by experiential content and local service discovery.
-C.2.6 Attack Interpretability
+A.2.4. Attack Interpretability
 We focus on why adversarial attacks exhibit both transferability and stealthiness.
-The transferability arises because adversarial texts are optimized to carry rich as well as general features and explicit attack objectives (e.g., ‘top-level consideration’, ‘a must-have’, ‘prime choice’) to mislead the decisions of the victim models, making them effective across models that share similar language understanding capabilities.
-The stealthiness is attributed to the fact that these texts are optimized to retain meaningful semantics and coherence, and mainly perturb the memories of the target item agents.
-C.2.7 Computing Resourses
-All experiments were run on a institution high-performance computing cluster equipped with NVIDIA H100 GPUs (each with 96GB VRAM), using CUDA version 12.4 and NVIDIA driver version 550.127.05.
-Each individual experimental run (from initialization to evaluation) completed in under 2 hours, with peak GPU memory usage remaining below 30GB.
-These resource requirements apply across all datasets, methods, and target systems evaluated in the paper.
-Appendix D Limitations
-While our work demonstrates the effectiveness of memory-based attacks against agentic recommender systems in controlled experimental settings, there are several limitations to be acknowledged.
-Our evaluation is conducted on publicly available benchmark datasets and simulated agent environments.
-Although these setups capture key characteristics of real-world agentic systems, they may not fully reflect the complexity and scale of industrial deployments.
-In particular, factors such as heterogeneous user behaviors, evolving content ecosystems, and industrial-scale memory management pose additional challenges that are beyond the scope of this study.
-Future work is needed to validate the generalizability and scalability of our approach in production environments.
-In addition, our study focuses on demonstrating the attack effectiveness under a black-box threat model, assuming limited access to internal model parameters or memory states.
-While this setting reflects practical constraints faced by external attackers, a deeper exploration of defense mechanisms and potential mitigation strategies is left for future investigation.
-Despite these limitations, we believe our findings offer valuable insights into the security risks posed by memory mechanisms in agentic recommender systems, and lay a foundation for further research on robust and trustworthy recommendation agents.
-Appendix E Broader Impact
-As agentic recommender systems are emerging as a promising paradigm for next-generation personalization [11, 8], our work contributes positively by proactively identifying potential security vulnerabilities in these systems.
-While such vulnerabilities could lead to potential negative societal impacts if misused, our research serves as a foundational step toward building more robust and trustworthy agentic RSs, enabling early risk awareness and informing effective mitigation strategies, thereby helping to minimize negative societal impacts before real-world deployment.
-Overall, we believe the positive impact of promoting secure and responsible development in this emerging area outweighs the potential risks.
+We attribute the transferability of our adversarial texts to the use of a surrogate model that simulates the victim agent’s behaviors, such as memory updates and recommendations.
+By optimizing on this surrogate, the texts are equipped with transferable update masking strategies and are encapsulated with generalizable attack objectives (e.g., ‘top-level consideration’, ‘a must-have’, ‘prime choice’) that can effectively mislead similar models.
+Their stealthiness arises from maintaining semantic meaningfulness and coherence, which conceals manipulative intent and subtly disrupts prompt parsing.